### v3

# Generalized Linear Bandits with Limited Adaptivity

Abstract
We study the generalized linear contextual bandit problem within the constraints of limited adaptivity. In this paper, we present two algorithms, B-GLinCB and RS-GLinCB, that address, respectively, two prevalent limited adaptivity settings. Given a budget $M$ on the number of policy updates, in the first setting, the algorithm needs to decide upfront $M$ rounds at which it will update its policy, while in the second setting it can adaptively perform $M$ policy updates during its course. For the first setting, we design an algorithm B-GLinCB, that incurs $\tilde{O}(\sqrt{T})$ regret when $M=\Omega\left(\log{\log T}\right)$ and the arm feature vectors are generated stochastically. For the second setting, we design an algorithm RS-GLinCB that updates its policy $\tilde{O}(\log^{2}T)$ times and achieves a regret of $\tilde{O}(\sqrt{T})$ even when the arm feature vectors are adversarially generated. Notably, in these bounds, we manage to eliminate the dependence on a key instance dependent parameter $\kappa$, that captures non-linearity of the underlying reward model. Our novel approach for removing this dependence for generalized linear contextual bandits might be of independent interest.
Keywords Generalized Linear Contextual Bandits, Limited Adaptivity, Optimal Design, Batched Bandits
1 Introduction
Contextual Bandits (CB) is an archetypal framework that models sequential decision making in time-varying environments. In this framework, the algorithm (decision maker) is presented, in each round, with a set of arms (represented as $d$-dimensional feature vectors), and it needs to decide which arm to play. Once an arm is played, a reward corresponding to the played arm is accrued. The regret of the round is defined as the difference between the maximum reward possible in that round and the reward of the played arm. The goal is to design a policy for selecting arms that minimizes cumulative regret (referred to as the regret of the algorithm) over a specified number of rounds, $T$. In the last few decades, much progress has been made in designing algorithms for special classes of reward models, e.g. linear model Auer (2002); Chu et al. (2011); Abbasi-Yadkori et al. (2011); Lattimore and Szepesvári (2020), logistic model Faury et al. (2020); Abeille et al. (2021); Faury et al. (2022); Zhang and Sugiyama (2023) and generalized linear models Filippi et al. (2010); Li et al. (2017).
However, despite this progress, there is a key challenge that prevents deployment of CB algorithms in the real world. Practical situations often allow for very limited adaptivity, i.e., do not allow CB algorithms to update their policy at all rounds. For example, in clinical trials Group et al. (1997), each trial involves administering medical treatments to a cohort of patients, with medical outcomes observed and collected for the entire cohort at the conclusion of the trial. This data is then used to design the treatment for the next phase of the trial. Similarly, in online advertising Schwartz et al. (2017) and recommendations Li et al. (2010), updating the policy after every iteration during deployment is often infeasible due to infrastructural constraints. A recent line of work Ruan et al. (2021); Ren and Zhou (2024); Han et al. (2020); Hanna et al. (2022); Perchet et al. (2015); Gao et al. (2019); Simchi-Levi and Xu (2022) tries to address this limitation by developing algorithms that try to minimize cumulative regret while ensuring that only a limited number of policy updates occur. Across these works, two settings (called M1 and M2 from here onwards) of limited adaptivity have been popular. Both M1, M2 provide a budget $M$ to the algorithm, determining the number of times it can update its policy. In M1 Ruan et al. (2021); Hanna et al. (2023), the algorithm is required to decide upfront a sub-sequence of $M$ rounds where policy updates will occur. While in M2 Abbasi-Yadkori et al. (2011); Ruan et al. (2021)), the algorithm is allowed to adaptively decide (during its course) when to update its policy.
Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting Ruan et al. (2021); Han et al. (2020), and optimal regret guarantees were obtained when the arm feature vectors were stochastically generated. Similarly, in their seminal work on linear bandits, Abbasi-Yadkori et al. (2011) developed algorithms for the M2 setting and proved optimal regret guarantees with no restrictions on the arm vectors. While these results provide tight regret guarantees for linear reward models, extending them to generalized linear models is quite a challenge. Straightforward extensions lead to sub-optimal regret with a significantly worse dependence on an instance dependent parameter $\kappa$ (See Section 2 for definition) that captures non-linearity of the problem instance. In fact, to the best of our knowledge, developing optimal algorithms for the CB problem with generalized linear reward models under the limited adaptivity settings M1, M2, is an open research question. This is the main focus of our work. We make the following contributions.
1.1 Our Contributions
$\bullet$ We propose B-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M1 setting of limited adaptivity. We prove that, when the arm feature vectors are generated stochastically, the regret of B-GLinCB at the end of $T$ rounds is $\tilde{O}(\sqrt{T})$, when $M=\Omega(\log\log{T})$. When $M=O(\log\log{T})$, we prove an $\tilde{O}(T^{2^{M-1}/(2^{M}-2)})$ regret guarantee. While the algorithm bears a slight resemblance to the one in Ruan et al. (2021), direct utilization of their key techniques (distributional optimal design) results in a regret guarantee that scales linearly with the instance dependent non-linearity $\kappa$. On the other hand, the leading terms in our regret guarantee for B-GLinCB have no dependence on $\kappa$. To achieve this, we make novel modifications to the key technique of distributional optimal design in Ruan et al. (2021). Along with this, the rounds for policy updates are also chosen more carefully (in a $\kappa$ dependent fashion) leading to a stronger regret guarantee.
$\bullet$ We propose RS-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M2 setting of limited adaptivity. RS-GLinCB builds on a similar algorithm in Abbasi-Yadkori et al. (2011) by adding a novel context-dependent criterion for determining if a policy update is needed. This new criterion allows us to prove optimal regret guarantee ($\tilde{O}(\sqrt{T})$) with only $O(\log^{2}T)$ updates to the policy. It is quite crucial for the generalized linear reward settings, since without it the resultant regret guarantees have a linear dependence on $\kappa$.
$\bullet$ Our work also resolves a conjecture in Lee et al. (2023) by proving an optimal ($\tilde{O}(\sqrt{T})$) regret guarantee (for the CB problem with logistic reward model) that does not depend polynomially on $S$ (the known upper bound on size of the model parameters, i.e. $\|\theta^{\star}\|\leq S$, See Section 2). RS-GLinCB is, to our knowledge, the first CB algorithm for generalized linear reward models, that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret. We also perform experiments in Section 5 that validate its superiority both in terms of regret and computational efficiency in comparison to other baseline algorithms proposed in Jun et al. (2017) and Faury et al. (2020).
1.2 Important Remarks on Contributions and Comparison with Prior Work
Remark 1.1 ($\kappa$-independence).
For both B-GLinCB and RS-GLinCB, our regret guarantees are free of $\kappa$ (in their leading term), an instance-dependent parameter that can be exponential in the size of the unknown parameter vector, i.e., $\|\theta^{\star}\|$ (See Section 2 for definition). Our contribution in this regard is two-fold. Not only do we prove $\kappa$-independent regret guarantees under the limited adaptivity constraint, we also characterize a broad class of generalized linear reward models for which a $\kappa$-independent regret guarantee can be achieved. Specifically, our results imply that the CB problem with generalized linear reward models originally proposed in Filippi et al. (2010) and subsequently studied in literature (Li et al., 2017; Jun et al., 2017; Russac et al., 2020) admits a $\kappa$-independent regret.
Remark 1.2 (Computational efficiency).
Efforts to reduce the total time complexity to be linear in $T$ have been active in the CB literature with generealized linear rewards models. For e.g., Jun et al. (2017) recently devised computationally efficient algorithms but they suffer from regret dependence on $\kappa$. Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models (Faury et al., 2020; Abeille et al., 2021), and the algorithms were subsequently made computationally efficient in Faury et al. (2022); Zhang and Sugiyama (2023). However, the techniques involved rely heavily on the structure of the logistic model and do not easily extend to more general models. To the best of our knowledge, ours is the first work that achieves optimal $\kappa$-independent regret guarantees for bounded generalized linear reward models while remaining computationally efficient [^1] While RS-GLinCB and B-GLinCB have total running time of $\widetilde{O}(T)$, their per-round complexity can reach $O(T)$. This stands in contrast to Faury et al. (2022), which maintains efficiency in both total and per-round time complexity..
Remark 1.3 (Self Concordance of bounded GLMs).
In order to prove $\kappa$-independent regret guarantees, we prove a key result about self concordance of bounded (almost surely) generalized linear models (Definition 2.1) in Lemma 2.2. This result was postulated in Filippi et al. (2010) for GLMs (with same definition as ours) but no proof was provided. While Faury et al. (2020; 2022) partially tackled this issue for logistic reward models [^2] In Faury (2021), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made, however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
2 Notations and Preliminaries
Notations: A policy $\pi$ is a function that maps any given arm set ${\mathcal{X}}$ to a probability distribution over the same set, i.e., $\pi({\mathcal{X}})\in\Delta({\mathcal{X}})$, where $\Delta({\mathcal{X}})$ is the probability simplex supported on ${\mathcal{X}}$. We will denote matrices in bold upper case (e.g. ${\mathbf{M}}$). $\left\lVert x\right\rVert$ denotes the $\ell_{2}$ norm of vector $x$. We write $\left\lVert x\right\rVert_{{\mathbf{M}}}$ to denote $\sqrt{x^{\top}{\mathbf{M}}x}$ for a positive semi-definite matrix ${\mathbf{M}}$ and vector $x$. For any two real numbers $a$ and $b$, we denote by $a\wedge b$ the minimum of $a$ and $b$. Throughout, $\widetilde{O}(\cdot)$ denotes big-O notation but suppresses log factors in all relevant parameters. For $m,n\in\mathbb{N}$ with $m<n$, we denote the set $\{1,\ldots,n\}$ by $[n]$ and $\{m,\ldots,n\}$ by $[m,n]$.
Definition 2.1 (GLM).
A Generalized Linear Model or GLM with parameter vector $\theta^{\star}\in\mathbb{R}^{d}$ is a real valued random variable $r$ that belongs to the exponential family with density function
$\mathbb{P}(r\mid x)=\exp\big{(}r\cdot\langle x,\theta^{*}\rangle-b\left(\langle x,\theta^{*}\rangle\right)+c\left(r\right)\big{)}$
Function $b$ (called the log-partition function) is assumed to be twice differentiable and $\dot{b}$ is assumed to be monotone. Further, we assume that $r\in[0,R]$ almost surely for some known $R\in\mathbb{R}$.
Important properties of GLMs such as $\operatorname*{\mathbb{E}}[r]=\dot{b}(\langle x,\theta^{\star}\rangle)$ and variance $\mathbb{V}[r]=\ddot{b}(\langle x,\theta^{\star}\rangle)$ are detailed in Appendix C. We define the link function $\mu$ as $\mu\left(\langle x,\theta^{*}\rangle\right)\coloneqq\operatorname*{\mathbb{E}}[r]$. Thus, $\mu$ is also monotone. We now present a key Lemma on GLMs (see Appendix C for details) that enable us to achieve optimal regret guarantees for our algorithms designed in Sections 3 and 4.
Lemma 2.2 (Self-Concordance of GLMs).
For any GLM supported on $[0,R]$ almost surely, the link function $\mu(\cdot)$ satisfies $\lvert\ddot{\mu}(z)\rvert\leq R\dot{\mu}(z)$, for all $z\in\mathbb{R}$.
Next we describe the two CB  problems with GLM rewards that we address in this paper. Let $T\in\mathbb{N}$ be the total number of rounds. At round $t\in[T]$, we receive an arm set ${\mathcal{X}}_{t}\subset\mathbb{R}^{d}$, with number of arms $K=|{\mathcal{X}}_{t}|$ and must select an arm $x_{t}\in{\mathcal{X}}_{t}$. Following this, we receive a reward $r_{t}$ sampled from the GLM distribution $\mathbb{P}(r|x_{t})$ with unknown $\theta^{*}$.
Problem 1: In this problem we assume that at each round $t$, the set of arms ${\mathcal{X}}_{t}\subset\mathbb{R}^{d}$ is drawn from an unknown distribution $\mathcal{D}$. Further, we assume the constraints of limited adaptivity setting M1, i.e., the algorithm is given a budget $M\in\mathbb{N}$ and needs to decide upfront the $M$ rounds at which it will update its policy. Let $\texttt{supp}\left({\mathcal{D}}\right)$ denote the support of distribution ${\mathcal{D}}$. We want to design an algorithm that minimizes the expected cumulative regret given as
$\textsc{R}_{T}=\operatorname*{\mathbb{E}}\big{[}\sum_{t=1}^{T}\max_{x\in{\mathcal{X}}_{t}}\mu\left(\langle x,\theta^{*}\rangle\right)\ -\ \sum_{t=1}^{T}\mu\left(\langle x_{t},\theta^{*}\rangle\right)\big{]}$
Here, the expectation is taken over the randomness of the algorithm, the distribution of rewards $r_{t}$, and the distribution of the arm set ${\mathcal{D}}$.
Problem 2: In this problem we do not make any assumptions on the arm feature vectors, i.e., the arm vectors can be adversarially chosen. However, we assume the constraints of limited adaptivity setting M2, i.e., the algorithm is given a budget $M\in\mathbb{N}$ and needs to adaptively decide the $M$ rounds at which it will update its policy (during its course). We want to design an algorithm that minimizes the cumulative regret given as
$\textsc{R}_{T}=\sum_{t=1}^{T}\max_{x\in{\mathcal{X}}_{t}}\mu\left(\langle x,\theta^{*}\rangle\right)\ -\ \sum_{t=1}^{T}\mu\left(\langle x_{t},\theta^{*}\rangle\right)$
2.1 Instance Dependent Non-Linearity Parameters
As in prior works Faury et al. (2020; 2022), we define instance dependent parameters that capture non-linearity of the underlying instance and critically impact our algorithm design. Algorithm 1 (B-GLinCB) that solves Problem $1$, requires three such parameters that are defined using the derivative of the link function $\dot{\mu}(\cdot)$. Specifically, for any arm set ${\mathcal{X}}$, write optimal arm $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\mu\left(\langle x,\theta^{*}\rangle\right)$ and define,
$\kappa\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}({\mathcal{D}})}\max_{x\in{\mathcal{X}}}\ \frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\kappa^{*}}}\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}\left({\mathcal{D}}\right)}\ {\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\widehat{\kappa}}}\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\right]$
Remark 2.3.
These quantities feature prominently in our regret analysis of Algorithm 1. In particular, the dominant term in our regret bound scales as $O(\sqrt{{T}/{{\kappa^{*}}}})$. We also note that ${\widehat{\kappa}}\geq{\kappa^{*}}$; in fact, for specific distributions ${\mathcal{D}}$, the gap between them can be significant. Hence, we also provide a regret upper bound of $O(\sqrt{{T}/{{\widehat{\kappa}}}})$. In this latter case, however, we incur a worse dependence on $d$. Section 3 provides a quantified form of this trade-off.
Algorithm 2 (RS-GLinCB) that solves Problem $2$, requires another such non-linearity parameter $\kappa$, defined as,
$\kappa\coloneqq\max_{x\in\cup_{t=1}^{T}{\mathcal{X}}_{t}}\frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)}$
We note that, here, $\kappa$ is defined considering the parameter vector $\theta^{*}$ in contrast to prior work on logistic bandits Faury et al. (2022), where its definition involved a maximization over all vectors $\theta$ with $\left\lVert\theta\right\rVert\leq S$ (known upper bound of $\|\theta^{\star}\|$). Hence, our definition of $\kappa$ is potentially much smaller and can lead to lower regret, compared to prior works. Standard to the CB literature with GLM rewards, we will assume that tight upper bounds on these parameters is known to the algorithms.
Assumption 2.4.
We make the following additional assumptions which are standard for the CB problem with linear or GLM reward models.
$\bullet$ For every round $t\in[T]$, and each arm $x\in{\mathcal{X}}_{t}$, $\left\lVert x\right\rVert\leq 1$.
$\bullet$ Let $\theta^{*}$ be the unknown parameter of the GLM reward, then $\|\theta^{*}\|\leq S$ for a known constant $S$.
2.2 Optimal Design Policies
G-optimal Design Given an arm set ${\mathcal{X}}$, the G-optimal design policy $\pi_{G}$ is the solution of the following optimization problem: $\min_{\lambda\in\Delta({\mathcal{X}})}\max_{x\in{\mathcal{X}}}{\left|\left|x\right|\right|}^{2}_{\mathbf{U}(\lambda)^{-1}}$, where $\mathbf{U}(\lambda)=\mathbb{E}_{x\sim\lambda}[xx^{\mathsf{T}}]$. Now consider the following optimization problem, also known as the D-optimal design problem: $\max_{\lambda\in\Delta({\mathcal{X}})}\log\mathsf{Det}(\mathbf{U}(\lambda))$. This is a concave maximization problem as opposed to the G-optimal design which is non-convex. We have the following equivalence theorem due to Kiefer and Wolfowitz (Kiefer and Wolfowitz, 1960):
Lemma 2.5 (Keifer-Wolfowitz).
Let ${\mathcal{X}}\subset\mathbb{R}^{d}$ be any set of arms and ${\mathbf{W}}_{G}$ be the expected design matrix, defined as ${\mathbf{W}}_{G}\coloneqq\operatorname*{\mathbb{E}}_{x\sim\pi_{G}({\mathcal{X}})}\left[xx^{\mathsf{T}}\right]$, with $\pi_{G}({\mathcal{X}})$ as the solution to the D-optimal design problem. Then, $\pi_{G}({\mathcal{X}})$ also solves the G-optimal design problem, and for all $x\in{\mathcal{X}}$, $\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}^{2}\leq d$.
Distributional optimal design Notably, the upper bound on $\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$ specified in Lemma 2.5 holds only for the arms $x$ in ${\mathcal{X}}$. When the arm set ${\mathcal{X}}_{t}$ varies from round to round, securing a guarantee analogous to Lemma 2.5 is generally challenging. Nonetheless, when the arm sets ${\mathcal{X}}_{t}$ are drawn from a distribution, it is possible to extend the guarantee, albeit with a worse dependence on $d$; see Section A.2 in Appendix A. Improving this dependence motivates the need of studying Distributional optimal design and towards this we utilize the results of Ruan et al. (2021).
The distributional design policy is defined using a collection of tuples =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i}):p_{1},\ldots,p_{n}\geq 0\text{ and }\sum_{i}p_{i}=1\}$, wherein each ${\mathbf{M}}_{i}$ is a $d\times d$ positive semi-definite matrix and $n\leq 4d\log{d}$. The collection $\mathcal{M}$ is detailed next. Let $\text{softmax}_{\alpha}\left(\left[s_{1},\ldots,s_{k}\right]\right)$ denote the probability distribution where the $i^{th}$ element is sampled with probability $\frac{s_{i}^{\alpha}}{\sum_{j=1}^{k}s_{j}^{\alpha}}$. For a specific =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i})\}_{i=1}^{n}$, and each $i\in[n]$ write $\pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)=\text{softmax}_{\alpha}(\{\left\lVert x\right\rVert_{{\mathbf{M}}_{i}}^{2}:x\in{\mathcal{X}}\})$. Finally, with $\pi_{G}$ as the G-optimal design policy (Section 2.2), we define the Distributional optimal design policy $\pi$ as
$\pi\left({\mathcal{X}}\right)=\begin{cases}\pi_{G}\left({\mathcal{X}}\right)&\text{ with probability $1/2$}\\ \pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)&\text{ with probability $p_{i}/2$}\end{cases}$
Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of (Ruan et al., 2021) to find the collection $\mathcal{M}$; see Algorithm 4 of Ruan et al. (2021). Overall, the computed $\mathcal{M}$ induces a policy $\pi$ that upholds the following guarantee.
Lemma 2.6 (Theorem 5, Ruan et al. (2021)).
Let $\pi$ be the Distributional optimal design policy that has been learnt from $s$ independent samples ${\mathcal{X}}_{1},\ldots{\mathcal{X}}_{s}\sim{\mathcal{D}}$. Also, let ${\mathbf{W}}$ denote the expected design matrix, ${\mathbf{W}}=\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi({\mathcal{X}})}\left[xx^{\mathsf{T}}\mid{\mathcal{X}}\right]\right]$. Then,
$\mathbb{P}\left\{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}\right]\leq O\left(\sqrt{d\log{d}}\right)\right\}\geq 1-\exp\left(O\left(d^{4}\log^{2}d\right)-sd^{-12}\cdot 2^{-16}\right).$
3 B-GLinCB
In this section, we present B-GLinCB (Algorithm 1) that solves Problem 1 described in Section 2, which enforces constraints of limited adaptivity setting M1. Given limited adaptivity budget $M\in\mathbb{N}$, our algorithm first computes the batch length for each of the $M$ batches (i.e., rounds where the policy remains constant). We build upon the batch length construction in Gao et al. (2019); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.
Batch Lengths: For each batch $k\in[M]$, let ${\mathcal{T}}_{k}$ denote all the consecutive rounds within the $k^{th}$ batch. We will refer to the first batch ${\mathcal{T}}_{1}$ as the warm-up batch. The batch lengths $\tau_{k}\coloneqq|{\mathcal{T}}_{k}|$, $k\in[M]$ are calculated as follows:
$\tau_{1}:=\left(\frac{\sqrt{\kappa}\ e^{3S}d^{2}\gamma^{2}}{S}\alpha\right)^{2/3},$ $\quad\tau_{2}:=\alpha,\quad\tau_{k}:=\alpha\sqrt{\tau_{k-1}},\;\text{for }k\in[3,M]$ (1)
where $\gamma\coloneqq 30RS\sqrt{d\log{T}}$  [^3] Recall that $R$ provides an upper bound on the stochastic rewards and $S$ is an upper bound on the norm of $\theta^{*}$. and $\alpha=T^{\frac{1}{2(1-2^{-M+1})}}$ if $M\leq\log{\log T}$ and $\alpha=2\sqrt{T}$ otherwise.
During the warm-up batch (Lines 2, 4), the algorithm follows the G-optimal design policy, $\pi_{G}$. At the end of the warm-up batch (Line 5), the algorithm computes the Maximum Likelihood Estimate (MLE), $\widehat{\theta}_{w}$, of $\theta^{*}$ [^4] In case the MLE lies outside the set $\{\theta^{*}:\left\lVert\theta^{*}\right\rVert\leq S\}$, we follow the projection step detailed in Appendix E., and design matrix $\mathbf{V}\coloneqq\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}$, with parameter $\lambda=20Rd\log{T}$.
Now, for each batch $k\geq 2$ and every round $t\in{\mathcal{T}}_{k}$, the algorithm updates ${\mathcal{X}}_{t}$ by eliminating arms from it using the confidence bounds (see Equation (5)) computed in the previous batches (Line 12). The algorithm next computes $\widetilde{{\mathcal{X}}}_{t}$, a scaled version of ${\mathcal{X}}_{t}$, as follows
$\widetilde{{\mathcal{X}}}_{t}\coloneqq\left\{\sqrt{{\dot{\mu}(\langle x,\widehat{\theta}_{w}\rangle)}/{\beta(x)}}\ \ \ x:\text{ for all }\ x\in{\mathcal{X}}_{t}\right\}.$ (2)
Finally, we use the distributional optimal design policy $\pi_{k}$, on the scaled arm set $\widetilde{{\mathcal{X}}}_{t}$, to sample the next arm (Line 13). At the end of every batch, we equally divide the batch ${\mathcal{T}}_{k}$ into two sets ${\mathcal{A}}$ and ${\mathcal{B}}$. We use samples from ${\mathcal{A}}$ to compute the estimator $\widehat{\theta}_{k}$ and the scaled design matrix ${\mathbf{H}}_{k}$. The rounds in ${\mathcal{B}}$ are used to compute $\pi_{k+1}$, the distributional optimal design policy for the next batch. It is important to note, while the policy $\pi_{k}$ is utilized in each round (Line 13) to draw arms, it is updated (to $\pi_{k+1}$) only at the end of the batch. Hence, conforming to setting M1, the algorithm updates the selection policy at $M$ rounds that were decided upfront.
Confidence Bounds: The scaled design matrix ${\mathbf{H}}_{k}$, an estimator of the Hessian, is computed at the end of each batch $k\in{2,\ldots,M}$ (Line 16):
${\mathbf{H}}_{k}=\sum_{t\in\mathcal{A}}\left({\dot{\mu}(\langle x_{t},\widehat{\theta}_{w}\rangle)}/{\beta(x_{t})}\right)x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}},\quad\text{where}\quad\beta(x)=\exp\left(R\min\left\{2S,\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right\}\right)$ (3)
where $\mathcal{A}$ is the first half of ${\mathcal{T}}_{k}$. Using this, we define the upper and lower confidence bounds ($UCB_{k}$ and $LCB_{k}$) computed at the end of batch ${\mathcal{T}}_{k}$:
$UCB_{k}(x)$$\coloneqq\begin{cases}\langle x,\widehat{\theta}_{w}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}&k=1\\ \langle x,\widehat{\theta}_{k}\rangle+\gamma\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}&k>1\end{cases},$ (4)
$LCB_{k}(x)$$\coloneqq\begin{cases}\langle x,\widehat{\theta}_{w}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}&k=1\\ \langle x,\widehat{\theta}_{k}\rangle-\gamma\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}&k>1\end{cases}$ (5)
Remark 3.1.
The confidence bounds employed by the algorithm exhibit a significant distinction between the first batch and subsequent batches. While the first batch’s bounds are influenced by the parameter $\kappa$, subsequent batches utilize $\kappa$-independent bounds. This difference arises from the use of the standard design matrix $\mathbf{V}$ in the first batch and a scaled design matrix ${\mathbf{H}}_{k}$ (equation 3) in later batches, leveraging the self-concordance property of GLM rewards to achieve $\kappa$-independence. Notably, the first batch’s confidence bounds influence the scaling factor $\beta(x)$ in later batches, creating a trade-off (addressed in the regret analysis in Appendix A) where an inaccurate estimate of $\widehat{\theta}_{w}$ can exponentially increase the scaling factor and confidence bounds.
In Theorem 3.2 and Corollary 3.3, we present our regret guarantee for B-GLinCB. Detailed proofs for both are provided in Appendix A. The computational efficiency of B-GLinCB is discussed in Appendix  D.
Algorithm 1 B-GLinCB: Batched Generalized Linear Bandits Algorithm
Input: Number of batches $M$ and horizon of play $T$.
1:  Initialize batches ${\mathcal{T}}_{1},\ldots,{\mathcal{T}}_{M}$, as defined in equation (1), and set $\lambda\coloneqq 20Rd\log{T}$.
2:  for rounds $t\in{\mathcal{T}}_{1}$ do
3:     Observe arm set ${\mathcal{X}}_{t}$, sample arm $x_{t}\sim\pi_{G}({\mathcal{X}}_{t})$, and observe reward $r_{t}$.
4:  end for
5:  Compute $\widehat{\theta}_{w}=\operatorname*{arg\,min}_{\begin{subarray}{c}\theta\end{subarray}}\sum_{s\in{\mathcal{T}}_{1}}\ell(\theta,x_{s},r_{s})$ and matrix $\mathbf{V}=\lambda{\mathbf{I}}+\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}$.
6:  Initialize policy $\pi_{1}$ as G-optimal design.
7:  for batches $k=2$ to $M$ do
8:     for  each round $t\in{\mathcal{T}}_{k}$ do
9:        Observe arm set ${\mathcal{X}}_{t}$.
10:        for $j=1$ to $k-1$ do
11:           Update arm set ${\mathcal{X}}_{t}\leftarrow{\mathcal{X}}_{t}\setminus\{x\in{\mathcal{X}}_{t}:{UCB_{j}}(x)<\max_{y\in{\mathcal{X}}_{t}}LCB_{j}(y)\}$.
12:        end for
13:        Scale ${\mathcal{X}}_{t}$, as in (2), to obtain $\widetilde{{\mathcal{X}}_{t}}$., then sample $x_{t}\sim\pi_{k-1}\left(\widetilde{{\mathcal{X}}_{t}}\right)$.
14:     end for
15:     Equally divide ${\mathcal{T}}_{k}$ into two sets $\mathcal{A}$ and $\mathcal{B}$.
16:     Define ${\mathbf{H}}_{k}=\lambda{\mathbf{I}}+\sum_{t\in\mathcal{A}}\frac{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}{\beta(x_{t})}x_{t}x_{t}^{\mathsf{T}}$, and $\widehat{\theta}_{k}=\operatorname*{arg\,min}_{\theta}\sum_{s\in\mathcal{A}}\ell(\theta,x_{s},r_{s})$.
17:     Compute Distributional optimal design policy $\pi_{k}$ using the arm sets $\{{\mathcal{X}}_{t}\}_{t\in\mathcal{B}}$.
18:  end for
Theorem 3.2.
Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$, where
$\textsc{R}_{1}$ $=O\left(RSd\ \left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)T^{\frac{1}{2\left(1-2^{1-M}\right)}}\log{T}\right)\text{ and }$
$\textsc{R}_{2}$ $=O\left(\kappa^{1/3}d^{2}e^{2S}(RS\log{T})^{2/3}T^{\frac{1}{3\left(1-2^{1-M}\right)}}\right).$
Corollary 3.3.
When the number of batches $M\geq\log{\log{T}}$, Algorithm 1 achieves a regret bound of
$\textsc{R}_{T}$ $\leq\widetilde{O}\left(\left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)dRS\sqrt{T}+d^{2}e^{2S}(S^{2}R^{2}\kappa T)^{1/3}\right).$
Remark 3.4.
Scaling the arm set (as in (2)) for optimal design is a crucial aspect of our algorithm, allowing us to obtain tight estimates of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemma A.10). This result relies on multiple novel ideas and techniques, including self-concordance for GLMs, matrix concentration, Bernstein-type concentration for the canonical exponential family (Lemma A.1), and application of distributional optimal design on scaled arm set.
Remark 3.5.
The $\kappa$-dependent batch construction is a crucial feature of our algorithm, enabling effective estimation of $\dot{\mu}{\langle\theta,x\rangle}$ at the end of the first batch. Since the first batch incurs regret linear in its length, achieving a $\kappa$-independent guarantee requires the first batch to be $o(\sqrt{T})$. We demonstrate that choosing $\tau_{1}=O(T^{\frac{1}{3}})$ is sufficient for this purpose (see Appendix A).
4 RS-GLinCB
Algorithm 2 RS-GLinCB: Rarely-Switching GLM Bandit Algorithm
1:  Initialize: $\mathbf{V}={\mathbf{H}}_{1}=\lambda{\mathbf{I}}$, ${\mathcal{T}}_{o}=\emptyset$, $\tau=1$, $\lambda\coloneqq d\log({T/\delta})/R^{2}$ and $\gamma\coloneqq 25RS\sqrt{d\log\left(\frac{T}{\delta}\right)}$.
2:  for rounds $t=1,\dots,T$ do
3:     Observe arm set ${\mathcal{X}}_{t}$.
[…]
5:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and observe reward $r_{t}$.
6:        Update ${\mathcal{T}}_{o}\leftarrow{\mathcal{T}}_{o}\cup\{t\}$, $\mathbf{V}\leftarrow\mathbf{V}+x_{t}x_{t}^{\mathsf{T}}$ and ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}$.
7:        Compute $\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\theta}\sum_{s\in{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$.
8:     else
[…]
10:           Set $\tau=t$ and $\widetilde{\theta}\leftarrow\operatorname*{arg\,min}_{\theta}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$ and
11:           $\widehat{\theta}_{\tau}\leftarrow\texttt{Project}(\widetilde{\theta})$.
12:        end if
13:        Update ${\mathcal{X}}_{t}\leftarrow{\mathcal{X}}_{t}\setminus\{x\in{\mathcal{X}}_{t}:UCB_{o}(x)<\max_{z\in{\mathcal{X}}_{t}}LCB_{o}(z)\}$.
14:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})$ and observe reward $r_{t}$.
15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{w}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
16:     end if
17:  end for
In this section we present RS-GLinCB (Algorithm 2) that solves Problem 2 described in Section 2, which enforces constraints of limited adaptivity setting M2. This algorithm incorporates a novel switching criterion (Line 4), extending the determinant-doubling approach of Abbasi-Yadkori et al. (2011). Additionally, we introduce an arm-elimination step (Line 13) to obtain tighter regret guarantees. Throughout this section, we set $\lambda=d\log({T/\delta})/R^{2}$ and $\gamma=25RS\sqrt{d\log\left({T}/{\delta}\right)}$.
At round $t$, on receiving an arm set ${\mathcal{X}}_{t}$, RS-GLinCB first checks the Switching Criterion I (Line 4). This criterion checks whether for any arm $x\in{\mathcal{X}}_{t}$ the quantity $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ is greater than a carefully chosen $\kappa$-dependent threshold. Here $\mathbf{V}$ is the design matrix corresponding to all arms that have been played in the rounds in ${\mathcal{T}}_{o}$ ($\coloneqq$ the set of rounds preceding round $t$, where Switching Criterion I was triggered). Under this criterion the arm that maximizes $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ is played (call this arm $x_{t}$) and the corresponding reward is obtained. Subsequently in Line 6, the set ${\mathcal{T}}_{o}$ is updated to include $t$; the design matrix $\mathbf{V}$ is updated as $\mathbf{V}\leftarrow\mathbf{V}+x_{t}x_{t}^{\intercal}$; and the scaled design matrix ${\mathbf{H}}_{t+1}$ is set to ${\mathbf{H}}_{t}$. The MLE is computed (Line 7) based on the data in the rounds in ${\mathcal{T}}_{o}$ to obtain $\widehat{\theta}_{o}$.
When Switching Criterion I is not triggered, the algorithm first checks (Line 9) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered). If Switching Criterion II is triggered at round $t$, then in Line 10, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$. Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^5] This optimization problem is non-convex. However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$,
$\min_{\theta}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}\right\rVert_{{\mathbf{H}}(\theta)}\text{s.t. }\left\lVert\theta-\widehat{\theta}_{w}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (6)
Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$. After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$. For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}\sqrt{d\log{({T}/{\delta})}}$, and plays the arm $x_{t}$ with the highest index (Line 14). After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 15) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$. With this, the round $t$ ends and the algorithm moves to the next round. Next, in Lemma 4.1 and Theorem 4.2 we present the guarantees on number of policy updates and regret, respectively, for RS-GLinCB. Detailed proofs for both are provided in Appendix B.
Lemma 4.1.
RS-GLinCB (Algorithm 2), during its entire execution, updates its policy at most $O(R^{4}S^{2}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
Theorem 4.2.
Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}d\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{RT}/{\delta}}\right)+\ \kappa d^{2}R^{5}S^{2}\log^{2}\left({{T}/{\delta}}\right)\big{)}.$
Remark 4.3.
Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting. Unlike existing literature (Faury et al., 2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.4 and B.5). We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.3), hence incurring a small regret in these rounds.
Remark 4.4.
Unlike (Abbasi-Yadkori et al., 2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$). The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret. This modification is a crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
Remark 4.5.
The feasible set for the optimization stated in 6 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability. Deviating from existing literature on GLM Bandits which projects the estimate into the ball set of radius $S$ ($\{\theta:\left\lVert\theta\right\rVert\leq S\}$), our projection step leads to tighter regret guarantees; notably, the leading $\sqrt{T}$ term is free of parameters $S$ (and $R$). This resolves the conjecture made in Lee et al. (2023) regarding the possibility of obtaining $S$-free regret in the $\sqrt{T}$ term in logistic bandits.
5 Experiments
Figure 1: Top: Cumulative Regret vs. number of rounds for Logistic (left) and Probit (right) reward models. Bottom: (left) Execution times of ECOLog and RS-GLinCB for different values of $\kappa$ (low $\kappa=9.3$ and high $\kappa=141.6$) for Logistic rewards. (right) Execution times of GLOC and RS-GLinCB for different values of $\kappa$ (low $\kappa=17.6$ and high $\kappa=202.3$) Probit rewards.
We tested the practicality of our algorithm RS-GLinCB against various baselines for logistic and generalized linear bandits. For these experiments, we adjusted the warm-up threshold constant in RS-GLinCB to $0.01$ and used data from both warm-up and non-warm-up rounds to estimate $\widetilde{\theta}$. These modifications do not affect the overall efficiency as $\widetilde{\theta}$ is calculated only $O(\log(T))$ times. The experiment code is available at https://github.com/nirjhar-das/GLBandit_Limited_Adaptivity.
Logistic. We compared RS-GLinCB against ECOLog (Faury et al., 2022) and GLOC (Jun et al., 2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting. The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$. Arms were sampled uniformly from the $d$-dimensional unit ball. We ran simulations for $T=20,000$ rounds, repeating them 10 times. RS-GLinCB showed the smallest regret with a flattened regret curve, as seen in Fig. 1 (top-left).
Probit. For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB (Filippi et al., 2010). The dimension was set to $d=5$ and number of arms per round to $K=20$. $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=3$. Arm features were generated similarly as in the logistic bandit simulation. We ran simulations for $T=5,000$ rounds, repeating them 10 times. RS-GLinCB outperformed both baselines, as shown in Fig. 1 (top-right).
Comparing Execution Times. We compared the execution times of RS-GLinCB and ECOLog. We created two logistic bandit instances with $d=5$ and $K=20$, and different $\kappa$ values. We ran both algorithms for $T=20,000$ rounds, repeating each run 20 times. For low $\kappa$, RS-GLinCB took about one-fifth of the time of ECOLog, and for high $\kappa$, slightly more than one-third, as seen in Fig. 1 (left-bottom). This demonstrates that RS-GLinCB has a significantly lower computational overhead compared to ECOLog. We also compared the execution times of RS-GLinCB and GLOC under the probit reward model, creating two bandit instances with $d=5$ and $K=20$, but with differing $\kappa$. We ran both algorithms for $T=20,000$ rounds, repeating each run 20 times. The result is shown in Fig. 1 (bottom-right). We observe that for low $\kappa$, RS-GLinCB takes less than half time of GLOC while for high $\kappa$, it takes about two-third time of GLOC. A more detailed discussion of these experiments is provided in Appendix D.
6 Conclusion and Future Work
The Contextual Bandit problem with GLM rewards is a ubiquitous framework for studying online decision-making with non-linear rewards. We study this problem with a focus on limited adaptivity. In particular, we design algorithms B-GLinCB and RS-GLinCB that obtain optimal regret guarantees for two prevalant limited adaptivity settings M1 and M2 respectively. A key feature of our guarantees are that their leading terms are independent of a instance dependent parameter $\kappa$ that captures non-linearity. To the best of our knowledge, our paper provides the first algorithm for CB algorithms for GLM rewards under limited adaptivity (and otherwise) that achieves $\kappa$-independent regret. The regret guarantee of RS-GLinCB, not only aligns with the best-known guarantees for Logistic Bandits but enhances them by removing the dependence on $S$ (upper bound on $\left\lVert\theta^{*}\right\rVert$) in the leading term of the regret and therefore resolves a conjecture in Lee et al. (2023). The batch learning algorithm B-GLinCB, for $M=\Omega(\log{(\log{T})})$, achieves a regret of $\widetilde{O}\left(dRS\left(\sqrt{{d/{\widehat{\kappa}}}}\land\sqrt{1/\kappa^{*}}\right)\sqrt{T}\right)$. We believe that the dependence on $d$ along with the ${\widehat{\kappa}}$ term is not tight and improving the dependence is a relevant direction for future work.
Appendix for “”
Appendix A Regret Analysis of B-GLinCB
Additional Notation: We write c to denote absolute constant(s) that appears throughout our analysis. Our analysis also utilizes the following function
$\gamma(\lambda)=24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}.$ (7)
Note that $\gamma(\lambda)$ is a ‘parameterized’ version of $\gamma$ (which was defined in section 3). In our proof, we present the arguments using this parameterized version. A direct minimization of the above expression in terms of $\lambda$ would not suffice since we need $\lambda$ to be sufficiently large for certain matrix concentration lemmas to hold (see Section A.2). However, later we show that setting $\lambda$ equal to $\textbf{c}Rd\log{T}$ leads to the desired bounds.
We use $\widetilde{x}$ to denote the scaled versions of the arms (see Line 13 of the algorithm); in particular,
$\widetilde{x}\coloneqq\sqrt{\frac{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}{\beta(x)}}x$ (8)
Furthermore, to capture the non-linearity of the problem, we introduce the term $\phi(\lambda)$:
$\phi(\lambda)\coloneqq\frac{\sqrt{\kappa}\ e^{3S}\ \gamma(\lambda)^{2}}{S}.$
Recall that the scaled data matrix ${\mathbf{H}}_{k}$ (for each batch $k$) was computed using $\widehat{\theta}_{w}$ as follows
${\mathbf{H}}_{k}=\sum_{t\in{\mathcal{T}}_{k}}\frac{\dot{\mu}\left(\langle x_{j},\widehat{\theta}_{w}\rangle\right)}{\beta(x_{t})}x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}.$
Following the definition of ${\mathbf{H}}_{k}$ and using the true vector $\theta^{*}$ we define
${\mathbf{H}}^{*}_{k}=\sum_{t\in{\mathcal{T}}_{k}}\dot{\mu}\left(\langle x_{t},\theta^{*}\rangle\right)x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}.$
We will show that ${\mathbf{H}}_{k}$ is an estimator of ${\mathbf{H}}^{*}_{k}$. Furthermore, we assume that the MLE estimator $\theta^{*}$ obtained by minimizing the log-loss objective always satisfies $\left\lVert\widehat{\theta}\right\rVert\leq S$. In case, that’s not true, one can use the non-convex projection described in Appendix E. The projected vector satisfies the same guarantees as described in the subsequent lemmas up to a multiplicative factor of $2$. Hence, the assumption $\left\lVert\widehat{\theta}\right\rVert\leq S$ is non-limiting.
Lemma A.1 (Bernstein’s Inequality).
Let $X_{1},\ldots,X_{n}$ be a sequence of independent random variables with $|X_{t}-\operatorname*{\mathbb{E}}\left[X_{t}\right]|\leq b$. Also, let sum $S:=\sum_{t=1}^{n}\left(X_{t}-\operatorname*{\mathbb{E}}\left[X_{t}\right]\right)$ and $v:=\sum_{t=1}^{m}\text{Var}[X_{t}]$. Then, for any $\delta\in[0,1]$, we have
$\mathbb{P}\left\{S\geq\sqrt{2v\log{\frac{1}{\delta}}}+\frac{2b}{3}\log{\frac{1}{\delta}}\right\}\leq\delta.$
Lemma A.2.
Let ${\mathcal{X}}=\left\{x_{1},x_{2},\dots,x_{s}\right\}\in\mathbb{R}^{d}$ be a set of vectors with $\|x_{t}\|\leq 1$, for all $t\in[s]$, and let scalar $\lambda\geq 0$. Also, let $r_{1},r_{2},\dots,r_{s}\in[0,R]$ be independent random variables distributed by the canonical exponential family; in particular, $\mathbb{E}\left[r_{s}\right]=\mu\left(\langle x_{s},\theta^{*}\rangle\right)$ for $\theta^{*}\in\mathbb{R}^{d}$. Further, let $\widehat{\theta}=\operatorname*{arg\,min}_{\theta}\sum_{s=1}^{t}\ell(\theta,x_{s},r_{s})$ be the maximum likelihood estimator of $\theta^{*}$ and let matrix
${{\mathbf{H}}^{*}}=\sum_{j=1}^{s}\dot{\mu}\left(\langle x_{j},\theta^{*}\rangle\right)x_{j}x_{j}^{\mathsf{T}}+\lambda{\mathbf{I}}.$
Then, with probability at least than $1-\frac{1}{T^{2}}$, the following inequality holds
$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}\leq 24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}$ (9)
Proof.
We first define the following quantities
[…]
${\mathbf{G}}$ $\coloneqq\sum_{j=1}^{s}\alpha(x,\theta^{*},\widehat{\theta})x_{j}x_{j}^{\mathsf{T}}+\frac{\lambda}{1+2RS}{\mathbf{I}}$
Using Lemma C.2 we have
${\mathbf{G}}\succeq\frac{1}{1+2RS}\ {{\mathbf{H}}^{*}}$ (10)
Hence we write
$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}$ $\leq\sqrt{(1+2RS)}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}}$
$=\sqrt{1+2RS}\left\lVert\left(\theta^{*}-\widehat{\theta}\right){\mathbf{G}}\right\rVert_{{\mathbf{G}}^{-1}}$
$=\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\langle\theta^{*},x_{j}\rangle-\langle\widehat{\theta},x_{j}\rangle\right)\alpha(x,\theta^{*},\widehat{\theta})x_{j}+\frac{\lambda}{1+2RS}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{\mathbf{G}}^{-1}}+\frac{\lambda}{\sqrt{1+2RS}}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}^{-1}}$
$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{\mathbf{G}}^{-1}}+2S\sqrt{\lambda}$
$\leq 3RS\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+2S\sqrt{\lambda}$ (Using (10) and assuming $RS\geq 1$)
Now by the optimality condition on $\widehat{\theta}$ we have $\sum_{j=1}^{s}\mu\left(\langle x_{j},\widehat{\theta}\rangle\right)x_{j}=\sum_{j=1}^{s}r_{j}x_{j}$ (see equation (3) Filippi et al. [2010]). Hence, we write
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}=\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ (11)
Let ${\mathcal{B}}$ denote the unit ball in $\mathbb{R}^{d}$. We can write
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}=\max_{y\in{\mathcal{B}}}\langle y,{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
We construct an $\varepsilon$-net for the unit ball, denoted as ${\mathcal{C}}_{\varepsilon}$. For any $y\in{\mathcal{B}}$, we define $y_{\varepsilon}\coloneqq\operatorname*{arg\,min}_{b\in{\mathcal{C}}_{\varepsilon}}\left\lVert b-y\right\rVert_{2}$. We can now write
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\!\!$
$=\max_{y\in{\mathcal{B}}}\langle y-y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle\!+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
$\leq\left\lVert y-y_{\varepsilon}\right\rVert_{2}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
$\leq\varepsilon\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
Rearranging, we obtain
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\frac{1}{1-\varepsilon}\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
Next, we use Lemma A.3 (stated below) with $\delta=T^{2}|{\mathcal{C}}_{\varepsilon}|$ and union bound over all vectors in ${\mathcal{C}}_{\varepsilon}$. We also observe that $|{\mathcal{C}}_{\varepsilon}|\leq\left(\frac{2}{\varepsilon}\right)^{d}$. Substituting $\epsilon=1/2$ and using Lemma A.3, we obtain that the following holds with probability greater than $1-\frac{1}{T^{2}}$,
$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq 3{\sqrt{\log{\left(T^{2}|{\mathcal{C}}_{\varepsilon}|\right)}}}+\frac{4R}{3\sqrt{\lambda}}\log{\left(T^{2}|{\mathcal{C}}_{\varepsilon}|\right)}$
$\leq 8\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)$
Substituting in equations (11), we get the desired inequality in the lemma statement. ∎
Lemma A.3.
Let $y$ be a fixed vector with $\left\lVert y\right\rVert\leq 1$. Then, with the notation stated in Lemma A.2, the following inequality holds with probability at least $1-\delta$
$\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)y^{\mathsf{T}}{{\mathbf{H}}^{*}}^{-1/2}x_{j}\leq\sqrt{2\log{\frac{1}{\delta}}}+\frac{2R}{3\sqrt{\lambda}}\log{\frac{1}{\delta}}.$
Proof.
Let us denote the $j^{th}$ term of the sum as $Z_{j}$. Note that each random variable $Z_{j}$ has variance $\text{Var}(Z_{j})=\dot{\mu}\left(\langle x_{j},\theta^{*}\rangle\right)\left(y^{\mathsf{T}}{{\mathbf{H}}^{*}}^{-1/2}x_{j}\right)^{2}$. Hence, we have
$\sum_{j=1}^{s}\text{Var}(Z_{j})$ $=\sum_{j=1}^{s}\dot{\mu}\left(\langle\theta^{*},x_{j}\rangle\right)\left(y^{\mathsf{T}}{{\mathbf{H}}^{*}}^{-1/2}x_{j}\right)^{2}$
$=y^{\mathsf{T}}y\leq 1.$
Moreover, each $Z_{j}$ is at most $\frac{R}{\sqrt{\lambda}}$ (since $\left\lVert x_{j}\right\rVert\leq 1$, ${{\mathbf{H}}^{*}}\succeq\lambda{\mathbf{I}}$ and $r\in[0,R]$). Now applying Lemma A.1, we have
$\mathbb{P}\left\{\sum_{j=1}^{s}Z_{j}\geq\sqrt{2\log{\frac{1}{\delta}}}+\frac{2R}{3\sqrt{\lambda}}\log{\frac{1}{\delta}}\right\}\leq\delta.$
Corollary A.4.
Let $x_{1},x_{2},\ldots,x_{\tau}$ be the sequence of arms pulled during the warm-up batch and let $\widehat{\theta}_{w}$ be the estimator of $\theta^{*}$ computed at the end of the batch. Then, for any vector $x$ and $\lambda\geq 0$ the following bound holds with probability greater than $1-\frac{1}{T^{2}}$
$|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma{\left(\lambda\right)}.$
Proof.
This result is derived directly from Lemma A.2 and the definition of (see ). By applying the lemma, we obtain
$|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{w}}\right\rVert_{{\mathbf{H}}^{*}}\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\gamma(\lambda)$
Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$. This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$ which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. ∎
Lemma A.5.
For each batch $k\geq 2$ and the scaled data matrix ${\mathbf{H}}_{k}$ computed at the end of batch, the following bound holds with probability at least $1-\frac{1}{T^{2}}$:
${\mathbf{H}}_{k}\preceq{\mathbf{H}}^{*}_{k}.$
Proof.
If the event stated in Lemma A.4 holds,
From Lemma C.2, we apply the multiplicative bound on $\dot{\mu}$ to obtain
$\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)$ $\leq\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\exp\left(R|\langle x,\widehat{\theta}_{w}-\theta^{*}\rangle|\right)$
Via Lemma A.4 we have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma(\lambda)$. Additionally, given that $\left\lVert\widehat{\theta}\right\rVert,\left\lVert\theta^{*}\right\rVert\leq S$ and $\left\lVert x\right\rVert\leq 1$ we also have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq 2S$. Hence, we write
$\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)$ $\leq\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\exp\left(R\min\{\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma(\lambda),2S\}\right)$
$\leq\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\beta(x)$
Substituting these results into the definitions of ${\mathbf{H}}_{k}$ and ${\mathbf{H}}^{*}_{k}$ proves the lemma statement. ∎
Claim A.6.
The Algorithm 1 runs for at most $\log{\log{T}}$ batches.
Proof.
When $M\leq\log{\log{T}}$ then the claim trivially holds. When $M\geq\log{\log{T}}+1$, we define the length of the second batch, $\tau_{2}$, as $2\sqrt{T}$. The length of the $M^{th}$ batch is
$\tau_{M}$ $=(2\sqrt{T})^{\sum_{k=1}^{M-1}\frac{1}{2^{k-1}}}$
$\geq 2T\ T^{\frac{-1}{2^{M-1}}}$
$\geq 2T\ T^{\frac{-1}{2^{\log{\log{T}}}}}$ ($M\geq\log{\log{T}}+1$)
$\geq T.$
Corollary A.7.
Let $\widehat{\theta}_{k}$ be the estimator of $\theta^{*}$ calculated at the end of the $k^{th}$ batch. Then for any vector $x$ the following holds with probability greater than $1-\frac{\log{\log T}}{T^{2}}$ for every batch $k>=2$.
$|\langle x,\theta^{*}-\widehat{\theta}\rangle|\leq\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}\gamma{\left(\lambda\right)}$
Proof.
This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 7). According to the lemma, we have
$|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|$ $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{w}}\right\rVert_{{\mathbf{H}}^{*}_{k}}$
$\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\gamma(\lambda)$
Using Lemma A.5, we can further bound $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\leq\left\lVert x\right\rVert_{{{\mathbf{H}}_{k}}^{-1}}$. Finally, a union bound over all batches and considering the fact that there are at most $\log{\log{T}}$ batches (Claim A.6) we establish the corollary’s claim. ∎
Claim A.8.
For any $x\in[0,M]$ the following holds
$e^{x}\leq\left(e^{M}-1\right)\frac{x}{M}+1.$
Proof.
The claim follows from the convexity of $e^{x}$. ∎
Lemma A.9.
Let $x\in{\mathcal{X}}$ be the selected in any round of batch $k\geq 2$ in the algorithm, and let $x^{*}$ be the optimal arm in the arm set ${\mathcal{X}}$, i.e., $x^{*}=\operatorname*{arg\,max}_{x\in}{\mathcal{X}}$. With probability greater than $1-\frac{\log{\log{T}}}{T^{2}}$, the following inequality holds-
$\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\leq 6\phi(\lambda)\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\!\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
Proof.
We begin by applying Taylor’s theorem, which yields the following for some $z\in\left[\langle x^{*},\theta^{*}\rangle,\langle x,\theta^{*}\rangle\right]$
$\left|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\right|$ (12)
$=\dot{\mu}(z)\left|\langle x,\theta^{*}\rangle-\langle x,\theta^{*}\rangle\right|$
$=\dot{\mu}(z)\left|\langle x^{*},\theta^{*}\rangle-\langle x^{*},\widehat{\theta}_{k-1}\rangle+\langle x^{*},\widehat{\theta}_{k-1}\rangle-\langle x,\widehat{\theta}_{k-1}\rangle+\langle x,\widehat{\theta}_{k-1}\rangle-\langle x,\theta^{*}\rangle\right|$
$\leq\dot{\mu}(z)\left(\left|\langle x^{*},\theta^{*}\rangle-\langle x^{*},\widehat{\theta}_{k-1}\rangle\right|+\left|\langle x^{*},\widehat{\theta}_{k-1}\rangle-\langle x,\widehat{\theta}_{k-1}\rangle\right|+\left|\langle x,\widehat{\theta}_{k-1}\rangle-\langle x,\theta^{*}\rangle\right|\right)$
$\leq 2\dot{\mu}(z)\left(\left\lVert x^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\gamma(\lambda)+\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}}\gamma(\lambda)\right)$ (via Lemma A.7)
$\leq 2\dot{\mu}(z)\gamma(\lambda)\left(\sqrt{\frac{\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\sqrt{\frac{\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
$\leq 2\gamma(\lambda)\sqrt{\dot{\mu}(z)}\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\sqrt{\dot{\mu}(z)}\gamma(\lambda)\sqrt{\frac{\dot{\mu}(z)\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}$ (13)
We now invoke Lemmas C.2 and A.4 to obtain
$\sqrt{\frac{\dot{\mu}(z)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}\beta(x)}$ $\leq\sqrt{\exp\left(\min\left\{2S,\left|z-\langle x,\widehat{\theta}_{w}\rangle\right|\right\}\right)\beta(x)}$ (by stated assumptions and Lemma C.2)
$\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-z|}{2}\right\}\right)\sqrt{\beta(x)}$
$\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-\langle x^{*},\theta^{*}\rangle|}{2}\right\}\right)\sqrt{\beta(x)}$ (since $z\in\left[\langle x^{*}_{t},\theta^{*}\rangle,\langle x,\theta^{*}\rangle\right]$ )
$\leq\exp\left(\min\left\{S,\frac{3\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+2\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)}{2}\right\}\right)\sqrt{\beta(x)}$ (using Lemma A.4)
$\leq\exp\left(\min\left\{2S,2\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)\right\}\right)$ (substituting the definition of $\beta(x)$ )
Similarly, we also have
$\sqrt{\dot{\mu}\left(z\right)}\leq\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\exp\left(\min\left\{S,\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)\right\}\right).$
Further, we can simplify each term in equation (13) as
$\sqrt{\dot{\mu}(z)}\left(\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
$\leq 2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\exp\left(\min\left\{3S,3\sqrt{\kappa}\gamma\left(\lambda\right)\left(\left\lVert x\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\right)\right\}\right)\right)$
$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
$\quad\quad+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
$\leq 6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
Finally, we substitute the above bound in (13) to obtain
$|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)|\leq$ $6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
$\quad\quad\quad+2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
For Phase $k$, the distribution of the remaining arms after the elimination step (${\mathcal{X}}$ in line 12 of the Algorithm 1) is represented as ${\mathcal{D}}_{k}$.
Lemma A.10.
During any round in batch $k$ of Algorithm 1, and for an absolute constant $c$, we have
$\operatorname*{\mathbb{E}}\left[|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)|\right]$ $\leq c\left(\frac{\phi(\lambda)d^{2}}{\sqrt{\tau_{1}\ \tau_{k-1}}}+\frac{\gamma(\lambda)}{\sqrt{\tau_{k-1}}}\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)\right)$
Proof.
The proof here invokes Lemma A.9. We begin by noting that
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}$ $\leq 4\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-}1}\right]$
$\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Jensen’s inequality)
$\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k-1}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Claim A.11)
$\leq c\left(\sqrt{\frac{d^{2}}{\tau_{1}}\cdot\frac{d^{2}}{\tau_{k-1}}}\right)$ (using Lemma A.17)
We also have
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)\right]$
$\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
$\leq 2\min\left\{\sqrt{\kappa^{*}}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right],\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}\right\}$ (using the definition of $\kappa^{*}$ for the first bound and Jensen for the second)
$\leq c\left(\sqrt{\frac{d\log{d}}{\kappa^{*}\tau_{k-1}}}\land\sqrt{\frac{d^{2}}{{\widehat{\kappa}}\tau_{k-1}}}\right)$ (using Lemma A.17)
Substituting the above bounds in Lemma A.9 we obtained the stated inequality. This completes the proof. ∎
A.1 Proof of Theorem 3.2
We trivially upper bound the regret incurred during the warm-up batch as $\tau_{1}R$; recall that $R$ denotes the upper bound on the rewards and $\tau_{1}$ denotes the length of the first (warm-up) batch; see equation (3)).
For each batch $k$ and an absolute constant $c$, Lemma A.10 gives us
$\operatorname*{\mathbb{E}}\left[\sum_{t\in{\mathcal{T}}_{k}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)\right]$ $\leq\tau_{k}\cdot c\left(\frac{\phi(\lambda)d^{2}}{\sqrt{\tau_{1}\ \tau_{k-1}}}+\frac{\gamma(\lambda)}{\sqrt{\tau_{k-1}}}\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)\right)$
$\leq c\left(\frac{\phi(\lambda)d^{2}}{\sqrt{\tau_{1}}}\alpha+\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)\gamma(\lambda)\alpha\right)$ (via (1))
Since there are at most $\log{\log T}$ batches, we can upper bound the regret as
$R_{T}\leq c\left(\tau_{1}R+\frac{\phi(\lambda)d^{2}}{\sqrt{\tau_{1}}}\alpha+\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)\gamma(\lambda)\alpha\right)\log{\log(T)}$
Setting $\tau_{1}=\left(\frac{\phi(\lambda)d^{2}\alpha}{R}\right)^{2/3}$ we get
$R_{T}\leq O\left(\left(\frac{\phi(\lambda)d^{2}\alpha}{R}\right)^{2/3}+\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)\gamma(\lambda)\alpha\right)\log{\log(T)}$
Now with the choice of $\lambda=20dR\log T$, we have
$\gamma(\lambda)\leq\gamma\quad\text{ where }\gamma=30RS\sqrt{d\log{T}}.$
Substituting $\alpha=T^{\frac{1}{2\left(1-2^{-M}\right)}}$ and $\phi(\lambda)=\frac{\sqrt{\kappa}\ e^{3S}\ \gamma^{2}}{S}$ we get
$R_{T}\leq O\left(\left(\sqrt{\kappa}d^{3}\ e^{3S}\ RST^{\frac{1}{2\left(1-2^{-M}\right)}}\log{T}\right)^{2/3}+\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)RST^{\frac{1}{2\left(1-2^{-M}\right)}}\sqrt{d\log{T}}\right).$
A.2 Optimal Design Guarantees
In this section, we study the optimal design policies utilized in different batches of the algorithm. Specifically, $\pi_{G}$ denotes the G-optimal design policy applied during the warm-up batch, while $\pi_{k}$ refers to the Distributional Optimal Design policy calculated at the end of batch $k$ ( and used in the $(k+1)^{th}$ batch). Recall that the distribution of the remaining arms after the elimination step ( ${\mathcal{X}}$ in line 12 of the Algorithm) is represented as ${\mathcal{D}}_{k}$. We define expected design matrices for each policy:
${\mathbf{W}}_{G}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi_{G}({\mathcal{X}})}\left[xx^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
${\mathbf{W}}_{k}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi_{k}({\mathcal{X}})}\left[\widetilde{x}\widetilde{x}^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
Recall, for all batches starting from the second batch ($k\geq 2$), we employ the scaled arm set, denoted as $\widetilde{{\mathcal{X}}}$, for learning and action selection under the Distributional optimal design policy. However, during the initial warm-up batch, we utilize the original, unscaled arm set.
Claim A.11.
The following holds for any positive semidefinite matrix ${\mathbf{A}}$ and any batch $k$-
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{A}\leq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{j}}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{A}\quad\forall j\in[k-1].$
This is due to the fact that the set of surviving arms in batch $k$ is always a smaller set than the previous batches.
Lemma A.12 (Lemma 4 Ruan et al. [2021]).
The expected data matrix ${\mathbf{W}}_{G}$ satisfies. We have
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}^{2}\right]\leq d^{2}$
Lemma A.13 (Theorem 5 Ruan et al. [2021]).
Let the Distributional optimal design $\pi$ which has been learnt from $s$ independent samples ${\mathcal{X}}_{1},\ldots{\mathcal{X}}_{s}\sim{\mathcal{D}}$ and let ${\mathbf{W}}$ denote the expected data matrix, ${\mathbf{W}}=\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi({\mathcal{X}})}\left[xx^{\mathsf{T}}|{\mathcal{X}}\right]\right]$. We have
$\mathbb{P}\left\{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}\right]\leq O\left(\sqrt{d\log{d}}\right)\right\}\geq 1-\exp\left(O\left(d^{4}\log^{2}d\right)-sd^{-12}\cdot 2^{-16}\right).$ (14)
Lemma A.14.
Under the notation of Lemma A.13, we have
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2d^{2}.$ (15)
Proof.
Recall that the Distributional optimal design policy samples according to the $\pi_{G}$ policy with half probability. Hence, we have
${\mathbf{W}}=\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi({\mathcal{X}})}\left[xx^{\mathsf{T}}|{\mathcal{X}}\right]\right]\succeq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\frac{1}{2}\operatorname*{\mathbb{E}}_{x\sim\pi_{G}({\mathcal{X}})}\left[xx^{\mathsf{T}}|{\mathcal{X}}\right]\right]=\frac{1}{2}{\mathbf{W}}_{G}$
Therefore,
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2d^{2}.$
Lemma A.15 (Matrix Chernoff Tropp et al. [2015], Ruan et al. [2021]).
Let $x_{1},x_{3},\ldots,x_{n}\sim{\mathcal{D}}$ be vectors, with $\left\lVert x_{t}\right\rVert\leq 1$, then we have
$\mathbb{P}\left\{3\varepsilon n{\mathbf{I}}+\sum_{i=1}^{n}x_{i}x_{i}^{\mathsf{T}}\succeq\frac{n}{8}\ \operatorname*{\mathbb{E}}_{x\sim{\mathcal{D}}}\left[xx^{\mathsf{T}}\right]\right\}\geq 1-2d\exp\left(\frac{-\varepsilon n}{8}\right)$
Corollary A.16.
In Algorithm 1 the warm-up matrix $\mathbf{V}$, with $\lambda\geq 16\log{(Td)}$, satisfies the following with probability greater than $1-\frac{1}{T^{2}}$.
$\mathbf{V}\succeq\frac{\tau_{1}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\operatorname*{\mathbb{E}}_{x\sim\pi_{G}({\mathcal{X}})}xx^{\mathsf{T}}$
Similarly ${\mathbf{H}}_{k}$, with $\lambda\geq 6\log{(Td)}$ satisfies the following for each batch $k\geq 2$ with probability greater than $1-\frac{1}{T^{2}}$
${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{G}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
Proof.
The results for both $\mathbf{V}$ and ${\mathbf{H}}_{k}$ are obtained directly by applying Lemma A.15 with $\varepsilon=\frac{\log(T)}{\tau_{1}}$ and $\varepsilon=\frac{\log(T)}{\tau_{k}}$, respectively. ∎
We note that the analysis of Ruan et al. [2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}}^{2}$. We obtain such a bound here and use it in the analysis.
Lemma A.17.
The following holds with probability greater than $1-\frac{1}{T^{2}}$
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{1}}\right)$ (16)
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{k}}\right)\quad\forall k\in[M]$ (17)
We also have that for sufficiently large $T\gtrsim O\left(d^{32}(\log{2T})^{2}\right)$, the following holds with probability greater than $1-\frac{\log{\log{T}}}{T}$
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}$ $\leq O\left(\sqrt{\frac{d}{\tau_{k}}}\right)\quad\forall k\in[M]$ (18)
Proof.
First, we note from Corollary A.16 that the following holds with high probability
$\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$
$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\geq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
We obtain the first to inequalities, (16) and (17), by a direct use of Corollary A.16. For (18) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M). Since, $T\geq d^{32}\log{2T}^{2}$ the event stated in Lemma A.13 holds with probability greater than $1-\frac{1}{T^{2}}$. ∎
Appendix B Regret Analysis of RS-GLinCB
Recall ${\mathcal{T}}_{o}$ denotes the set of warm-up rounds. We write $\tau_{w}$ to denote the size of the set $\tau_{w}=|{\mathcal{T}}_{o}|$. We define the following (scaled) data matrix
${\mathbf{H}}^{*}_{w}=\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}\ .$
We will specify the regularizer $\lambda$ later. We also define
$\gamma\coloneqq\textbf{c}RS\sqrt{d\log{\frac{T}{\delta}}}$
Below, we state the main concentration bound used in the proof.
Lemma B.1 (Theorem 1 of Faury et al. [2020]).
Let $\{\mathcal{F}_{t}\}_{t=1}^{\infty}$ be a filtration. Let $\{x_{t}\}_{t=1}^{\infty}$ be a stochastic process in $B_{1}(d)$ such that $x_{t}$ is $\mathcal{F}_{t}$ measurable. Let $\{\eta_{t}\}_{t=1}^{\infty}$ be a martingale difference sequence such that $\eta_{t}$ is $\mathcal{F}_{t}$ measurable. Furthermore, assume we have $\lvert\eta_{t}\rvert\leq 1$ almost surely, and denote $\sigma_{t}^{2}=\mathbb{V}[\eta_{t}|\mathcal{F}_{t}]$. Let $\lambda>0$ and for any $t\geq 1$ define:
$S_{t}=\sum_{s=1}^{t-1}\eta_{s}x_{s}\quad\quad\quad{\mathbf{H}}_{t}=\sum_{s=1}^{t-1}\sigma_{s}^{2}x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$
Then, for any $\delta\in(0,1]$,
$\mathbb{P}\left[\exists t\geq 1:\lVert S_{t}\rVert_{{{\mathbf{H}}}^{-1}_{t}}\geq\frac{\sqrt{\lambda}}{2}+\frac{2}{\sqrt{\lambda}}\log\left(\frac{\det({\mathbf{H}}_{t})^{\frac{1}{2}}}{\lambda^{d/2}\delta}\right)+\frac{2}{\sqrt{\lambda}}d\log(2)\right]\leq\delta$
Lemma B.2.
At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the warm-up rounds ${\mathcal{T}}_{o}$. With probability at least $1-\delta$ we have
$\lVert\widehat{\theta}_{w}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma.$
Proof.
Let us define the matrix ${\mathbf{G}}_{w}=\sum_{s\in{\mathcal{T}}_{o}}\alpha(x,\theta^{*},\widehat{\theta}_{o})x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$. First, we note that by self-concordance property of $\mu$ (Lemma C.2), ${\mathbf{G}}_{w}\succeq\frac{1}{1+2RS}{\mathbf{H}}^{*}_{w}$. Hence,
$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq\sqrt{1+2RS}\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{G}}_{w}}$
$=\sqrt{1+2RS}\left\lVert{\mathbf{G}}_{w}\left(\widehat{\theta}_{o}-\theta^{*}\right)\right\rVert_{{\mathbf{G}}_{w}^{-1}}$
$=\sqrt{1+2RS}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\langle\widehat{\theta}_{o},x_{s}\rangle-\langle\theta^{*},x_{s}\rangle\right)\alpha(x_{s},\widehat{\theta}_{o},\theta^{*})x_{s}+\lambda\widehat{\theta}_{o}-\lambda_{t}\theta^{*}\right\rVert_{{\mathbf{G}}_{w}^{-1}}$
$=\sqrt{1+2RS}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle\widehat{\theta}_{o},x_{s}\rangle\right)-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}+\lambda\widehat{\theta}_{o}-\lambda_{t}\theta^{*}\right\rVert_{{\mathbf{G}}_{w}^{-1}}$ (Taylor’s theorem)
$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle\widehat{\theta}_{o},x_{s}\rangle\right)-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}+\lambda\widehat{\theta}_{o}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$ (${\mathbf{G}}_{w}\succeq\frac{1}{1+2RS}{\mathbf{H}}^{*}_{w}$)
Since $\widehat{\theta}_{o}$ is the maximum likelihood estimate, by optimality condition, we have the following relation: $\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\widehat{\theta}_{o}\rangle\right)x_{s}+\lambda\widehat{\theta}_{o}=\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}$. Substituting this above, we get
$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+\lambda\left\lVert\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta_{s}x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+S\sqrt{\lambda}.$ (${\mathbf{H}}^{*}_{w}\succeq\lambda_{t}{\mathbf{I}}$, $\left\lVert\theta^{*}\right\rVert_{2}\leq S$)
where $\eta_{s}\coloneqq\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)$.
We will now apply Lemma B.1 $\eta_{s}$ scaled by $R$. First note that $\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta_{s}x_{s}\right\rVert_{({\mathbf{H}}^{*}_{w})^{-1}}=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\frac{\eta_{s}}{R}x_{s}\right\rVert_{(R^{2}{\mathbf{H}}^{*}_{w})^{-1}}$ which, in turn ensures that the noise variable is upper bounded by $1$. Applying B.1 we get
$\lVert\widehat{\theta}_{o}-\theta^{*}$ $\rVert_{{\mathbf{H}}^{*}_{w}}\leq S\sqrt{R^{2}\lambda}$
$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{w}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
Simplifying constants and setting $\sqrt{R^{2}\lambda}=\textbf{c}\sqrt{d\log(T)+\log(1/\delta)}$, we have $\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}\leq\textbf{c}RS\sqrt{d\log(T/\delta)+\log(1/\delta)}\leq\textbf{c}RS\sqrt{d\log{(T/\delta)}}$. ∎
Next, we obtain a bound on the number of warm-up rounds.
Lemma B.3.
Algorithm 2, during its entire execution, enters the enters the warmup segment at most $2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$ times.
Proof.
Via the warm-up condition (Line 6) $\lVert x\rVert_{\mathbf{V}^{-1}}^{2}>1/(R^{2}\kappa\gamma^{2})$. Let $\mathbf{V}_{m}$ be the sequence of $\mathbf{V}$ matrices (line 7 of 2) for $m\in{\mathcal{T}}_{o}$. That is, $\mathbf{V}_{1}=\lambda{\mathbf{I}},\mathbf{V}_{m}=\sum_{s\in[m-1]\setminus{\mathcal{T}}_{o}}x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$). Thus,
$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\geq\frac{\tau_{w}}{R^{2}\kappa\gamma^{2}}$ (19)
Furthermore, by the Elliptic Potential Lemma (Lemma B.11) we have
$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\leq 2d\log\left(1+\frac{\tau_{w}}{\lambda d}\right)$ (20)
Combining (20) and (19) we have
$\tau_{w}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{w}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
We define ${\mathcal{E}}_{w}$ be the event define in Lemma B.2, that is, ${\mathcal{E}}_{w}=\{\lVert\widehat{\theta}_{w}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma_{t}\}$.
Lemma B.4.
If in round $t$ the warmup criteria is not satisfied and event ${\mathcal{E}}_{w}$, holds, we have
$\lvert\langle x,\widehat{\theta}_{o}-\theta^{*}\rangle\rvert\leq\frac{1}{R}\quad\text{ for all }x\in{\mathcal{X}}_{t}.$
Proof.
$\lvert\langle x,\widehat{\theta}_{o}-\theta^{*}\rangle\rvert$ $\leq\lVert x\rVert_{{\mathbf{H}}^{*-1}_{w}}\cdot\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ (Cauchy-Schwarz)
$\leq\lVert x\rVert_{{\mathbf{H}}^{*-1}_{w}}\gamma$ (via Lemma B.2)
$\leq\gamma\sqrt{\kappa}\lVert x\rVert_{\mathbf{V}_{w}^{-1}}$ ($\mathbf{V}_{w}\preceq\kappa{\mathbf{H}}^{*}_{w}$)
$\leq\gamma\sqrt{\kappa}\frac{1}{\sqrt{R^{2}\kappa\gamma^{2}}}$ (warm-up criteria is not satisfied)
$\leq\frac{1}{R}$
Recall that ${\mathbf{H}}_{t}$ is defined in line 15 of Algorithm 2. Further, we define
${\mathbf{H}}^{*}_{t}=\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$
Corollary B.5.
Under event ${\mathcal{E}}_{w}$, ${\mathbf{H}}_{t}\preceq{\mathbf{H}}_{t}^{*}\preceq e^{2}{\mathbf{H}}_{t}$
Proof.
For a given $s\in[t-1]\setminus{\mathcal{T}}_{o}$, let $\widehat{\theta}_{o}^{s}$ denote the value of $\widehat{\theta}_{o}$ in that round. Then, for all $x\in{\mathcal{X}}_{s}$, by Lemma C.2,
$\dot{\mu}\left(\langle x,\widehat{\theta}_{o}^{s}\rangle\right)\exp(-R\lvert\langle x,\widehat{\theta}_{o}^{s}-\theta^{*}\rangle\rvert)\leq\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\leq\dot{\mu}\left(\langle x,\widehat{\theta}_{o}^{s}\rangle\right)\exp(R\lvert\langle x,\widehat{\theta}_{o}^{s}-\theta^{*}\rangle\rvert)$
Applying lemma B.4, gives $e^{-1}\dot{\mu}\left(\langle x,\widehat{\theta}_{o}^{s}\rangle\right)\leq\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\leq e^{1}\dot{\mu}\left(\langle x,\widehat{\theta}_{o}^{s}\rangle\right)$. Thus,
${\mathbf{H}}_{t}=\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}e^{-1}\dot{\mu}\left(\langle x_{s},\widehat{\theta}_{o}^{s}\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}\preceq\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}={\mathbf{H}}_{t}^{*}$
. Further, ${\mathbf{H}}_{t}^{*}\preceq\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}e^{2}\frac{\dot{\mu}\left(\langle x_{s},\widehat{\theta}_{o}^{s}\rangle\right)}{e}x_{s}x_{s}^{\intercal}+e^{2}\lambda{\mathbf{I}}=e^{2}{\mathbf{H}}_{t}$. ∎
Now we define the following quantities:
$g_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda_{\tau}\theta$
${{\mathbf{H}}}_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$
$\Theta$ =$=\left\{\theta|\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}\right\}$
$\tilde{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})$
$\beta$ $\coloneqq\textbf{c}\sqrt{d\log(T/\delta)}$
Moreover, recall the following definition $\widehat{\theta}_{\tau}$:
$\widehat{\theta}_{\tau}=\operatorname*{arg\,min}_{\theta\in\Theta}\left\lVert g_{\tau}(\theta)-g_{\tau}(\tilde{\theta})\right\rVert_{{\mathbf{H}}_{\tau}(\theta)^{-1}}$ (21)
Lemma B.6.
Under event ${\mathcal{E}}_{w}$, $\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{\mathbf{V}}\leq 2\gamma\sqrt{\kappa}$
Proof.
First, we observe from Lemma B.2 that $\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma$. Using $\mathbf{V}\preceq\kappa{\mathbf{H}}^{*}_{w}$, we can write $\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}$. This implies that $\theta^{*}\in\Theta$. Now, $\widehat{\theta}_{\tau}\in\Theta$ by virtue of being a feasible solution to the optimization in (21). Thus,
$\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{\mathbf{V}}$ $=\left\lVert\widehat{\theta}_{\tau}-\widehat{\theta}_{o}+\widehat{\theta}_{o}-\theta^{*}\right\rVert_{\mathbf{V}}$
$\leq\left\lVert\widehat{\theta}_{\tau}-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}+\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{\mathbf{V}}$ (triangle inequality)
$\leq 2\gamma\sqrt{\kappa}$
Lemma B.7.
Let $\delta\in(0,1)$. Then, under event ${\mathcal{E}}_{w}$, with probability $1-\delta$, $\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta\coloneqq\textbf{c}\sqrt{d\log(T/\delta)}$.
Proof.
We have for all main rounds $s\in[\tau-1]\setminus{\mathcal{T}}_{o}$,
$\lvert\langle x_{s},\theta^{*}-\widehat{\theta}_{\tau}\rangle\rvert$ $\leq\left\lVert x_{s}\right\rVert_{\mathbf{V}^{-1}}\left\lVert\theta^{*}-\widehat{\theta}_{\tau}\right\rVert_{\mathbf{V}}$ (Cauchy-Schwarz)
$\leq\left\lVert x_{s}\right\rVert_{\mathbf{V}^{-1}}2\gamma\sqrt{\kappa}$ (by Lemma B.6)
$\leq 2\gamma\sqrt{\kappa}\frac{1}{R\gamma\sqrt{\kappa}}$ (warm up criterion not satisfied)
$=\frac{2}{R}$
Also note that $\theta^{*}\in\Theta$. Hence,
$\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}$ $\leq 2(1+2)\left\lVert g_{\tau}(\theta^{*})-\sum_{s\in[\tau-1]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}$ (by Lemma E.1)
$\leq 6\textbf{c}\sqrt{d\log(T/\delta)}$ (by Lemma B.1)
Let the event in lemma B.7 be denoted by ${\mathcal{E}}_{\tau}$, or in other words, ${\mathcal{E}}_{\tau}=\{\lVert\widehat{\theta}_{\tau}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta\}$.
Corollary B.8.
If round $t$ is a main round for Algorithm 2, then, under event ${\mathcal{E}}_{\tau}$, $\lvert\langle x,\widehat{\theta}_{\tau}-\theta^{*}\rangle\rvert\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{-1}}$
Proof.
$\lvert\langle x,\widehat{\theta}_{\tau}-\theta^{*}\rangle\rvert$ $\leq\lVert x\rVert_{{\mathbf{H}}_{\tau}^{*-1}}\cdot\lVert\widehat{\theta}_{\tau}-\theta^{*}\rVert_{{\mathbf{H}}_{\tau}^{*}}$ (Cauchy-Schwartz)
$\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{*-1}}$ (by Lemma B.7)
$\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{-1}}$ (${\mathbf{H}}_{\tau}\preceq{\mathbf{H}}_{\tau}^{*}$)
Lemma B.9.
For main rounds in Algorithm 2, let $x_{t}$ be the played arm defined via line 14 Algorithm 2. Then, under event ${\mathcal{E}}_{\tau}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq 2\sqrt{2}\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}~{}.$
Proof.
$\langle x^{*}_{t},\theta^{*}\rangle-\langle x_{t},\theta^{*}\rangle$ $\leq\left(\langle x^{*}_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x^{*}_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}\right)$ (by Corollary B.8)
$\leq\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)$ (optimistic $x_{t}$, see line 14 algo. 2)
$=2\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}$
$\leq 2\sqrt{2}\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}$ (Lemma B.12, $\frac{\det({\mathbf{H}}_{\tau}^{-1})}{\det({\mathbf{H}}_{t}^{-1})}=\frac{\det({\mathbf{H}}_{t})}{\det({\mathbf{H}}_{\tau}})\leq 2$)
Lemma B.10.
For non-warm up rounds, the arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$ (b) for $x_{t},x^{*}_{t}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
Proof.
Suppose $x^{\prime}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}LCB_{w}(x)$. Now, we have, for all $x\in{\mathcal{X}}_{t}$,
$\lvert\langle x,\widehat{\theta}_{o}-\theta^{*}\rangle\rvert$ $\leq\left\lVert x\right\rVert_{{\mathbf{H}}^{*-1}_{w}}\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}$ (Cauchy-Schwarz)
$\leq\gamma\left\lVert x\right\rVert_{{\mathbf{H}}^{*-1}_{w}}$ (by Lemma B.2)
$\leq\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ ($\mathbf{V}\preceq\kappa{\mathbf{H}}^{*}_{w}$)
Thus, for $x^{*}_{t}$, we can write,
$\langle x^{*}_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{*}_{t},\theta^{*}\rangle$
$\geq\langle x^{\prime},\theta^{*}\rangle$ (optimality of $x^{*}_{t}$)
$\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
Thus, $x^{*}_{t}\in{\mathcal{X}}^{\prime}_{t}$.
Since $x_{t}$ is also in ${\mathcal{X}}^{\prime}_{t}$,
$\langle x_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
$\geq\langle x_{t}^{*},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ ($x^{\prime}$ has max LCB)
Again, using the fact that $\langle x^{*}_{t},\widehat{\theta}_{o}\rangle\geq\langle x^{*}_{t},\theta^{*}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ and $\langle x_{t},\widehat{\theta}_{o}\rangle\leq\langle x_{t},\theta^{*}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$, we obtain,
$\langle x_{t},\theta^{*}\rangle+2\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{*}_{t},\theta^{*}\rangle-2\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$
$\implies\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ $\leq 2\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}+2\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$
Finally, since this is a non warm-up round, $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}<\frac{1}{R\gamma\sqrt{\kappa}}$ for all $x\in{\mathcal{X}}_{t}$. Plugging this above,
$\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
B.1 Proof of Theorem 4.2
We prove Theoreom 4.2 in this section
Proof.
Firstly, we will assume throughout the proof that ${\mathcal{E}}_{w}\cap{\mathcal{E}}_{\tau}$ holds, which happens with probability at least $1-\delta$. Thus, regret of Algorithm 2 is upper bounded as:
$\textsc{R}_{T}$ $=\sum_{t\in[T]}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$
$\leq R\tau_{w}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$ (Upper bound of $R$ for warm-up rounds)
$\leq\textbf{c}R^{3}\kappa\gamma^{2}\log(T/\delta)+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ (some $z\in[\langle x_{t},\theta^{*}\rangle,\langle x_{t}^{*},\theta^{*}\rangle]$; lemma B.3)
Now, let $R_{1}(T)=\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$. Hereon, we will slightly abuse notation ${\mathbf{H}}_{\tau}$ to denote the ${\mathbf{H}}_{\tau}$ matrix last updated before time $t$ for each time step $t\in[T]$. This will be clear from the context as we will only use ${\mathbf{H}}_{\tau}$ term-wise. With this, we upper bound $R_{1}(T)$ as follows:
$R_{1}(T)$ $\leq\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)2\beta\lVert x_{t}\rVert_{{\mathbf{H}}_{\tau}^{-1}}$ (by Lemma B.9)
$\leq\sqrt{2}\beta\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)2\lVert x_{t}\rVert_{{\mathbf{H}}_{t}^{-1}}$ (Lemma B.12, $\frac{\det({\mathbf{H}}_{\tau}^{-1})}{\det({\mathbf{H}}_{t}^{-1})}=\frac{\det({\mathbf{H}}_{t})}{\det({\mathbf{H}}_{\tau}})\leq 2$)
$\leq 2\sqrt{2}\beta\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)e^{1}\lVert x_{t}\rVert_{{\mathbf{H}}_{t}^{*-1}}$ (by Lemma B.5)
$\leq 2e\sqrt{2}\beta\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\sqrt{\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\dot{\mu}\left(\langle x_{t},\theta^{*}\rangle\right)}\exp(R\langle x_{t}^{*}-x_{t},\theta^{*}\rangle)\lVert x_{t}\rVert_{{\mathbf{H}}_{t}^{*-1}}$ (by Lemma C.2)
$\leq 2e\sqrt{2}\beta\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\sqrt{\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}\sqrt{\dot{\mu}\left(\langle x_{t},\theta^{*}\rangle\right)}e^{4}\left\lVert x_{t}\right\rVert_{{\mathbf{H}}_{t}^{*-1}}$ (by Lemma B.10)
$=2e^{5}\sqrt{2}\beta\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\sqrt{\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}\sqrt{\dot{\mu}\left(\langle x_{t},\theta^{*}\rangle\right)}\left\lVert x_{t}\right\rVert_{{\mathbf{H}}_{t}^{*-1}}$
$=2e^{5}\sqrt{2}\beta\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\sqrt{\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}\left\lVert\tilde{x}_{t}\right\rVert_{{\mathbf{H}}_{t}^{*-1}}$ ($\tilde{x}_{t}=\sqrt{\dot{\mu}\left(\langle x_{t},\theta^{*}\rangle\right)}x_{t}$)
$\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\left\lVert\tilde{x}_{t}\right\rVert_{{\mathbf{H}}_{t}^{*-1}}^{2}}$ (Cauchy-Schwarz)
$\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot 2d\log\left(1+\frac{RT}{\lambda d}\right)}$ (Elliptic Potential Lemma; $\left\lVert\tilde{x}_{t}\right\rVert_{2}\leq R$)
$\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}.$
Putting things back,
$\textsc{R}_{T}\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}+\textbf{c}R^{5}S^{2}\kappa\log(T/\delta)^{2}.$
B.2 Proof of Lemma 4.1
We first restate Lemma 4.1.
Lemma (4.1).
Algorithm 2, during its entire execution, updates its policy at most $O(R^{4}S^{2}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
Proof.
Note that in Algorithm 2, policy changes happen only in the warm-up rounds and when the determinant of ${\mathbf{H}}_{t}$ doubles in the main round. Thus, total number of policy switches is upper bounded by the number of warm-up rounds and the number of times $\det({\mathbf{H}}_{t})$ doubles. The first quantity is bounded by Lemma B.3 while the second quantity is bounded by Lemma B.14. Thus in total, the number of policy changes in Algorithm A is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
B.3 Some Useful Lemmas
Lemma B.11 (Elliptic Potential Lemma (Lemma 10 Abbasi-Yadkori et al. [2011])).
Let $x_{1},x_{2},\ldots x_{t}$ be a sequence of vectors in $\mathbb{R}^{d}$ and let $\left\lVert x_{s}\right\rVert_{2}\leq L$ for all $s\in[t]$. Further, let $\mathbf{V}_{s}=\sum_{m=1}^{s-1}x_{m}x_{m}^{\intercal}+\lambda{\mathbf{I}}$. Suppose $\lambda\geq L^{2}$. Then,
$\sum_{s=1}^{t}\left\lVert x_{s}\right\rVert_{\mathbf{V}_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{L^{2}t}{\lambda d}\right)$ (22)
Lemma B.12 (Lemma 12 of Abbasi-Yadkori et al. [2011]).
Let $A\succeq B\succ 0$. Then
$\sup_{x\neq 0}\frac{x^{\intercal}Ax}{x^{\intercal}Bx}\leq\frac{\det(A)}{\det(B)}$
Lemma B.13 (Lemma 10 of Abbasi-Yadkori et al. [2011]).
Let $\{x_{s}\}_{s=1}^{t}$ be a set of vectors. Define the sequence $\{\mathbf{V}_{s}\}_{s=1}^{t}$ as $\mathbf{V}_{1}=\lambda{\mathbf{I}}$, $\mathbf{V}_{s+1}=\mathbf{V}_{s}+x_{s}x_{s}^{\intercal}$ for $s\in[t-1]$. Further, let $\left\lVert x_{s}\right\rVert_{2}\leq L\ \forall\ s\in[t]$. Then,
$\det(\mathbf{V}_{t})\leq\left(\lambda+tL^{2}/d\right)^{d}~{}.$
Lemma B.14.
Let $\{x_{s}\}_{s=1}^{t}$ be a set of vectors. Define the sequence $\{\mathbf{V}_{s}\}_{s=1}^{t}$ as $\mathbf{V}_{1}=\lambda{\mathbf{I}}$, $\mathbf{V}_{s+1}=\mathbf{V}_{s}+x_{s}x_{s}^{\intercal}$ for $s\in[t-1]$. Further, let $\left\lVert x_{s}\right\rVert_{2}\leq L\ \forall\ s\in[t]$. Define the set $\{1=\tau_{1},\tau_{2}\ldots\tau_{m}=t\}$ such that: $\det(\mathbf{V}_{\tau_{i+1}})\geq 2\det(\mathbf{V}_{\tau_{i}})$ but $\det(\mathbf{V}_{\tau_{i+1}-1})<2\det(\mathbf{V}_{\tau_{i}})$ for $i\in\{2,\ldots m-1\}$. Then, the number of time doubling happens,i.e.$m$ is at most $O(d\log(t))$.
Proof.
By Lemma B.13, $\det(\mathbf{V}_{t})\leq\left(\lambda+tL^{2}/d\right)^{d}$. But we have that from definition of $\tau_{i}$’s
$\det(\mathbf{V}_{t})$ $\geq\det(\mathbf{V}_{\tau_{m-1}})$
$\geq 2\det(\mathbf{V}_{\tau_{m-2}})$
$\vdots$
$\geq 2^{m-2}\det(\mathbf{V}_{\tau_{1}})$
$=2^{m-2}\det(\mathbf{V}_{1})$
$=2^{m-2}\lambda^{d}$ ($\mathbf{V}_{1}=\lambda{\mathbf{I}}$)
Thus, $2^{m-2}\lambda^{d}\leq\left(\lambda+tL^{2}/d\right)^{d}$ which implies that
$2^{m-2}$ $\leq\left(1+\frac{tL^{2}}{\lambda d}\right)^{d}$
Hence, $m\leq O(d\log(t))~{}.$
Appendix C Useful Properties of GLMs
Recall that a Generalized Linear Model is characterized by a canonical exponential family, i.e., the random variable $r$ has density function $p_{z}\left(r\right)=\exp\left(rz-b\left(z\right)+c\left(r\right)\right)$, with parameter $z$, log-partition function $b(\cdot)$, and a function $c$. Further, $\dot{b}(z)=\mu(z)$ is also called the link function.
Hereon, we will assume that the random variable has a bounded non-negative support, i.e., $r\in[0,R]$ almost surely. Now, we state the following key Lemmas on GLMs
Lemma C.1 (Self-Concordance for GLMs).
For distributions in the exponential family the function $\mu(\cdot)$ satisfies that for all $z\in\mathbb{R}$, $\lvert\ddot{\mu}(z)\rvert\leq R\dot{\mu}(z)$.
Proof.
Indeed,
$\lvert\dddot{b}(z)\rvert$ $=\lvert\operatorname*{\mathbb{E}}[(r-\operatorname*{\mathbb{E}}[r])^{3}]\rvert$ (Lemma C.3)
$\leq\operatorname*{\mathbb{E}}\left[\lvert(r-\operatorname*{\mathbb{E}}[r])^{3}\rvert\right]$ (Jensen’s inequality)
$=\operatorname*{\mathbb{E}}\left[\lvert r-\operatorname*{\mathbb{E}}[r]\rvert\cdot(r-\operatorname*{\mathbb{E}}[r])^{2}\right]$
$\leq\operatorname*{\mathbb{E}}[R(r-\operatorname*{\mathbb{E}}[r])^{2}]$ ($r,\operatorname*{\mathbb{E}}[r]\in[0,R]$)
$=R\operatorname*{\mathbb{E}}[(r-\operatorname*{\mathbb{E}}[r])^{2}]$
$=R\ddot{b}(z)$ (Lemma C.3)
As a consequence, we have the following simple modification of the self-concordance results of Faury et al. [2020].
Lemma C.2.
For an exponential distribution with log-partition function $b(\cdot)$, for all $z_{1},z_{2}\in\mathbb{R}$, letting $\mu(z)\coloneqq\dot{b}(z)$, following holds:
$\alpha(z_{1},z_{2})\coloneqq\int_{v=0}^{1}\dot{\mu}\left(z_{1}+v\left(z_{2}-z_{1}\right)\right)$ $\geq\frac{\dot{\mu}\left(z\right)}{1+R|z_{1}-z_{2}|}\quad\text{for $z\in\{z_{1},z_{2}\}$}$ (23)
$\frac{\dot{\mu}(z_{2})}{e^{R|z_{2}-z_{1}|}}$ $\leq\dot{\mu}\left(z_{1}\right)\leq e^{R|z_{2}-z_{1}|}\dot{\mu}\left(z_{2}\right)$ (24)
[…]
Proof.
[…]
$-R\dot{\mu}(v)\leq$ $\ddot{\mu}(v)\leq R\dot{\mu}(v)$ (Lemma C.1)
$-R\leq$ $\frac{\ddot{\mu}(v)}{\dot{\mu}(v)}\leq R$
[…]
$-R(z-z_{1})\leq$ $\log\left(\frac{\dot{\mu}(z)}{\dot{\mu}(z_{1})}\right)\leq R(z-z_{1})$
$\dot{\mu}(z_{1})\exp(-R(z-z_{1}))\leq$ $\dot{\mu}(z)\leq\dot{\mu}(z_{1})\exp(R(z-z_{1}))$
Putting $z=z_{2}$ establishes 24. To show 23, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
[…]
which $\text{which gives}\quad\dot{\mu}(z_{1})\frac{1-\exp(-R(z_{2}-z_{1}))}{R(z_{2}-z_{1})}\leq$ $\alpha(z_{1},z_{2})\leq\dot{\mu}(z_{1})\frac{\exp(R(z_{2}-z_{1}))-1}{R(z_{2}-z_{1})}$
Next, we use the fact that for $x>0$, $e^{-x}\leq(1+x)^{-1}$ which on rearranging gives $(1-e^{-x})/x\geq 1/(1+x)$. Applying this inequality to the LHS above finishes the proof. Note that similar exercise can be repeated with $z_{2}\leq z_{1}$ to get the same result for $z_{2}$.
For 25, we have, by application of 24, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
[…]
$=\dot{\mu}(z_{1})\left(\frac{1}{R|z_{1}-z_{2}|}+\frac{\exp(-R|z_{1}-z_{2}|)-1}{R^{2}|z_{1}-z_{2}|^{2}}\right)$
$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of [Abeille et al., 2021])
Next we state some nice properties of the GLM family that is the key in deriving Lemma C.1.
Lemma C.3 (Properties of GLMs).
For any random variable $r$ that is distributed by a canonical exponential family, we have
  1. 1.
    $\operatorname*{\mathbb{E}}\left[r\right]=\mu\left(z\right)=\dot{b}\left(z\right)$
  2. 2.
    $\mathbb{V}[r]=\operatorname*{\mathbb{E}}\left[\left(r-\operatorname*{\mathbb{E}}\left[r\right]\right)^{2}\right]=\dot{\mu}\left(z\right)=\ddot{b}\left(z\right)$
  3. 3.
    $\operatorname*{\mathbb{E}}\left[\left(r-\operatorname*{\mathbb{E}}[r]\right)^{3}\right]=\dddot{b}(z)$
Proof.
  1. 1.
[…]
  2. 2.
[…]
    $\ddot{b}(z)$ $=-\dot{b}(z)\exp(-b(z))f(z)+\exp(-b(z))\dot{f}(z)$
    $=-\operatorname*{\mathbb{E}}[r]^{2}+\exp(-b(z))\int_{r}r^{2}\exp(rz+c(r))dr$
    $=-\operatorname*{\mathbb{E}}[r]^{2}+\int_{r}r^{2}\exp(rz-b(z)+c(r))dr$
    $=-\operatorname*{\mathbb{E}}[r]^{2}+\operatorname*{\mathbb{E}}[r^{2}]=\mathbb{V}[r]$
  3. 3.
[…]
    $\dddot{b}(z)$ $=-2\dot{b}(z)\ddot{b}(z)-\dot{b}(z)\exp(-b(z))f(z)+\exp(-b(z))\dot{f}(z)$
[…]
    $=-2\operatorname*{\mathbb{E}}[r]\mathbb{V}[r]-\operatorname*{\mathbb{E}}[r]\operatorname*{\mathbb{E}}[r^{2}]+\operatorname*{\mathbb{E}}[r^{3}]$
    Now, let us expand $\operatorname*{\mathbb{E}}[(r-\operatorname*{\mathbb{E}}[r])^{3}]$.
    $\operatorname*{\mathbb{E}}[(r-\operatorname*{\mathbb{E}}[r])^{3}]$ $=\operatorname*{\mathbb{E}}[r^{3}-3r^{2}\operatorname*{\mathbb{E}}[r]+3r\operatorname*{\mathbb{E}}[r]^{2}-\operatorname*{\mathbb{E}}[r]^{3}]$
    $=\operatorname*{\mathbb{E}}[r^{3}]-3\operatorname*{\mathbb{E}}[r]\operatorname*{\mathbb{E}}[r^{2}]+3\operatorname*{\mathbb{E}}[r]^{3}-\operatorname*{\mathbb{E}}[r]^{3}$
    $=\operatorname*{\mathbb{E}}[r^{3}]-\operatorname*{\mathbb{E}}[r]\operatorname*{\mathbb{E}}[r^{2}]-2\operatorname*{\mathbb{E}}[r]\left(-\operatorname*{\mathbb{E}}[r^{2}]+\operatorname*{\mathbb{E}}[r]^{2}\right)$
    $=\operatorname*{\mathbb{E}}[r^{3}]-\operatorname*{\mathbb{E}}[r]\operatorname*{\mathbb{E}}[r^{2}]-2\operatorname*{\mathbb{E}}[r]\mathbb{V}[r]$
Corollary C.4.
For all exponential family, $b(\cdot)$ is a convex function.
Proof.
Indeed, note that $\ddot{b}(z)=\mathbb{V}[r]$ which is always non-negative. Thus, $\ddot{b}(z)\geq 0$ implying that $b(\cdot)$ is convex. ∎
Remark C.5.
In [Faury, 2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set. Thereafter the author notes that the techniques developed in [Faury, 2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter). However, the claim regarding self-concordance of GLMs is not true in general. There are classes of GLMs whose parameters may be restricted in a bounded set, but for them no constant $a$ exists. One such example is the exponential distribution. The link function $\mu$ for exponential distribution is given as $\mu(z)=-\frac{1}{z}$. If we allow $z$ to lie in the set $(-c,0)$ for some positive $c$, then we have $\mu(z)$ strictly increasing (satisfying our assumption on monotonicity of $\mu$, thus a valid example). However, for this GLM,
$\dot{\mu}(z)=\frac{1}{z^{2}}\quad\quad\ddot{\mu}(z)=-\frac{2}{z^{3}}$
Note that $\ddot{\mu}(z)$ is positive for the assumed support of $z$. Suppose this GLM is self-concordant, then we must have some positive constant $a$ such that
$|\ddot{\mu}(z)|=-\frac{2}{z^{3}}\leq a\dot{\mu}(z)=a\frac{1}{z^{2}}~{}.$
Simplifying, we obtain the following relation:
$-\frac{2}{z}\leq a~{}.$
However, since $z\in(-c,0)$, we have $\lim_{z\to 0}-\frac{2}{z}\to\infty$. Hence, no constant $a$ is possible. By this counterexample it can be seen that bounded parameter set is not enough to guarantee self-concordance of GLMs. In this work, we give a characterization of self-concordance of GLMs with bounded support of the random variable. It will be interesting to understand a complete characterization of self-concordance of GLMs.
Appendix D Computational Cost
Consider a log-loss minimization oracle that returns the unconstrained MLE for a given GLM class with a computational complexity of $C_{opt}\cdot n$, when the log-loss is computed over $n$ data points. Let the maximum number of arms available every round be $K$. Furher, let the computational cost of an oracle that solves the non-convex optimization 6 be $NC_{opt}$.
Computational Cost of B-GLinCB: In the B-GLinCB algorithm, we employ the log-loss oracle at the end of each batch. The estimator $\widehat{\theta}$ calculated at the end of a batch of length $\tau$ incurs a computational cost of $C_{opt}\tau$. Furthermore, this oracle is invoked for a maximum of $M\leq\log{\log{T}}$ batches. Additionally, the computation of the distributional optimal design at the end of each batch is efficient in $d$ ($\text{poly}(d)$). Moreover, in every round, the algorithm solves the $D/G-$ Optimal Design problem (requiring $O(d\log{d})$ computation) and runs elimination based on prior (at most $\log{\log T}$) phases. Hence, the amortized cost per round of B-GLinCB is $O(K\log{\log T}+d\log{d}+C_{opt})$.
Computational Cost of RS-GLinCB: In the RS-GLinCB algorithm, the estimator $\widehat{\theta}_{w}$ is computed during each warmup round. Additionally, during non-warmup rounds, the estimator $\widehat{\theta}$ is computed a maximum of $O(\log(T))$ times. These computations involve utilizing both the log-loss oracle and the non-convex projection oracle. Furthermore, in each non-warmup round, the algorithm executes an elimination step. This yields an amortized time complexity of $O(C_{opt}\log{T}+NC_{opt}\log^{2}(T)+K)$ per round.
Performance in Practice: As evident from Fig. 1, RS-GLinCB has much better computational performance in practice. We ran all the experiments on an Azure Data Science VM equipped with AMD EPYC 7V13 64-Core Processor (clock speed of 2.45 GHz) and Linux Ubuntu 20.04 LTS operating system. It was ensured that no other application processes were running while we tested the performance. We implemented and tested our code in Python, and measured the execution times using time.time() command. We allowed no operations for 10 seconds after every run to let the CPU temperature come back to normal, in case the execution heats up the CPU, thereby causing subsequent runs to slow down.
Comparison with ECOLog [Faury et al., 2022] shows that execution time for RS-GLinCB is significantly smaller. We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round. On an implementation level, this translates into more function calls and computation. Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
While comparing with GLOC [Jun et al., 2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes. Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$. On the other hand, in accordance with Lemma 4.1, the computational cost of RS-GLinCB increases with $\kappa$. However, after a few initial rounds, when neither of the switching criteria are triggered, RS-GLinCB does not need to solve any computationally intensive optimization problem, hence these rounds execute very fast. In practice, with typical data distribution, RS-GLinCB reaches this stage much before what the worst-case guarantees show, hence we see it perform better than GLOC.
Appendix E Projection
We describe the projection step used in Algorithms 1 and 2. We present arguments similar to the ones made in Appendix B.3 of Faury et al. [2020]. We write
${\mathbf{H}}(\theta)=\sum_{s=1}^{t}\dot{\mu}\left(\langle\theta,x_{s}\rangle\right)x_{s}x_{s}^{\mathsf{T}}+\lambda{\mathbf{I}}$
Recall, ${\mathbf{H}}^{*}={\mathbf{H}}(\theta^{*})$. Let $\widehat{\theta}$ be the MLE estimator of $\theta^{*}$ calculated after the sequence arm pulls $x_{1},x_{2},\ldots,x_{t}$. Let $r_{1},r_{2},\ldots,r_{t}$ be the corresponding observed rewards. We project $\widehat{\theta}$ to a set $\Theta$ by solving the following optimization problem
$\widetilde{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widehat{\theta}\rangle\right))x_{s}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}$ (26)
Lemma E.1.
Using the notations described above, if $\theta^{*}\in\Theta$ and $\max_{i\in[t]}|\langle x_{i},\widetilde{\theta}-\theta^{*}\rangle|\leq c/R$, then we have
$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}(\theta^{*})}\leq 2(1+c)\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta^{*}\rangle\right)-r_{s})x_{s}\right\rVert_{{\mathbf{H}}(\theta^{*})^{-1}}$
Proof.
First, we note that by self-concordance property of $\mu$ (lemma C.2), for any $s\in[t]$,
$\alpha(x_{s},\widetilde{\theta},\theta^{*})$ $\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{1+R\lvert\langle x_{s},\widetilde{\theta}-\theta^{*}\rangle\rvert}$
$\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{1+R(c/R)}$ ($\max_{i\in[s]}|\langle x_{s},\widetilde{\theta}-\theta^{*}\rangle|\leq c/R$)
$=\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{1+c}$
Similarly, we have $\alpha(x_{s},\widetilde{\theta},\theta^{*})\geq\frac{\dot{\mu}\left(\langle x_{s},\widetilde{\theta}\rangle\right)}{1+c}$.
Let us define the matrix ${\mathbf{G}}=\sum_{s\in[t]}\alpha(x,\widetilde{\theta},\theta^{*})x_{s}x_{s}^{\intercal}$. Using the above fact, we obtain the relation: ${\mathbf{G}}\succeq\frac{1}{1+c}{\mathbf{H}}^{*}$ and ${\mathbf{G}}\succeq\frac{1}{1+c}{{\mathbf{H}}}(\widetilde{\theta})$. Also define the vector $g(\theta)=\sum_{s\in[t]}\mu\left(\langle\theta,x_{s}\rangle\right)x_{s}$. Now,
$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}$ $\leq\sqrt{1+c}\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{G}}}$ (${\mathbf{H}}^{*}\preceq(\sqrt{1+c}){\mathbf{G}}$)
$=\sqrt{1+c}\left\lVert{\mathbf{G}}\left(\widetilde{\theta}-\theta^{*}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\alpha(x_{s},\widetilde{\theta},\theta^{*})\langle\widetilde{\theta}-\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}$
$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\mu\left(\langle x_{s},\widetilde{\theta}\rangle\right)-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}$ (Taylor’s theorem)
$=\sqrt{1+c}\left\lVert\left(\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}\right)-\left(\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
Let $g(\theta)=\sum_{s=1}^{t}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}$ for any $\theta$. Therefore, we have,
$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}$ $\leq\sqrt{1+c}\left\lVert g(\widetilde{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
$=\sqrt{1+c}\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})+g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
$\leq\sqrt{1+c}\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}^{-1}}+\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}\right)$ ($\triangle$ inequality)
$\leq(1+c)\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})\right\rVert_{{{\mathbf{H}}(\widetilde{\theta})}^{-1}}+\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}\right)$ (${\mathbf{H}}^{*-1}\succeq(\sqrt{1+c}){\mathbf{G}}^{-1}$)
$\leq 2(1+c)\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}$ (by (26))
$=2(1+c)\left\lVert g(\theta^{*})-\sum_{s\in[t]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}$ ($\widehat{\theta}$ is the unconstrained MLE, $g(\widehat{\theta})=\sum_{s\in[t]}r_{s}x_{s}$.)
E.1 Convex Relaxation
The optimization problem in (26) is a non-convex optimization problem and therefore it is not clear what is the computational complexity of the problem. However, it is possible to substitute this optimization problem with a convex one, whose computational complexity can be better tractable. The process is similar to the one detailed in [Abeille et al., 2021, section 6]. Here we briefly outline the steps.
Let ${\mathcal{L}}_{t}(\theta)=\sum_{s=1}^{t}\ell(\theta,x_{s},r_{s})$ and $\breve{\theta}$ be defined as follows:
$\breve{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}{\mathcal{L}}_{t}(\theta)$ (27)
Note that when the set $\Theta$ is a convex set, then the above optimization problem is convex by property of the log-likelihood function of GLMs. Hence it can be solved efficiently. With this projected $\breve{\theta}$, we have the following guarantee:
Lemma E.2.
Suppose $\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}\leq\gamma$ and $\lambda=\gamma/R$. If $\theta^{*}\in\Theta$ and $\max_{i\in[t]}|\langle x_{i},\breve{\theta}-\theta^{*}\rangle|\leq c/R$, then we have
$\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}(\theta^{*})}\leq\textbf{c}\sqrt{(2+c)R^{3}S}\gamma$
Proof.
First we note that by self-concordance property of $\mu$, for any $s\in[t]$,
$\tilde{\alpha}(x_{s},\theta^{*},\breve{\theta})$ $\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R|\langle x_{s},\breve{\theta}-\theta^{*}\rangle|}$ (Lemma C.2)
$\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R(c/R)}$ ($\max_{i\in[s]}|\langle x_{s},\widetilde{\theta}-\theta^{*}\rangle|\leq c/R$)
$=\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+c}$
Let us define $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\coloneqq\sum_{s=1}^{t}\tilde{\alpha}(x_{s},\theta^{*},\theta)x_{s}x_{s}^{\intercal}$. Using the above fact, we obtain $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\succeq\frac{1}{2+c}{\mathbf{H}}^{*}$.
We now follow closely the proof outlined in Appendix B.3 of [Abeille et al., 2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
${\mathcal{L}}_{t}(\theta)-{\mathcal{L}}_{t}(\theta^{*})-\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\theta-\theta^{*}\rangle$ $=\left\lVert\theta-\theta^{*}\right\rVert_{\tilde{{\mathbf{G}}}(\theta,\theta^{*})}^{2}$
$\geq\frac{1}{2+c}\left\lVert\theta-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}$
Taking absolute value on both sides, and substituting $\theta=\breve{\theta}$,
$\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}$ $\leq(2+c)\left(|{\mathcal{L}}_{t}(\breve{\theta})-{\mathcal{L}}_{t}(\theta^{*})|+|\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\breve{\theta}-\theta^{*}\rangle|\right)$ ($\triangle$-inequality)
$\leq(2+c)\left(|{\mathcal{L}}_{t}(\breve{\theta})-{\mathcal{L}}_{t}(\theta^{*})|+\left\lVert\nabla{\mathcal{L}}_{t}(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}\right)$ (Cauchy-Schwarz)
$=(2+c)\left(|{\mathcal{L}}_{t}(\breve{\theta})-{\mathcal{L}}_{t}(\theta^{*})|+\left\lVert g(\theta^{*})-\sum_{s\in[t]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}\right)$
Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$. By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of [Abeille et al., 2021]), we have, for $\theta^{*}$.
${\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})$ $\leq\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}(\theta^{*},\widehat{\theta})^{-1}}^{2}$
$\leq\frac{R}{\sqrt{\lambda}}\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}^{2}+\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$
$\leq\frac{R}{\sqrt{\lambda}}\gamma^{2}+\gamma$ (Lemma B.1)
$\leq 2R^{3}S\gamma$ (recall $\sqrt{R^{2}\lambda}=\gamma/RS$)
We also have, by definition of $\breve{\theta}$, whenever $\theta^{*}\in\Theta$, ${\mathcal{L}}_{t}(\breve{\theta})\leq{\mathcal{L}}_{t}(\theta^{*})$, therefore we have ${\mathcal{L}}_{t}(\breve{\theta})-{\mathcal{L}}_{t}(\widehat{\theta})\leq{\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})\leq 2R^{3}S\gamma$ Thus, we have,
$\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}\leq(2+c)\left(4R^{3}S\gamma+\gamma\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}\right)$
Using the inequality that for some $x^{2}\leq bx+c\implies x\leq b+\sqrt{c}$, we have,
$\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}$ $\leq(2+c)\gamma+\sqrt{(2+c)4R^{3}S\gamma}$
$=\textbf{c}\sqrt{(2+c)R^{3}S}\gamma$

### v4

@@
 (2021)), the algorithm is allowed to adaptively decide (during its course) when to update its policy.
-Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting Ruan et al.
++–+++++++++++ Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting Ruan et al.
 (2021); Han et al.
@@
 (2021).
-Along with this, the rounds for policy updates are also chosen more carefully (in a $\kappa$ dependent fashion) leading to a stronger regret guarantee.
+Along with this, the rounds for policy updates are also chosen more carefully (in a $\kappa$ dependent fashion), leading to a stronger regret guarantee.
 $\bullet$ We propose RS-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M2 setting of limited adaptivity.
@@
 This new criterion allows us to prove optimal regret guarantee ($\tilde{O}(\sqrt{T})$) with only $O(\log^{2}T)$ updates to the policy.
-It is quite crucial for the generalized linear reward settings, since without it the resultant regret guarantees have a linear dependence on $\kappa$.
+It is quite crucial for the generalized linear reward settings since, without it, the resultant regret guarantees have a linear dependence on $\kappa$.
 $\bullet$ Our work also resolves a conjecture in Lee et al.
-(2023) by proving an optimal ($\tilde{O}(\sqrt{T})$) regret guarantee (for the CB problem with logistic reward model) that does not depend polynomially on $S$ (the known upper bound on size of the model parameters, i.e.
-$\|\theta^{\star}\|\leq S$, See Section 2).
-RS-GLinCB is, to our knowledge, the first CB algorithm for generalized linear reward models, that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret.
+(2024) by proving an optimal ($\tilde{O}(\sqrt{T})$) regret guarantee (for the CB problem with logistic reward model) that does not depend polynomially on $S$ (the known upper bound on the size of the model parameters, i.e.
+$\|\theta^{\star}\|\leq S$, See Section 2111This requires a non-convex projection.
+We discuss its convex relaxation in Appendix E..
+RS-GLinCB is, to our knowledge, the first CB algorithm for generalized linear reward models that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret.
 We also perform experiments in Section 5 that validate its superiority both in terms of regret and computational efficiency in comparison to other baseline algorithms proposed in Jun et al.
@@
 Specifically, our results imply that the CB problem with generalized linear reward models originally proposed in Filippi et al.
-(2010) and subsequently studied in literature (Li et al., 2017; Jun et al., 2017; Russac et al., 2020) admits a $\kappa$-independent regret.
+(2010) and subsequently studied in literature Li et al.
+(2017); Jun et al.
+(2017); Russac et al.
+(2020) admits a $\kappa$-independent regret.
 Remark 1.2 (Computational efficiency).
@@
 (2017) recently devised computationally efficient algorithms but they suffer from regret dependence on $\kappa$.
-Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models (Faury et al., 2020; Abeille et al., 2021), and the algorithms were subsequently made computationally efficient in Faury et al.
+Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models Faury et al.
+(2020); Abeille et al.
+(2021), and the algorithms were subsequently made computationally efficient in Faury et al.
 (2022); Zhang and Sugiyama (2023).
@@
 This result was postulated in Filippi et al.
-(2010) for GLMs (with same definition as ours) but no proof was provided.
+(2010) for GLMs (with the same definition as ours), but no proof was provided.
 While Faury et al.
-(2020; 2022) partially tackled this issue for logistic reward models [^2] In Faury (2021), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made, however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
+(2020; 2022) partially tackled this issue for logistic reward models [^3] In Faury (2021), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made; however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
 2 Notations and Preliminaries
@@
 Further, we assume that $r\in[0,R]$ almost surely for some known $R\in\mathbb{R}$.
-Important properties of GLMs such as $\operatorname*{\mathbb{E}}[r]=\dot{b}(\langle x,\theta^{\star}\rangle)$ and variance $\mathbb{V}[r]=\ddot{b}(\langle x,\theta^{\star}\rangle)$ are detailed in Appendix C.
-We define the link function $\mu$ as $\mu\left(\langle x,\theta^{*}\rangle\right)\coloneqq\operatorname*{\mathbb{E}}[r]$.
+Important properties of GLMs such as $\operatorname*{\mathbb{E}}[r\mid x]=\dot{b}(\langle x,\theta^{\star}\rangle)$ and variance $\mathbb{V}[r\mid x]=\ddot{b}(\langle x,\theta^{\star}\rangle)$ are detailed in Appendix C.
+We define the link function $\mu$ as $\mu\left(\langle x,\theta^{*}\rangle\right)\coloneqq\operatorname*{\mathbb{E}}[r\mid x]$.
 Thus, $\mu$ is also monotone.
-We now present a key Lemma on GLMs (see Appendix C for details) that enable us to achieve optimal regret guarantees for our algorithms designed in Sections 3 and 4.
+We now present a key Lemma on GLMs (see Appendix C for details) that enables us to achieve optimal regret guarantees for our algorithms designed in Sections 3 and 4.
 Lemma 2.2 (Self-Concordance of GLMs).
@@
 $\textsc{R}_{T}=\sum_{t=1}^{T}\max_{x\in{\mathcal{X}}_{t}}\mu\left(\langle x,\theta^{*}\rangle\right)\ -\ \sum_{t=1}^{T}\mu\left(\langle x_{t},\theta^{*}\rangle\right)$
+[…]
 2.1 Instance Dependent Non-Linearity Parameters
@@
 (2020; 2022), we define instance dependent parameters that capture non-linearity of the underlying instance and critically impact our algorithm design.
-Algorithm 1 (B-GLinCB) that solves Problem $1$, requires three such parameters that are defined using the derivative of the link function $\dot{\mu}(\cdot)$.
+The performance of Algorithm 1 (B-GLinCB) that solves Problem $1$, can be quantified using three such parameters that are defined using the derivative of the link function $\dot{\mu}(\cdot)$.
 Specifically, for any arm set ${\mathcal{X}}$, write optimal arm $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\mu\left(\langle x,\theta^{*}\rangle\right)$ and define,
-$\kappa\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}({\mathcal{D}})}\max_{x\in{\mathcal{X}}}\ \frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\kappa^{*}}}\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}\left({\mathcal{D}}\right)}\ {\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\widehat{\kappa}}}\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\right]$
+$\kappa\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}({\mathcal{D}})}\max_{x\in{\mathcal{X}}}\ \frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\kappa^{*}}}\coloneqq\max_{{\mathcal{X}}\in\texttt{supp}\left({\mathcal{D}}\right)}\ {\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)},\hskip 10.00002pt\frac{1}{{\widehat{\kappa}}}\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\right]$ (1)
 Remark 2.3.
@@
 Section 3 provides a quantified form of this trade-off.
-Algorithm 2 (RS-GLinCB) that solves Problem $2$, requires another such non-linearity parameter $\kappa$, defined as,
-$\kappa\coloneqq\max_{x\in\cup_{t=1}^{T}{\mathcal{X}}_{t}}\frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)}$
+Algorithm 2 (RS-GLinCB) that solves Problem $2$, requires another such non-linearity parameter $\kappa$ [^4] We overload the notation to match that in the literature.
+$\kappa$ in the context of Problem 1 is defined via (1), while $\kappa$ in the context of Problem 2 by (2)., defined as,
+$\kappa\coloneqq\max_{x\in\cup_{t=1}^{T}{\mathcal{X}}_{t}}\frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)}$ (2)
 We note that, here, $\kappa$ is defined considering the parameter vector $\theta^{*}$ in contrast to prior work on logistic bandits Faury et al.
 (2022), where its definition involved a maximization over all vectors $\theta$ with $\left\lVert\theta\right\rVert\leq S$ (known upper bound of $\|\theta^{\star}\|$).
-Hence, our definition of $\kappa$ is potentially much smaller and can lead to lower regret, compared to prior works.
+Hence, $\kappa$ as defined here is potentially much smaller and can lead to lower regret, compared to prior works.
 Standard to the CB literature with GLM rewards, we will assume that tight upper bounds on these parameters is known to the algorithms.
@@
 2.2 Optimal Design Policies
-G-optimal Design Given an arm set ${\mathcal{X}}$, the G-optimal design policy $\pi_{G}$ is the solution of the following optimization problem: $\min_{\lambda\in\Delta({\mathcal{X}})}\max_{x\in{\mathcal{X}}}{\left|\left|x\right|\right|}^{2}_{\mathbf{U}(\lambda)^{-1}}$, where $\mathbf{U}(\lambda)=\mathbb{E}_{x\sim\lambda}[xx^{\mathsf{T}}]$.
+G-optimal Design Given an arm set ${\mathcal{X}}$, the G-optimal design policy $\pi_{G}$ is the solution of the following optimization problem: $\operatorname*{arg\,min}_{\lambda\in\Delta({\mathcal{X}})}\max_{x\in{\mathcal{X}}}{\left|\left|x\right|\right|}^{2}_{\mathbf{U}(\lambda)^{-1}}$, where $\mathbf{U}(\lambda)=\mathbb{E}_{x\sim\lambda}[xx^{\mathsf{T}}]$.
 Now consider the following optimization problem, also known as the D-optimal design problem: $\max_{\lambda\in\Delta({\mathcal{X}})}\log\mathsf{Det}(\mathbf{U}(\lambda))$.
 This is a concave maximization problem as opposed to the G-optimal design which is non-convex.
-We have the following equivalence theorem due to Kiefer and Wolfowitz (Kiefer and Wolfowitz, 1960):
+We have the following equivalence theorem due to Kiefer and Wolfowitz Kiefer and Wolfowitz (1960):
 Lemma 2.5 (Keifer-Wolfowitz).
@@
 (2021).
-The distributional design policy is defined using a collection of tuples =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i}):p_{1},\ldots,p_{n}\geq 0\text{ and }\sum_{i}p_{i}=1\}$, wherein each ${\mathbf{M}}_{i}$ is a $d\times d$ positive semi-definite matrix and $n\leq 4d\log{d}$.
+The distributional optimal design policy is defined using a collection of tuples =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i}):p_{1},\ldots,p_{n}\geq 0\text{ and }\sum_{i}p_{i}=1\}$, wherein each ${\mathbf{M}}_{i}$ is a $d\times d$ positive semi-definite matrix and $n\leq 4d\log{d}$.
 The collection $\mathcal{M}$ is detailed next.
-Let $\text{softmax}_{\alpha}\left(\left[s_{1},\ldots,s_{k}\right]\right)$ denote the probability distribution where the $i^{th}$ element is sampled with probability $\frac{s_{i}^{\alpha}}{\sum_{j=1}^{k}s_{j}^{\alpha}}$.
+Let $\text{softmax}_{\alpha}\left(\left\{s_{1},\ldots,s_{k}\right\}\right)$ denote the probability distribution where the $i^{th}$ element is sampled with probability $\frac{s_{i}^{\alpha}}{\sum_{j=1}^{k}s_{j}^{\alpha}}$.
 For a specific =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i})\}_{i=1}^{n}$, and each $i\in[n]$ write $\pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)=\text{softmax}_{\alpha}(\{\left\lVert x\right\rVert_{{\mathbf{M}}_{i}}^{2}:x\in{\mathcal{X}}\})$.
@@
 $\pi\left({\mathcal{X}}\right)=\begin{cases}\pi_{G}\left({\mathcal{X}}\right)&\text{ with probability $1/2$}\\ \pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)&\text{ with probability $p_{i}/2$}\end{cases}$
-Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of (Ruan et al., 2021) to find the collection $\mathcal{M}$; see Algorithm 4 of Ruan et al.
+Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of Ruan et al.
+(2021) to find the collection $\mathcal{M}$; see Algorithm 4 of Ruan et al.
 (2021).
@@
 3 B-GLinCB
-In this section, we present B-GLinCB (Algorithm 1) that solves Problem 1 described in Section 2, which enforces constraints of limited adaptivity setting M1.
-Given limited adaptivity budget $M\in\mathbb{N}$, our algorithm first computes the batch length for each of the $M$ batches (i.e., rounds where the policy remains constant).
-We build upon the batch length construction in Gao et al.
-(2019); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.
-Batch Lengths: For each batch $k\in[M]$, let ${\mathcal{T}}_{k}$ denote all the consecutive rounds within the $k^{th}$ batch.
-We will refer to the first batch ${\mathcal{T}}_{1}$ as the warm-up batch.
-The batch lengths $\tau_{k}\coloneqq|{\mathcal{T}}_{k}|$, $k\in[M]$ are calculated as follows:
-$\tau_{1}:=\left(\frac{\sqrt{\kappa}\ e^{3S}d^{2}\gamma^{2}}{S}\alpha\right)^{2/3},$ $\quad\tau_{2}:=\alpha,\quad\tau_{k}:=\alpha\sqrt{\tau_{k-1}},\;\text{for }k\in[3,M]$ (1)
-where $\gamma\coloneqq 30RS\sqrt{d\log{T}}$  [^3] Recall that $R$ provides an upper bound on the stochastic rewards and $S$ is an upper bound on the norm of $\theta^{*}$.
-and $\alpha=T^{\frac{1}{2(1-2^{-M+1})}}$ if $M\leq\log{\log T}$ and $\alpha=2\sqrt{T}$ otherwise.
-During the warm-up batch (Lines 2, 4), the algorithm follows the G-optimal design policy, $\pi_{G}$.
-At the end of the warm-up batch (Line 5), the algorithm computes the Maximum Likelihood Estimate (MLE), $\widehat{\theta}_{w}$, of $\theta^{*}$ [^4] In case the MLE lies outside the set $\{\theta^{*}:\left\lVert\theta^{*}\right\rVert\leq S\}$, we follow the projection step detailed in Appendix E., and design matrix $\mathbf{V}\coloneqq\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}$, with parameter $\lambda=20Rd\log{T}$.
-Now, for each batch $k\geq 2$ and every round $t\in{\mathcal{T}}_{k}$, the algorithm updates ${\mathcal{X}}_{t}$ by eliminating arms from it using the confidence bounds (see Equation (5)) computed in the previous batches (Line 12).
-The algorithm next computes $\widetilde{{\mathcal{X}}}_{t}$, a scaled version of ${\mathcal{X}}_{t}$, as follows
-$\widetilde{{\mathcal{X}}}_{t}\coloneqq\left\{\sqrt{{\dot{\mu}(\langle x,\widehat{\theta}_{w}\rangle)}/{\beta(x)}}\ \ \ x:\text{ for all }\ x\in{\mathcal{X}}_{t}\right\}.$ (2)
-Finally, we use the distributional optimal design policy $\pi_{k}$, on the scaled arm set $\widetilde{{\mathcal{X}}}_{t}$, to sample the next arm (Line 13).
-At the end of every batch, we equally divide the batch ${\mathcal{T}}_{k}$ into two sets ${\mathcal{A}}$ and ${\mathcal{B}}$.
-We use samples from ${\mathcal{A}}$ to compute the estimator $\widehat{\theta}_{k}$ and the scaled design matrix ${\mathbf{H}}_{k}$.
-The rounds in ${\mathcal{B}}$ are used to compute $\pi_{k+1}$, the distributional optimal design policy for the next batch.
-It is important to note, while the policy $\pi_{k}$ is utilized in each round (Line 13) to draw arms, it is updated (to $\pi_{k+1}$) only at the end of the batch.
-Hence, conforming to setting M1, the algorithm updates the selection policy at $M$ rounds that were decided upfront.
-Confidence Bounds: The scaled design matrix ${\mathbf{H}}_{k}$, an estimator of the Hessian, is computed at the end of each batch $k\in{2,\ldots,M}$ (Line 16):
-${\mathbf{H}}_{k}=\sum_{t\in\mathcal{A}}\left({\dot{\mu}(\langle x_{t},\widehat{\theta}_{w}\rangle)}/{\beta(x_{t})}\right)x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}},\quad\text{where}\quad\beta(x)=\exp\left(R\min\left\{2S,\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right\}\right)$ (3)
-where $\mathcal{A}$ is the first half of ${\mathcal{T}}_{k}$. Using this, we define the upper and lower confidence bounds ($UCB_{k}$ and $LCB_{k}$) computed at the end of batch ${\mathcal{T}}_{k}$:
-$UCB_{k}(x)$$\coloneqq\begin{cases}\langle x,\widehat{\theta}_{w}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}&k=1\\ \langle x,\widehat{\theta}_{k}\rangle+\gamma\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}&k>1\end{cases},$ (4)
-$LCB_{k}(x)$$\coloneqq\begin{cases}\langle x,\widehat{\theta}_{w}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}&k=1\\ \langle x,\widehat{\theta}_{k}\rangle-\gamma\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}&k>1\end{cases}$ (5)
-Remark 3.1.
-The confidence bounds employed by the algorithm exhibit a significant distinction between the first batch and subsequent batches.
-While the first batch’s bounds are influenced by the parameter $\kappa$, subsequent batches utilize $\kappa$-independent bounds.
-This difference arises from the use of the standard design matrix $\mathbf{V}$ in the first batch and a scaled design matrix ${\mathbf{H}}_{k}$ (equation 3) in later batches, leveraging the self-concordance property of GLM rewards to achieve $\kappa$-independence.
-Notably, the first batch’s confidence bounds influence the scaling factor $\beta(x)$ in later batches, creating a trade-off (addressed in the regret analysis in Appendix A) where an inaccurate estimate of $\widehat{\theta}_{w}$ can exponentially increase the scaling factor and confidence bounds.
-In Theorem 3.2 and Corollary 3.3, we present our regret guarantee for B-GLinCB.
-Detailed proofs for both are provided in Appendix A.
-The computational efficiency of B-GLinCB is discussed in Appendix  D.
 Algorithm 1 B-GLinCB: Batched Generalized Linear Bandits Algorithm
@@
 18:  end for
+In this section, we present B-GLinCB (Algorithm 1) that solves Problem 1 described in Section 2, which enforces constraints of limited adaptivity setting M1.
+Given limited adaptivity budget $M\in\mathbb{N}$, our algorithm first computes the batch length for each of the $M$ batches (i.e., determine rounds where the policy remains constant).
+We build upon the batch length construction in Gao et al.
+(2019); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.
+[^5] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm.
+Batch Lengths: For each batch $k\in[M]$, let ${\mathcal{T}}_{k}$ denote all the rounds within the $k^{th}$ batch.
+We will refer to the first batch ${\mathcal{T}}_{1}$ as the warm-up batch.
+The batch lengths $\tau_{k}\coloneqq|{\mathcal{T}}_{k}|$, $k\in[M]$ are calculated as follows:
+$\tau_{1}:=\left(\frac{\sqrt{\kappa}\ e^{3S}d^{2}\gamma^{2}}{S}\alpha\right)^{2/3},$ $\quad\tau_{2}:=\alpha,\quad\tau_{k}:=\alpha\sqrt{\tau_{k-1}},\;\text{for }k\in[3,M]$ (3)
+where $\gamma\coloneqq 30RS\sqrt{d\log{T}}$  [^6] Recall that $R$ provides an upper bound on the stochastic rewards and $S$ is an upper bound on the norm of $\theta^{*}$.
+and $\alpha=T^{\frac{1}{2(1-2^{-M+1})}}$ if $M\leq\log{\log T}$ and $\alpha=2\sqrt{T}$ otherwise.
+During the warm-up batch (Lines 2, 4), the algorithm follows the G-optimal design policy, $\pi_{G}$.
+At the end of the warm-up batch (Line 5), the algorithm computes the Maximum Likelihood Estimate (MLE), $\widehat{\theta}_{w}$, of $\theta^{*}$ [^7] In case the MLE lies outside the set $\{\theta^{*}:\left\lVert\theta^{*}\right\rVert\leq S\}$, we apply the projection step detailed in Appendix E., and design matrix $\mathbf{V}\coloneqq\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}$, with parameter $\lambda=20Rd\log{T}$.
+Now, for each batch $k\geq 2$ and every round $t\in{\mathcal{T}}_{k}$, the algorithm updates ${\mathcal{X}}_{t}$ by eliminating arms from it using the confidence bounds (see Equation (7)) computed in the previous batches (Line 12).
+The algorithm next computes $\widetilde{{\mathcal{X}}}_{t}$, a scaled version of ${\mathcal{X}}_{t}$, as follows, with $\beta(x)$ define in equation  (5),
+$\widetilde{{\mathcal{X}}}_{t}\coloneqq\left\{\sqrt{{\dot{\mu}(\langle x,\widehat{\theta}_{w}\rangle)}/{\beta(x)}}\ \ \ x:\ x\in{\mathcal{X}}_{t}\right\}.$ (4)
+Finally, we use the distributional optimal design policy $\pi_{k}$, on the scaled arm set $\widetilde{{\mathcal{X}}}_{t}$, to sample the next arm (Line 13).
+At the end of every batch, we equally divide the batch ${\mathcal{T}}_{k}$ into two sets ${\mathcal{A}}$ and ${\mathcal{B}}$.
+We use samples from ${\mathcal{A}}$ to compute the estimator $\widehat{\theta}_{k}$ and the scaled design matrix ${\mathbf{H}}_{k}$.
+The rounds in ${\mathcal{B}}$ are used to compute $\pi_{k+1}$, the distributional optimal design policy for the next batch.
+It is important to note while the policy $\pi_{k}$ is utilized in each round (Line 13) to draw arms, it is updated (to $\pi_{k+1}$) only at the end of the batch.
+Hence, conforming to setting M1, the algorithm updates the selection policy at $M$ rounds that were decided upfront.
+Confidence Bounds: The scaled design matrix ${\mathbf{H}}_{k}$, an estimator of the Hessian, is computed at the end of each batch $k\in{2,\ldots,M}$ (Line 16):
+${\mathbf{H}}_{k}=\sum_{t\in\mathcal{A}}\left({\dot{\mu}(\langle x_{t},\widehat{\theta}_{w}\rangle)}/{\beta(x_{t})}\right)x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}},\quad\text{where}\quad\beta(x)=\exp\left(R\min\left\{2S,\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right\}\right)$ (5)
+where $\mathcal{A}$ is the first half of ${\mathcal{T}}_{k}$. Using this, we define the upper and lower confidence bounds ($UCB_{k}$ and $LCB_{k}$) computed at the end of batch ${\mathcal{T}}_{k}$:
+$UCB_{k}(x)$$\coloneqq\begin{cases}\langle x,\widehat{\theta}_{w}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}&k=1\\ \langle x,\widehat{\theta}_{k}\rangle+\gamma\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}&k>1\end{cases},$ (6)
+$LCB_{k}(x)$$\coloneqq\begin{cases}\langle x,\widehat{\theta}_{w}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}&k=1\\ \langle x,\widehat{\theta}_{k}\rangle-\gamma\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}&k>1\end{cases}$ (7)
+Remark 3.1.
+The confidence bounds employed by the algorithm exhibit a significant distinction between the first batch and subsequent batches.
+While the first batch’s bounds are influenced by the parameter $\kappa$, subsequent batches utilize $\kappa$-independent bounds.
+This difference arises from the use of the standard design matrix $\mathbf{V}$ in the first batch and a scaled design matrix ${\mathbf{H}}_{k}$ (equation 5) in later batches, leveraging the self-concordance property of GLM rewards to achieve $\kappa$-independence.
+Notably, the first batch’s confidence bounds influence the scaling factor $\beta(x)$ in later batches, creating a trade-off (addressed in the regret analysis in Appendix A) where an inaccurate estimate of $\widehat{\theta}_{w}$ can exponentially increase the scaling factor and confidence bounds.
+In Theorem 3.2 and Corollary 3.3, we present our regret guarantee for B-GLinCB.
+Detailed proofs for both are provided in Appendix A.
+The computational efficiency of B-GLinCB is discussed in Appendix  D.
 Theorem 3.2.
-Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$, where
+Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$  [^8] Note that $\textsc{R}_{T}$ is expected regret. See the 2., where
 $\textsc{R}_{1}$ $=O\left(RSd\ \left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)T^{\frac{1}{2\left(1-2^{1-M}\right)}}\log{T}\right)\text{ and }$
@@
 Remark 3.5.
-The $\kappa$-dependent batch construction is a crucial feature of our algorithm, enabling effective estimation of $\dot{\mu}{\langle\theta,x\rangle}$ at the end of the first batch.
+The $\kappa$-dependent batch construction is a crucial feature of our algorithm, enabling effective estimation of $\dot{\mu}{\left(\langle x,\theta^{*}\rangle\right)}$ at the end of the first batch.
 Since the first batch incurs regret linear in its length, achieving a $\kappa$-independent guarantee requires the first batch to be $o(\sqrt{T})$.
@@
 14:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})$ and observe reward $r_{t}$.
-15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{w}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
+15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{o}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
 16:     end if
@@
 Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^9] This optimization problem is non-convex.
-However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$,
+However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
 $\min_{\theta}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}\right\rVert_{{\mathbf{H}}(\theta)}\text{s.t.
-}\left\lVert\theta-\widehat{\theta}_{w}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (6)
+}\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (8)
 Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$.
-After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$.
+After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ [^10] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm..
 Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$.
@@
 Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting.
-Unlike existing literature (Faury et al., 2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.4 and B.5).
+Unlike existing literature Faury et al.
+(2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.3 and B.4).
 We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.11), hence incurring a small regret in these rounds.
 Remark 4.4.
-Unlike (Abbasi-Yadkori et al., 2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$).
+Unlike Abbasi-Yadkori et al.
+(2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$).
 The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret.
-This modification is a crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
+This modification is crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
 Remark 4.5.
-The feasible set for the optimization stated in 6 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability.
+The feasible set for the optimization stated in 8 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability.
 Deviating from existing literature on GLM Bandits which projects the estimate into the ball set of radius $S$ ($\{\theta:\left\lVert\theta\right\rVert\leq S\}$), our projection step leads to tighter regret guarantees; notably, the leading $\sqrt{T}$ term is free of parameters $S$ (and $R$).
@@
 (2024) regarding the possibility of obtaining $S$-free regret in the $\sqrt{T}$ term in logistic bandits.
+Remark 4.6.
+The regret guarantees of the logistic bandit algorithms in Abeille et al.
+(2021); Lee et al.
+(2024) have a second-order term that is minimum of an arm-geometry dependent quantity (see Theorem 3 of Lee et al.
+(2024)) and a $\kappa$-dependent term similar to our regret guarantee.
+Although our analysis is not able to accommodate this arm-geometry dependent quantity, we underscore that our algorithm is computationally efficient while the above works are not.
+In fact, to the best of our knowledge, the other known efficient algorithms for logistic bandits Faury et al.
+(2022); Zhang and Sugiyama (2023) also do not achieve the arm-geometry dependent regret term.
+It can be interesting to design an efficient algorithm that is able to achieve the same guarantees in the second-order regret term as in Abeille et al.
+(2021); Lee et al.
+(2024).
 5 Experiments
@@
 Bottom: (left) Execution times of ECOLog and RS-GLinCB for different values of $\kappa$ (low $\kappa=9.3$ and high $\kappa=141.6$) for Logistic rewards.
-(right) Execution times of GLOC and RS-GLinCB for different values of $\kappa$ (low $\kappa=17.6$ and high $\kappa=202.3$) Probit rewards.
+(right) Execution times of GLOC and RS-GLinCB for different values of $\kappa$ (low $\kappa=17.6$ and high $\kappa=202.3$) for Probit rewards.
 We tested the practicality of our algorithm RS-GLinCB against various baselines for logistic and generalized linear bandits.
-For these experiments, we adjusted the warm-up threshold constant in RS-GLinCB to $0.01$ and used data from both warm-up and non-warm-up rounds to estimate $\widetilde{\theta}$.
+For these experiments, we adjusted the Switching Criterion I threshold constant in RS-GLinCB to $0.01$ and used data from both Switching Criteria (I and II) rounds to estimate $\widetilde{\theta}$.
 These modifications do not affect the overall efficiency as $\widetilde{\theta}$ is calculated only $O(\log(T))$ times.
@@
 Logistic.
-We compared RS-GLinCB against ECOLog (Faury et al., 2022) and GLOC (Jun et al., 2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting.
+We compared RS-GLinCB against ECOLog Faury et al.
+(2022) and GLOC Jun et al.
+(2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting.
 The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$.
@@
 Probit.
-For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB (Filippi et al., 2010).
+For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB Filippi et al.
+(2010).
 The dimension was set to $d=5$ and number of arms per round to $K=20$.
@@
 In particular, we design algorithms B-GLinCB and RS-GLinCB that obtain optimal regret guarantees for two prevalant limited adaptivity settings M1 and M2 respectively.
-A key feature of our guarantees are that their leading terms are independent of a instance dependent parameter $\kappa$ that captures non-linearity.
-To the best of our knowledge, our paper provides the first algorithm for CB algorithms for GLM rewards under limited adaptivity (and otherwise) that achieves $\kappa$-independent regret.
+A key feature of our guarantees are that their leading terms are independent of an instance dependent parameter $\kappa$ that captures non-linearity.
+To the best of our knowledge, our paper provides the first algorithms for the CB problem with GLM rewards under limited adaptivity (and otherwise) that achieve $\kappa$-independent regret.
 The regret guarantee of RS-GLinCB, not only aligns with the best-known guarantees for Logistic Bandits but enhances them by removing the dependence on $S$ (upper bound on $\left\lVert\theta^{*}\right\rVert$) in the leading term of the regret and therefore resolves a conjecture in Lee et al.
@@
 We believe that the dependence on $d$ along with the ${\widehat{\kappa}}$ term is not tight and improving the dependence is a relevant direction for future work.
-Appendix for “”
+Acknowledgments and Disclosure of Funding
+Siddharth Barman gratefully acknowledges the support of the Walmart Center for Tech Excellence (CSR WMGT-23-0001) and a SERB Core research grant (CRG/2021/006165).
+Appendix
 Appendix A Regret Analysis of B-GLinCB
-Additional Notation: We write c to denote absolute constant(s) that appears throughout our analysis. Our analysis also utilizes the following function
+A.1 Additional Notation
+We write c to denote absolute constant(s) that appears throughout our analysis. Our analysis also utilizes the following function
 $\gamma(\lambda)=24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}.$ (9)
@@
 ${\mathbf{H}}^{*}_{k}=\sum_{t\in{\mathcal{T}}_{k}}\dot{\mu}\left(\langle x_{t},\theta^{*}\rangle\right)x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}.$
-We will show that ${\mathbf{H}}_{k}$ is an estimator of ${\mathbf{H}}^{*}_{k}$.
+We will show that ${\mathbf{H}}_{k}$ dominates ${\mathbf{H}}^{*}_{k}$ with high probability.
 Furthermore, we assume that the MLE estimator $\theta^{*}$ obtained by minimizing the log-loss objective always satisfies $\left\lVert\widehat{\theta}\right\rVert\leq S$.
@@
 Hence, the assumption $\left\lVert\widehat{\theta}\right\rVert\leq S$ is non-limiting.
+A.2 Concentration Inequalities and Confidence Intervals
 Lemma A.1 (Bernstein’s Inequality).
@@
 $\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}$ $\leq\sqrt{(1+2RS)}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}}$
-$=\sqrt{1+2RS}\left\lVert\left(\theta^{*}-\widehat{\theta}\right){\mathbf{G}}\right\rVert_{{\mathbf{G}}^{-1}}$
+$=\sqrt{1+2RS}\left\lVert{\mathbf{G}}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
 $=\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\langle\theta^{*},x_{j}\rangle-\langle\widehat{\theta},x_{j}\rangle\right)\alpha(x,\theta^{*},\widehat{\theta})x_{j}+\frac{\lambda}{1+2RS}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
@@
 $=\max_{y\in{\mathcal{B}}}\langle y-y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle\!+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
-$\leq\left\lVert y-y_{\varepsilon}\right\rVert_{2}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
-$\leq\varepsilon\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
+$\leq\max_{y\in{\mathcal{B}}}\left\lVert y-y_{\varepsilon}\right\rVert_{2}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
+$\leq\max_{y\in{\mathcal{B}}}\varepsilon\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+\langle y_{\varepsilon},{{\mathbf{H}}^{*}}^{-1/2}\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\rangle$
 Rearranging, we obtain
@@
 Proof.
-This result is derived directly from Lemma A.2 and the definition of (see ). By applying the lemma, we obtain
+This result is derived directly from Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). By applying the lemma, we obtain
 $|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{w}}\right\rVert_{{\mathbf{H}}^{*}}\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\gamma(\lambda)$
 Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$.
-This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$ which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$.
+This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$  [^11] For logistic rewards, we note that instead of using the worst case bound of $\kappa$, one can make use of an upper bound on $S\coloneqq\left\lVert\theta\right\rVert$ as done in Mason et al.
+[2022] That is, we lower bound $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\geq\dot{\mu}\left({S\left\lVert x\right\rVert}\right)$ and use ${\mathbf{H}}^{*}\succeq\sum_{t}\dot{\mu}\left({S\left\lVert x_{t}\right\rVert}\right)x_{t}x_{t}^{\intercal}$.
+which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$.
 ∎
+A.3 Preliminary Lemmas
 Lemma A.5.
@@
 $\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)$ $\leq\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\exp\left(R|\langle x,\widehat{\theta}_{w}-\theta^{*}\rangle|\right)$
-Via Lemma A.4 we have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma(\lambda)$.
+Via Corollary A.4 we have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma(\lambda)$.
 Additionally, given that $\left\lVert\widehat{\theta}\right\rVert,\left\lVert\theta^{*}\right\rVert\leq S$ and $\left\lVert x\right\rVert\leq 1$ we also have $|\langle x,\widehat{\theta}_{w}\rangle-\langle x,\theta^{*}\rangle|\leq 2S$.
@@
 Let $\widehat{\theta}_{k}$ be the estimator of $\theta^{*}$ calculated at the end of the $k^{th}$ batch.
-Then for any vector $x$ the following holds with probability greater than $1-\frac{\log{\log T}}{T^{2}}$ for every batch $k>=2$.
-$|\langle x,\theta^{*}-\widehat{\theta}\rangle|\leq\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}\gamma{\left(\lambda\right)}$
+Then for any vector $x$ the following holds with probability greater than $1-\frac{\log{\log T}}{T^{2}}$ for every batch $k\geq 2$.
+$|\langle x,\theta^{*}-\widehat{\theta}_{k}\rangle|\leq\left\lVert x\right\rVert_{{\mathbf{H}}_{k}^{-1}}\gamma{\left(\lambda\right)}$
 Proof.
-This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 7). According to the lemma, we have
-$|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|$ $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{w}}\right\rVert_{{\mathbf{H}}^{*}_{k}}$
+This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). According to the lemma, we have
+$|\langle x,\theta^{*}-\widehat{\theta}_{k}\rangle|$ $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{k}}\right\rVert_{{\mathbf{H}}^{*}_{k}}$
 $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\gamma(\lambda)$
@@
 Lemma A.9.
-Let $x\in{\mathcal{X}}$ be the selected in any round of batch $k\geq 2$ in the algorithm, and let $x^{*}$ be the optimal arm in the arm set ${\mathcal{X}}$, i.e., $x^{*}=\operatorname*{arg\,max}_{x\in}{\mathcal{X}}$.
+Let $x\in{\mathcal{X}}$ be the selected in any round of batch $k\geq 2$ in the algorithm, and let $x^{*}$ be the optimal arm in the arm set ${\mathcal{X}}$, i.e., $x^{*}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\mu\left(\langle x,\theta^{*}\rangle\right)$.
 With probability greater than $1-\frac{\log{\log{T}}}{T^{2}}$, the following inequality holds-
-$\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\leq 6\phi(\lambda)\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\!\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
+$\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\leq 6\phi(\lambda)\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\!\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
 Proof.
-We begin by applying Taylor’s theorem, which yields the following for some $z\in\left[\langle x^{*},\theta^{*}\rangle,\langle x,\theta^{*}\rangle\right]$
+We begin by applying Taylor’s theorem, which yields the following for some $z$ between $\langle x,\theta^{*}\rangle$ and $\langle x^{*},\theta^{*}\rangle$
 $\left|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)\right|$ (14)
@@
 $\leq\dot{\mu}(z)\left(\left|\langle x^{*},\theta^{*}\rangle-\langle x^{*},\widehat{\theta}_{k-1}\rangle\right|+\left|\langle x^{*},\widehat{\theta}_{k-1}\rangle-\langle x,\widehat{\theta}_{k-1}\rangle\right|+\left|\langle x,\widehat{\theta}_{k-1}\rangle-\langle x,\theta^{*}\rangle\right|\right)$
-$\leq 2\dot{\mu}(z)\left(\left\lVert x^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\gamma(\lambda)+\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}}\gamma(\lambda)\right)$ (via Lemma A.7)
-$\leq 2\dot{\mu}(z)\gamma(\lambda)\left(\sqrt{\frac{\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\sqrt{\frac{\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
-$\leq 2\gamma(\lambda)\sqrt{\dot{\mu}(z)}\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\sqrt{\dot{\mu}(z)}\gamma(\lambda)\sqrt{\frac{\dot{\mu}(z)\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}$ (13)
+$\leq 2\dot{\mu}(z)\left(\left\lVert x^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\gamma(\lambda)+\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\gamma(\lambda)\right)$ (via Corollary A.7)
+$\leq 2\dot{\mu}(z)\gamma(\lambda)\left(\sqrt{\frac{\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\sqrt{\frac{\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
+$\leq 2\gamma(\lambda)\sqrt{\dot{\mu}(z)}\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+2\sqrt{\dot{\mu}(z)}\gamma(\lambda)\sqrt{\frac{\dot{\mu}(z)\beta(x)}{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ (15)
 We now invoke Lemmas C.2 and A.4 to obtain
@@
 $\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-z|}{2}\right\}\right)\sqrt{\beta(x)}$
-$\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-\langle x^{*},\theta^{*}\rangle|}{2}\right\}\right)\sqrt{\beta(x)}$ (since $z\in\left[\langle x^{*}_{t},\theta^{*}\rangle,\langle x,\theta^{*}\rangle\right]$ )
-$\leq\exp\left(\min\left\{S,\frac{3\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+2\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)}{2}\right\}\right)\sqrt{\beta(x)}$ (using Lemma A.4)
+$\leq\exp\left(\min\left\{S,\frac{|\langle x,\theta^{*}\rangle-\langle x,\widehat{\theta}_{w}\rangle|+|\langle x,\theta^{*}\rangle-\langle x^{*},\theta^{*}\rangle|}{2}\right\}\right)\sqrt{\beta(x)}$ (since $z\in\left[\langle x,\theta^{*}\rangle,\langle x^{*}_{t},\theta^{*}\rangle\right]$ )
+$\leq\exp\left(\min\left\{S,\frac{3\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+2\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)}{2}\right\}\right)\sqrt{\beta(x)}$ (using Lemma A.4 and the elimination criteria)
 $\leq\exp\left(\min\left\{2S,2\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)+\sqrt{\kappa}\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\gamma\left(\lambda\right)\right\}\right)$ (substituting the definition of $\beta(x)$ )
@@
 Further, we can simplify each term in equation (15) as
-$\sqrt{\dot{\mu}(z)}\left(\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
-$\leq 2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}\exp\left(\min\left\{3S,3\sqrt{\kappa}\gamma\left(\lambda\right)\left(\left\lVert x\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\right)\right\}\right)\right)$
-$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
-$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
-$\quad\quad+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
-$\leq 6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}$
+$\sqrt{\dot{\mu}(z)}\left(\sqrt{\frac{\dot{\mu}(z)\beta(x^{*})}{\dot{\mu}\left(\langle x^{*},\widehat{\theta}_{w}\rangle\right)}}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
+$\leq 2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\exp\left(\min\left\{3S,3\sqrt{\kappa}\gamma\left(\lambda\right)\left(\left\lVert x\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}\right)\right\}\right)\right)$
+$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ (via Claim A.8)
+$\leq 6\frac{\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\kappa}\ e^{3S}\ \gamma(\lambda)}{S}^{2}\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$
+$\quad\quad+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$
+$\leq 6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+2\gamma(\lambda)\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$
 Finally, we substitute the above bound in (15) to obtain
-$|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)|\leq$ $6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
-$\quad\quad\quad+2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)$
+$|\mu\left(\langle x^{*},\theta^{*}\rangle\right)-\mu\left(\langle x,\theta^{*}\rangle\right)|\leq$ $6\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\phi(\lambda)\left(\left\lVert x^{*}\right\rVert_{\mathbf{V}^{-1}}+\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\right)\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
+$\quad\quad\quad+2\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)$
 For Phase $k$, the distribution of the remaining arms after the elimination step (${\mathcal{X}}$ in line 12 of the Algorithm 1) is represented as ${\mathcal{D}}_{k}$.
@@
 The proof here invokes Lemma A.9. We begin by noting that
-$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}}$ $\leq 4\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-}1}\right]$
+$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ $\leq 4\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
 $\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Jensen’s inequality)
@@
 We also have
-$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}}\right)\right]$
-$\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
-$\leq 2\min\left\{\sqrt{\kappa^{*}}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right],\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\left\lVert x\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}\right\}$ (using the definition of $\kappa^{*}$ for the first bound and Jensen for the second)
+$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\left(\left\lVert\widetilde{x}^{*}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}+\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right)\right]$
+$\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\sqrt{\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
+$\leq 2\min\left\{\sqrt{\frac{1}{\kappa^{*}}}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right],\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\dot{\mu}\left(\langle x^{*},\theta^{*}\rangle\right)\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}\right\}$ (using the definition of $\kappa^{*}$ for the first bound and Jensen for the second)
 $\leq c\left(\sqrt{\frac{d\log{d}}{\kappa^{*}\tau_{k-1}}}\land\sqrt{\frac{d^{2}}{{\widehat{\kappa}}\tau_{k-1}}}\right)$ (using Lemma A.17)
 Substituting the above bounds in Lemma A.9 we obtained the stated inequality. This completes the proof. ∎
-A.1 Proof of Theorem 3.2
+A.4 Proof of Theorem 3.2
 We trivially upper bound the regret incurred during the warm-up batch as $\tau_{1}R$; recall that $R$ denotes the upper bound on the rewards and $\tau_{1}$ denotes the length of the first (warm-up) batch; see equation (5)).
@@
 $R_{T}\leq O\left(\left(\sqrt{\kappa}d^{3}\ e^{3S}\ RST^{\frac{1}{2\left(1-2^{-M}\right)}}\log{T}\right)^{2/3}+\left(\frac{d}{\sqrt{{\widehat{\kappa}}}}\land\sqrt{\frac{\ d\log d}{\kappa^{*}}}\right)RST^{\frac{1}{2\left(1-2^{-M}\right)}}\sqrt{d\log{T}}\right).$
-A.2 Optimal Design Guarantees
+A.5 Optimal Design Guarantees
 In this section, we study the optimal design policies utilized in different batches of the algorithm.
@@
 Therefore,
-$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2d^{2}.$
+$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}_{G}}^{2}\right]\leq 2d^{2}.$
 Lemma A.15 (Matrix Chernoff Tropp et al. [2015], Ruan et al. [2021]).
 Let $x_{1},x_{3},\ldots,x_{n}\sim{\mathcal{D}}$ be vectors, with $\left\lVert x_{t}\right\rVert\leq 1$, then we have
-$\mathbb{P}\left\{3\varepsilon n{\mathbf{I}}+\sum_{i=1}^{n}x_{i}x_{i}^{\mathsf{T}}\succeq\frac{n}{8}\ \operatorname*{\mathbb{E}}_{x\sim{\mathcal{D}}}\left[xx^{\mathsf{T}}\right]\right\}\geq 1-2d\exp\left(\frac{-\varepsilon n}{8}\right)$
+$\mathbb{P}\left\{3\varepsilon n{\mathbf{I}}+\sum_{i=1}^{n}x_{i}x_{i}^{\mathsf{T}}\succeq\frac{n}{8}\ \operatorname*{\mathbb{E}}_{x\sim{\mathcal{D}}}\left[xx^{\mathsf{T}}\right]\right\}\geq 1-2d\exp\left(-\frac{\varepsilon n}{8}\right)$
 Corollary A.16.
@@
 Similarly ${\mathbf{H}}_{k}$, with $\lambda\geq 6\log{(Td)}$ satisfies the following for each batch $k\geq 2$ with probability greater than $1-\frac{1}{T^{2}}$
-${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{G}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
+${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
 Proof.
@@
 We note that the analysis of Ruan et al.
-[2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}}^{2}$.
+[2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$.
 We obtain such a bound here and use it in the analysis.
@@
 $\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{1}}\right)$ (18)
-$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{k}}\right)\quad\forall k\in[M]$ (17)
+$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{k}}\right)\quad\forall k\in[M]$ (19)
 We also have that for sufficiently large $T\gtrsim O\left(d^{32}(\log{2T})^{2}\right)$, the following holds with probability greater than $1-\frac{\log{\log{T}}}{T}$
@@
 $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$
-$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\geq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
-We obtain the first to inequalities, (16) and (17), by a direct use of Corollary A.16.
+$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
+We obtain the first two inequalities, (18) and (19), by a direct use of Corollary A.16.
 For (20) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M).
@@
 Appendix B Regret Analysis of RS-GLinCB
-Recall ${\mathcal{T}}_{o}$ denotes the set of warm-up rounds. We write $\tau_{w}$ to denote the size of the set $\tau_{w}=|{\mathcal{T}}_{o}|$. We define the following (scaled) data matrix
+Recall ${\mathcal{T}}_{o}$ denotes the set of rounds when switching criterion I is satisfied.
+We write $\tau_{o}$ to denote the size of the set $\tau_{o}=|{\mathcal{T}}_{o}|$.
+We define the following (scaled) data matrix
 ${\mathbf{H}}^{*}_{w}=\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}\ .$
@@
 $\mathbb{P}\left[\exists t\geq 1:\lVert S_{t}\rVert_{{{\mathbf{H}}}^{-1}_{t}}\geq\frac{\sqrt{\lambda}}{2}+\frac{2}{\sqrt{\lambda}}\log\left(\frac{\det({\mathbf{H}}_{t})^{\frac{1}{2}}}{\lambda^{d/2}\delta}\right)+\frac{2}{\sqrt{\lambda}}d\log(2)\right]\leq\delta$
+B.1 Confidence Sets for Switching Criterion I rounds
 Lemma B.2.
-At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the warm-up rounds ${\mathcal{T}}_{o}$.
-With probability at least $1-\delta$ we have
-$\lVert\widehat{\theta}_{w}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma.$
+At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$. With probability at least $1-\delta$ we have
+$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma.$
 Proof.
@@
 $\lVert\widehat{\theta}_{o}-\theta^{*}$ $\rVert_{{\mathbf{H}}^{*}_{w}}\leq S\sqrt{R^{2}\lambda}$
-$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{w}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
+$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{o}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
 Simplifying constants and setting $\sqrt{R^{2}\lambda}=\textbf{c}\sqrt{d\log(T)+\log(1/\delta)}$, we have $\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}\leq\textbf{c}RS\sqrt{d\log(T/\delta)+\log(1/\delta)}\leq\textbf{c}RS\sqrt{d\log{(T/\delta)}}$.
 ∎
-Next, we obtain a bound on the number of warm-up rounds.
+B.2 Confidence Sets for non-Switching Criterion I rounds
+We define ${\mathcal{E}}_{w}$ be the event defined in Lemma B.2, that is, ${\mathcal{E}}_{w}=\{\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma\}$.
 Lemma B.3.
-Algorithm 2, during its entire execution, enters the enters the warmup segment at most $2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$ times.
-Proof.
-Via the warm-up condition (Line 6) $\lVert x\rVert_{\mathbf{V}^{-1}}^{2}>1/(R^{2}\kappa\gamma^{2})$.
-Let $\mathbf{V}_{m}$ be the sequence of $\mathbf{V}$ matrices (line 7 of 2) for $m\in{\mathcal{T}}_{o}$.
-That is, $\mathbf{V}_{1}=\lambda{\mathbf{I}},\mathbf{V}_{m}=\sum_{s\in[m-1]\setminus{\mathcal{T}}_{o}}x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$).
-Thus,
-$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\geq\frac{\tau_{w}}{R^{2}\kappa\gamma^{2}}$ (19)
-Furthermore, by the Elliptic Potential Lemma (Lemma B.11) we have
-$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\leq 2d\log\left(1+\frac{\tau_{w}}{\lambda d}\right)$ (20)
-Combining (20) and (19) we have
-$\tau_{w}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{w}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
-We define ${\mathcal{E}}_{w}$ be the event define in Lemma B.2, that is, ${\mathcal{E}}_{w}=\{\lVert\widehat{\theta}_{w}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma_{t}\}$.
-Lemma B.4.
-If in round $t$ the warmup criteria is not satisfied and event ${\mathcal{E}}_{w}$, holds, we have
+If in round $t$ the switching criteria I is not satisfied and the event ${\mathcal{E}}_{w}$ holds, we have
 $\lvert\langle x,\widehat{\theta}_{o}-\theta^{*}\rangle\rvert\leq\frac{1}{R}\quad\text{ for all }x\in{\mathcal{X}}_{t}.$
@@
 ∎
-Now we define the following quantities:
+Recall that $\tau$ is the round when the Switching Criterion II (Line 9) is satisfied. Now we define the following quantities:
 $g_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda_{\tau}\theta$
 ${{\mathbf{H}}}_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$
-$\Theta$ =$=\left\{\theta|\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}\right\}$
+$\Theta$ =$=\left\{\theta:\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}\right\}$
 $\tilde{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})$
@@
 Lemma B.6.
-Let $\delta\in(0,1)$.
-Then, under event ${\mathcal{E}}_{w}$, with probability $1-\delta$, $\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta\coloneqq\textbf{c}\sqrt{d\log(T/\delta)}$.
+Let $\delta\in(0,1)$. Then, under event ${\mathcal{E}}_{w}$, with probability $1-\delta$, $\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta$.
 Proof.
-We have for all main rounds $s\in[\tau-1]\setminus{\mathcal{T}}_{o}$,
+We have for all rounds $s\in[\tau-1]\setminus{\mathcal{T}}_{o}$,
 $\lvert\langle x_{s},\theta^{*}-\widehat{\theta}_{\tau}\rangle\rvert$ $\leq\left\lVert x_{s}\right\rVert_{\mathbf{V}^{-1}}\left\lVert\theta^{*}-\widehat{\theta}_{\tau}\right\rVert_{\mathbf{V}}$ (Cauchy-Schwarz)
@@
 Let the event in lemma B.6 be denoted by ${\mathcal{E}}_{\tau}$, or in other words, ${\mathcal{E}}_{\tau}=\{\lVert\widehat{\theta}_{\tau}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta\}$.
+B.3 Bounding the instantaneous regret
+In this subsection, we will only consider rounds $t\in[T]$ which does not satisfy Switching Criterion I.
+Let $x_{t}\in{\mathcal{X}}_{t}$ be the played arm defined via line 14 Algorithm 2.
+Further, let $x^{*}_{t}\in{\mathcal{X}}_{t}$ be the best available arm in that round.
 Corollary B.7.
-If round $t$ is a main round for Algorithm 2, then, under event ${\mathcal{E}}_{\tau}$, $\lvert\langle x,\widehat{\theta}_{\tau}-\theta^{*}\rangle\rvert\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{-1}}$
+Under the event ${\mathcal{E}}_{\tau}$, for all $x\in{\mathcal{X}}_{t}$, we have, $\lvert\langle x,\widehat{\theta}_{\tau}-\theta^{*}\rangle\rvert\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{-1}}~{}.$
 Proof.
@@
 Lemma B.8.
-For main rounds in Algorithm 2, let $x_{t}$ be the played arm defined via line 14 Algorithm 2.
-Then, under event ${\mathcal{E}}_{\tau}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq 2\sqrt{2}\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}~{}.$
+Under event ${\mathcal{E}}_{\tau}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq 2\sqrt{2}\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}~{}.$
 Proof.
@@
 Lemma B.9.
-For non-warm up rounds, the arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$ (b) for $x_{t},x^{*}_{t}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
+The arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$, (b) $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
 Proof.
-Suppose $x^{\prime}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}LCB_{w}(x)$. Now, we have, for all $x\in{\mathcal{X}}_{t}$,
+Suppose $x^{\prime}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}LCB_{o}(x)$. Now, we have, for all $x\in{\mathcal{X}}_{t}$,
 $\lvert\langle x,\widehat{\theta}_{o}-\theta^{*}\rangle\rvert$ $\leq\left\lVert x\right\rVert_{{\mathbf{H}}^{*-1}_{w}}\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}$ (Cauchy-Schwarz)
@@
 $\leq\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ ($\mathbf{V}\preceq\kappa{\mathbf{H}}^{*}_{w}$)
-Thus, for $x^{*}_{t}$, we can write,
-$\langle x^{*}_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{*}_{t},\theta^{*}\rangle$
-$\geq\langle x^{\prime},\theta^{*}\rangle$ (optimality of $x^{*}_{t}$)
-$\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
-Thus, $x^{*}_{t}\in{\mathcal{X}}^{\prime}_{t}$.
-Since $x_{t}$ is also in ${\mathcal{X}}^{\prime}_{t}$,
-$\langle x_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
-$\geq\langle x_{t}^{*},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ ($x^{\prime}$ has max LCB)
+Thus, $UCB_{o}(x^{*}_{t})=\langle x^{*}_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}\geq\langle x^{*}_{t},\theta^{*}\rangle\geq\langle x^{\prime},\theta^{*}\rangle\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}=LCB_{o}(x^{\prime})~{},$ where the second inequality is due to optimality of $x^{*}_{t}$.
+Hence, $x^{*}_{t}$ is not eliminated, implying $x^{*}_{t}\in{\mathcal{X}}^{\prime}_{t}$.
+This completes the proof of (a).
+Since $x_{t}$ is also in ${\mathcal{X}}^{\prime}_{t}$ (by definition),
+$UCB_{o}(x_{t})=\langle x_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
+$\geq\langle x_{t}^{*},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ ($x^{\prime}$ has max $LCB_{o}(\cdot)$)
 Again, using the fact that $\langle x^{*}_{t},\widehat{\theta}_{o}\rangle\geq\langle x^{*}_{t},\theta^{*}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ and $\langle x_{t},\widehat{\theta}_{o}\rangle\leq\langle x_{t},\theta^{*}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$, we obtain,
 $\langle x_{t},\theta^{*}\rangle+2\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{*}_{t},\theta^{*}\rangle-2\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$
-$\implies\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ $\leq 2\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}+2\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$
-Finally, since this is a non warm-up round, $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}<\frac{1}{R\gamma\sqrt{\kappa}}$ for all $x\in{\mathcal{X}}_{t}$. Plugging this above,
+[…]
+Finally, since Switching Criterion I is not satisfied in this round, $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}<\frac{1}{R\gamma\sqrt{\kappa}}$ for all $x\in{\mathcal{X}}_{t}$. Plugging this above,
 $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
-B.1 Proof of Theorem 4.2
-We prove Theoreom 4.2 in this section
+B.4 Proof of Theorem 4.2
+In this subsection, we complete the proof of the regret bound of RS-GLinCB(Algorithm 2).
+We first restate Theoreom 4.2 and then prove it.
+For every round $t\in[T]$, we use $x_{t}\in{\mathcal{X}}_{t}$ to denote the arm played by the algorithm and $x^{*}_{t}$ to denote the best available arm in that round.
+Theorem B.10 (Theorem 4.2).
+Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}d\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{RT}/{\delta}}\right)+\ \kappa d^{2}R^{5}S^{2}\log^{2}\left({{T}/{\delta}}\right)\big{)}.$
 Proof.
@@
 $\textsc{R}_{T}$ $=\sum_{t\in[T]}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$
-$\leq R\tau_{w}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$ (Upper bound of $R$ for warm-up rounds)
+$\leq R\tau_{o}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$ (Upper bound of $R$ for rounds in ${\mathcal{T}}_{o}$)
 $\leq\textbf{c}R^{3}\kappa\gamma^{2}\log(T/\delta)+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ (some $z\in[\langle x_{t},\theta^{*}\rangle,\langle x_{t}^{*},\theta^{*}\rangle]$; lemma B.11)
@@
 $\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\left\lVert\tilde{x}_{t}\right\rVert_{{\mathbf{H}}_{t}^{*-1}}^{2}}$ (Cauchy-Schwarz)
-$\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot 2d\log\left(1+\frac{RT}{\lambda d}\right)}$ (Elliptic Potential Lemma; $\left\lVert\tilde{x}_{t}\right\rVert_{2}\leq R$)
+$\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot 2d\log\left(1+\frac{RT}{\lambda d}\right)}$ (Lemma B.12; $\left\lVert\tilde{x}_{t}\right\rVert_{2}\leq R$)
 $\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}.$
@@
 $\textsc{R}_{T}\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}+\textbf{c}R^{5}S^{2}\kappa\log(T/\delta)^{2}.$
-B.2 Proof of Lemma 4.1
-We first restate Lemma 4.1.
+B.5 Bounding number of policy updates: Proof of Lemma 4.1
+We first obtain a bound on the number of rounds when Switching Criterion I is satisfied.
+Then we restate Lemma 4.1 and present its proof.
+Here, we use ${\mathcal{T}}_{o}$ to denote the collection of all rounds till $T$ for which Switching Criterion I is satisfied.
+Lemma B.11.
+Algorithm 2, during its entire execution, satisfies the Switching Criterion I at most $2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$ times.
+Proof.
+Recall that Switching Criterion I (Line 4) is satisfied, when $\lVert x\rVert_{\mathbf{V}^{-1}}^{2}>1/(R^{2}\kappa\gamma^{2})$ for some $x\in{\mathcal{X}}_{t}$.
+Let $\mathbf{V}_{m}$ be the sequence of $\mathbf{V}$ matrices (line 6 of Algorithm 2) for $m\in{\mathcal{T}}_{o}$.
+That is, $\mathbf{V}_{1}=\lambda{\mathbf{I}},\mathbf{V}_{m}=\sum_{s\in[m-1]\cap{\mathcal{T}}_{o}}x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
+In these rounds, by Line 5 of Algorithm 2, we have that the arm played $x_{t}$ is such that $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}$.
+Therefore,
+$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x_{t}\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\geq\frac{\tau_{o}}{R^{2}\kappa\gamma^{2}}$ (22)
+Furthermore, by the Elliptic Potential Lemma (Lemma B.12) we have
+$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x_{t}\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\leq 2d\log\left(1+\frac{\tau_{o}}{\lambda d}\right)$ (23)
+Combining (23) and (22) we have
+$\tau_{o}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{o}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
 Lemma (4.1).
@@
 Proof.
-Note that in Algorithm 2, policy changes happen only in the warm-up rounds and when the determinant of ${\mathbf{H}}_{t}$ doubles in the main round.
-Thus, total number of policy switches is upper bounded by the number of warm-up rounds and the number of times $\det({\mathbf{H}}_{t})$ doubles.
-The first quantity is bounded by Lemma B.3 while the second quantity is bounded by Lemma B.14.
-Thus in total, the number of policy changes in Algorithm A is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
-B.3 Some Useful Lemmas
+Note that in Algorithm 2, policy changes happen only in the rounds when either of the Switching Criteria are triggered.
+The number of times Switching Criterion I is triggered is bounded by Lemma B.11.
+On the other hand, the number of times Switching Criterion II is triggered is equal to the number of times determinant of ${\mathbf{H}}_{t}$ doubles, which is bounded by Lemma B.15.
+Thus in total, the number of policy changes in Algorithm 2 is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
+B.6 Some Useful Lemmas
 Lemma B.12 (Elliptic Potential Lemma (Lemma 10 Abbasi-Yadkori et al. [2011])).
@@
 Define the set $\{1=\tau_{1},\tau_{2}\ldots\tau_{m}=t\}$ such that: $\det(\mathbf{V}_{\tau_{i+1}})\geq 2\det(\mathbf{V}_{\tau_{i}})$ but $\det(\mathbf{V}_{\tau_{i+1}-1})<2\det(\mathbf{V}_{\tau_{i}})$ for $i\in\{2,\ldots m-1\}$.
-Then, the number of time doubling happens,i.e.$m$ is at most $O(d\log(t))$.
+Then, the number of time doubling happens,i.e.$m$, is at most $O(d\log(t))$.
 Proof.
@@
 $\dot{\mu}(z_{1})\exp(-R(z-z_{1}))\leq$ $\dot{\mu}(z)\leq\dot{\mu}(z_{1})\exp(R(z-z_{1}))$
-Putting $z=z_{2}$ establishes 24. To show 23, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
+Putting $z=z_{2}$ establishes 26. To show 25, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
 […]
@@
 Note that similar exercise can be repeated with $z_{2}\leq z_{1}$ to get the same result for $z_{2}$.
-For 25, we have, by application of 24, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
+For 27, we have, by application of 26, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
 […]
@@
 $=\dot{\mu}(z_{1})\left(\frac{1}{R|z_{1}-z_{2}|}+\frac{\exp(-R|z_{1}-z_{2}|)-1}{R^{2}|z_{1}-z_{2}|^{2}}\right)$
-$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of [Abeille et al., 2021])
+$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of Abeille et al. [2021])
 Next we state some nice properties of the GLM family that is the key in deriving Lemma C.1.
@@
 Remark C.5.
-In [Faury, 2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set.
-Thereafter the author notes that the techniques developed in [Faury, 2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter).
+In Faury [2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set.
+Thereafter the author notes that the techniques developed in Faury [2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter).
 However, the claim regarding self-concordance of GLMs is not true in general.
@@
 Let the maximum number of arms available every round be $K$.
-Furher, let the computational cost of an oracle that solves the non-convex optimization 6 be $NC_{opt}$.
+Furher, let the computational cost of an oracle that solves the non-convex optimization 8 be $NC_{opt}$.
 Computational Cost of B-GLinCB: In the B-GLinCB algorithm, we employ the log-loss oracle at the end of each batch.
@@
 Hence, the amortized cost per round of B-GLinCB is $O(K\log{\log T}+d\log{d}+C_{opt})$.
-Computational Cost of RS-GLinCB: In the RS-GLinCB algorithm, the estimator $\widehat{\theta}_{w}$ is computed during each warmup round.
-Additionally, during non-warmup rounds, the estimator $\widehat{\theta}$ is computed a maximum of $O(\log(T))$ times.
+Computational Cost of RS-GLinCB: In the RS-GLinCB algorithm, the estimator $\widehat{\theta}_{o}$ is computed each time Switching Criterion I is triggered.
+Additionally, during rounds when Switching Criterion I is not triggered, the estimator $\widehat{\theta}_{\tau}$ is computed a maximum of $O(\log(T))$ times.
 These computations involve utilizing both the log-loss oracle and the non-convex projection oracle.
-Furthermore, in each non-warmup round, the algorithm executes an elimination step.
-This yields an amortized time complexity of $O(C_{opt}\log{T}+NC_{opt}\log^{2}(T)+K)$ per round.
+Furthermore, in each non-Switching Criterion I round, the algorithm executes an elimination step.
+This yields an amortized time complexity of $O(C_{opt}\log{(T)}+NC_{opt}\log^{2}(T)+K)$ per round.
 Performance in Practice: As evident from Fig.
@@
 We allowed no operations for 10 seconds after every run to let the CPU temperature come back to normal, in case the execution heats up the CPU, thereby causing subsequent runs to slow down.
-Comparison with ECOLog [Faury et al., 2022] shows that execution time for RS-GLinCB is significantly smaller.
+Comparison with ECOLog Faury et al.
+[2022] shows that execution time for RS-GLinCB is significantly smaller.
 We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round.
@@
 Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
-While comparing with GLOC [Jun et al., 2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes.
+While comparing with GLOC Jun et al.
+[2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes.
 Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$.
@@
 Using the above fact, we obtain $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\succeq\frac{1}{2+c}{\mathbf{H}}^{*}$.
-We now follow closely the proof outlined in Appendix B.3 of [Abeille et al., 2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
+We now follow closely the proof outlined in Appendix B.3 of Abeille et al. [2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
 ${\mathcal{L}}_{t}(\theta)-{\mathcal{L}}_{t}(\theta^{*})-\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\theta-\theta^{*}\rangle$ $=\left\lVert\theta-\theta^{*}\right\rVert_{\tilde{{\mathbf{G}}}(\theta,\theta^{*})}^{2}$
@@
 Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$.
-By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of [Abeille et al., 2021]), we have, for $\theta^{*}$.
+By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of Abeille et al.
+[2021]), we have, for $\theta^{*}$.
 ${\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})$ $\leq\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}(\theta^{*},\widehat{\theta})^{-1}}^{2}$

### v5

@@
 In the last few decades, much progress has been made in designing algorithms for special classes of reward models, e.g.
-linear model Auer (2002); Chu et al.
-(2011); Abbasi-Yadkori et al.
-(2011); Lattimore and Szepesvári (2020), logistic model Faury et al.
-(2020); Abeille et al.
-(2021); Faury et al.
-(2022); Zhang and Sugiyama (2023) and generalized linear models Filippi et al.
-(2010); Li et al.
-(2017).
+linear model Aue (02); CLRS (11); AYPS (11); LS (20), logistic model FACF (20); AFC (21); FAJC (22); ZS (23) and generalized linear models FCGS (10); LLZ (17).
 However, despite this progress, there is a key challenge that prevents deployment of CB algorithms in the real world.
 Practical situations often allow for very limited adaptivity, i.e., do not allow CB algorithms to update their policy at all rounds.
-For example, in clinical trials Group et al.
-(1997), each trial involves administering medical treatments to a cohort of patients, with medical outcomes observed and collected for the entire cohort at the conclusion of the trial.
+For example, in clinical trials ABCED (23), each trial involves administering medical treatments to a cohort of patients, with medical outcomes observed and collected for the entire cohort at the conclusion of the trial.
 This data is then used to design the treatment for the next phase of the trial.
-Similarly, in online advertising Schwartz et al.
-(2017) and recommendations Li et al.
-(2010), updating the policy after every iteration during deployment is often infeasible due to infrastructural constraints.
-A recent line of work Ruan et al.
-(2021); Ren and Zhou (2024); Han et al.
-(2020); Hanna et al.
-(2022); Perchet et al.
-(2015); Gao et al.
-(2019); Simchi-Levi and Xu (2022) tries to address this limitation by developing algorithms that try to minimize cumulative regret while ensuring that only a limited number of policy updates occur.
+Similarly, in online advertising SBF (17) and recommendations LCLS (10), updating the policy after every iteration during deployment is often infeasible due to infrastructural constraints.
+A recent line of work RYZ (21); RZ (24); HZZ+ (20); HYF (22); PRCS (15); GHRZ (19); SLX (22) tries to address this limitation by developing algorithms that try to minimize cumulative regret while ensuring that only a limited number of policy updates occur.
 Across these works, two settings (called M1 and M2 from here onwards) of limited adaptivity have been popular.
 Both M1, M2 provide a budget $M$ to the algorithm, determining the number of times it can update its policy.
-In M1 Ruan et al.
-(2021); Hanna et al.
-(2023), the algorithm is required to decide upfront a sub-sequence of $M$ rounds where policy updates will occur.
-While in M2 Abbasi-Yadkori et al.
-(2011); Ruan et al.
-(2021)), the algorithm is allowed to adaptively decide (during its course) when to update its policy.
-+–+++++++++++ Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting Ruan et al.
-(2021); Han et al.
-(2020), and optimal regret guarantees were obtained when the arm feature vectors were stochastically generated.
-Similarly, in their seminal work on linear bandits, Abbasi-Yadkori et al.
-(2011) developed algorithms for the M2 setting and proved optimal regret guarantees with no restrictions on the arm vectors.
+In M1 RYZ (21); HYF (23), the algorithm is required to decide upfront a sub-sequence of $M$ rounds where policy updates will occur.
+While in M2 AYPS (11); RYZ (21)), the algorithm is allowed to adaptively decide (during its course) when to update its policy.
+Limited adaptivity algorithms were recently proposed for the CB problem with linear reward models under the M1 setting RYZ (21); HZZ+ (20), and optimal regret guarantees were obtained when the arm feature vectors were stochastically generated.
+Similarly, in their seminal work on linear bandits, AYPS (11) developed algorithms for the M2 setting and proved optimal regret guarantees with no restrictions on the arm vectors.
 While these results provide tight regret guarantees for linear reward models, extending them to generalized linear models is quite a challenge.
@@
 When $M=O(\log\log{T})$, we prove an $\tilde{O}(T^{2^{M-1}/(2^{M}-2)})$ regret guarantee.
-While the algorithm bears a slight resemblance to the one in Ruan et al.
-(2021), direct utilization of their key techniques (distributional optimal design) results in a regret guarantee that scales linearly with the instance dependent non-linearity $\kappa$.
+While the algorithm bears a slight resemblance to the one in RYZ (21), direct utilization of their key techniques (distributional optimal design) results in a regret guarantee that scales linearly with the instance dependent non-linearity $\kappa$.
 On the other hand, the leading terms in our regret guarantee for B-GLinCB have no dependence on $\kappa$.
-To achieve this, we make novel modifications to the key technique of distributional optimal design in Ruan et al.
-(2021).
+To achieve this, we make novel modifications to the key technique of distributional optimal design in RYZ (21).
 Along with this, the rounds for policy updates are also chosen more carefully (in a $\kappa$ dependent fashion), leading to a stronger regret guarantee.
 $\bullet$ We propose RS-GLinCB, an algorithm that solves the CB problem for bounded (almost surely) generalized linear reward models (Definition 2.1) under the M2 setting of limited adaptivity.
-RS-GLinCB builds on a similar algorithm in Abbasi-Yadkori et al.
-(2011) by adding a novel context-dependent criterion for determining if a policy update is needed.
+RS-GLinCB builds on a similar algorithm in AYPS (11) by adding a novel context-dependent criterion for determining if a policy update is needed.
 This new criterion allows us to prove optimal regret guarantee ($\tilde{O}(\sqrt{T})$) with only $O(\log^{2}T)$ updates to the policy.
 It is quite crucial for the generalized linear reward settings since, without it, the resultant regret guarantees have a linear dependence on $\kappa$.
-$\bullet$ Our work also resolves a conjecture in Lee et al.
-(2024) by proving an optimal ($\tilde{O}(\sqrt{T})$) regret guarantee (for the CB problem with logistic reward model) that does not depend polynomially on $S$ (the known upper bound on the size of the model parameters, i.e.
-$\|\theta^{\star}\|\leq S$, See Section 2111This requires a non-convex projection.
-We discuss its convex relaxation in Appendix E..
-RS-GLinCB is, to our knowledge, the first CB algorithm for generalized linear reward models that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret.
-We also perform experiments in Section 5 that validate its superiority both in terms of regret and computational efficiency in comparison to other baseline algorithms proposed in Jun et al.
-(2017) and Faury et al.
-(2020).
+$\bullet$ To the best of our knowledge, RS-GLinCB is the first CB algorithm for generalized linear reward models that is both computationally efficient (amortized $O(\log T)$ computation per round) and incurs optimal regret.
+We also perform experiments in Section 5 that validate its superiority both in terms of regret and computational efficiency in comparison to other baseline algorithms proposed in JBNW (17) and FACF (20).
 1.2 Important Remarks on Contributions and Comparison with Prior Work
@@
 Not only do we prove $\kappa$-independent regret guarantees under the limited adaptivity constraint, we also characterize a broad class of generalized linear reward models for which a $\kappa$-independent regret guarantee can be achieved.
-Specifically, our results imply that the CB problem with generalized linear reward models originally proposed in Filippi et al.
-(2010) and subsequently studied in literature Li et al.
-(2017); Jun et al.
-(2017); Russac et al.
-(2020) admits a $\kappa$-independent regret.
+Specifically, our results imply that the CB problem with generalized linear reward models originally proposed in FCGS (10) and subsequently studied in literature LLZ (17); JBNW (17); RCG (20) admits a $\kappa$-independent regret.
 Remark 1.2 (Computational efficiency).
 Efforts to reduce the total time complexity to be linear in $T$ have been active in the CB literature with generealized linear rewards models.
-For e.g., Jun et al.
-(2017) recently devised computationally efficient algorithms but they suffer from regret dependence on $\kappa$.
-Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models Faury et al.
-(2020); Abeille et al.
-(2021), and the algorithms were subsequently made computationally efficient in Faury et al.
-(2022); Zhang and Sugiyama (2023).
+For e.g., JBNW (17) recently devised computationally efficient algorithms but they suffer from regret dependence on $\kappa$.
+Optimal ($\kappa$-independent) guarantees were recently achieved for logistic reward models FACF (20); AFC (21), and the algorithms were subsequently made computationally efficient in FAJC (22); ZS (23).
 However, the techniques involved rely heavily on the structure of the logistic model and do not easily extend to more general models.
 To the best of our knowledge, ours is the first work that achieves optimal $\kappa$-independent regret guarantees for bounded generalized linear reward models while remaining computationally efficient [^1] While RS-GLinCB and B-GLinCB have total running time of $\widetilde{O}(T)$, their per-round complexity can reach $O(T)$.
-This stands in contrast to Faury et al.
-(2022), which maintains efficiency in both total and per-round time complexity..
+This stands in contrast to FAJC (22), which maintains efficiency in both total and per-round time complexity..
 Remark 1.3 (Self Concordance of bounded GLMs).
 In order to prove $\kappa$-independent regret guarantees, we prove a key result about self concordance of bounded (almost surely) generalized linear models (Definition 2.1) in Lemma 2.2.
-This result was postulated in Filippi et al.
-(2010) for GLMs (with the same definition as ours), but no proof was provided.
-While Faury et al.
-(2020; 2022) partially tackled this issue for logistic reward models [^3] In Faury (2021), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made; however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
+This result was postulated in FCGS (10) for GLMs (with the same definition as ours), but no proof was provided.
+While FACF (20); FAJC (22) partially tackled this issue for logistic reward models [^2] In Fau (21), a claim about $\kappa$ independent regret for all generalized linear models with bounded $\theta^{*}$ was made; however, we can construct counterexamples to this claim (see Appendix C, Remark C.5)., in our work, we prove self concordance for much more general generalized linear models.
 2 Notations and Preliminaries
@@
 2.1 Instance Dependent Non-Linearity Parameters
-As in prior works Faury et al.
-(2020; 2022), we define instance dependent parameters that capture non-linearity of the underlying instance and critically impact our algorithm design.
+As in prior works FACF (20); FAJC (22), we define instance dependent parameters that capture non-linearity of the underlying instance and critically impact our algorithm design.
 The performance of Algorithm 1 (B-GLinCB) that solves Problem $1$, can be quantified using three such parameters that are defined using the derivative of the link function $\dot{\mu}(\cdot)$.
@@
 $\kappa\coloneqq\max_{x\in\cup_{t=1}^{T}{\mathcal{X}}_{t}}\frac{1}{\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)}$ (2)
-We note that, here, $\kappa$ is defined considering the parameter vector $\theta^{*}$ in contrast to prior work on logistic bandits Faury et al.
-(2022), where its definition involved a maximization over all vectors $\theta$ with $\left\lVert\theta\right\rVert\leq S$ (known upper bound of $\|\theta^{\star}\|$).
+We note that, here, $\kappa$ is defined considering the parameter vector $\theta^{*}$ in contrast to prior work on logistic bandits FAJC (22), where its definition involved a maximization over all vectors $\theta$ with $\left\lVert\theta\right\rVert\leq S$ (known upper bound of $\|\theta^{\star}\|$).
 Hence, $\kappa$ as defined here is potentially much smaller and can lead to lower regret, compared to prior works.
@@
 This is a concave maximization problem as opposed to the G-optimal design which is non-convex.
-We have the following equivalence theorem due to Kiefer and Wolfowitz Kiefer and Wolfowitz (1960):
+We have the following equivalence theorem due to Kiefer and Wolfowitz KW (60):
 Lemma 2.5 (Keifer-Wolfowitz).
@@
 Nonetheless, when the arm sets ${\mathcal{X}}_{t}$ are drawn from a distribution, it is possible to extend the guarantee, albeit with a worse dependence on $d$; see Section A.5 in Appendix A.
-Improving this dependence motivates the need of studying Distributional optimal design and towards this we utilize the results of Ruan et al.
-(2021).
+Improving this dependence motivates the need of studying Distributional optimal design and towards this we utilize the results of RYZ (21).
 The distributional optimal design policy is defined using a collection of tuples =$\mathcal{M}=\{(p_{i},{\mathbf{M}}_{i}):p_{1},\ldots,p_{n}\geq 0\text{ and }\sum_{i}p_{i}=1\}$, wherein each ${\mathbf{M}}_{i}$ is a $d\times d$ positive semi-definite matrix and $n\leq 4d\log{d}$.
@@
 $\pi\left({\mathcal{X}}\right)=\begin{cases}\pi_{G}\left({\mathcal{X}}\right)&\text{ with probability $1/2$}\\ \pi_{{\mathbf{M}}_{i}}\left({\mathcal{X}}\right)&\text{ with probability $p_{i}/2$}\end{cases}$
-Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of Ruan et al.
-(2021) to find the collection $\mathcal{M}$; see Algorithm 4 of Ruan et al.
-(2021).
+Given a collection of arm sets $\left\{{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{s}\right\}$ (called core set) sampled from the distribution ${\mathcal{D}}$, we utilize Algorithm 2 of RYZ (21) to find the collection $\mathcal{M}$; see Algorithm 4 of RYZ (21).
 Overall, the computed $\mathcal{M}$ induces a policy $\pi$ that upholds the following guarantee.
-Lemma 2.6 (Theorem 5, Ruan et al. (2021)).
+Lemma 2.6 (Theorem 5, RYZ (21)).
 Let $\pi$ be the Distributional optimal design policy that has been learnt from $s$ independent samples ${\mathcal{X}}_{1},\ldots{\mathcal{X}}_{s}\sim{\mathcal{D}}$.
@@
 4:  end for
-5:  Compute $\widehat{\theta}_{w}=\operatorname*{arg\,min}_{\begin{subarray}{c}\theta\end{subarray}}\sum_{s\in{\mathcal{T}}_{1}}\ell(\theta,x_{s},r_{s})$ and matrix $\mathbf{V}=\lambda{\mathbf{I}}+\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}$.
-6:  Initialize policy $\pi_{1}$ as G-optimal design.
+5:  Compute $\widehat{\theta}_{w}=\operatorname*{arg\,min}_{\begin{subarray}{c}\theta\end{subarray}}\sum_{s\in{\mathcal{T}}_{1}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$ and matrix $\mathbf{V}=\lambda{\mathbf{I}}+\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}$.
+6:  Initialize policy $\pi_{2}$ as G-optimal design.
 7:  for batches $k=2$ to $M$ do
@@
 12:        end for
-13:        Scale ${\mathcal{X}}_{t}$, as in (4), to obtain $\widetilde{{\mathcal{X}}_{t}}$., then sample $x_{t}\sim\pi_{k-1}\left(\widetilde{{\mathcal{X}}_{t}}\right)$.
+13:        Scale ${\mathcal{X}}_{t}$, as in (4), to obtain $\widetilde{{\mathcal{X}}_{t}}$., then sample $x_{t}\sim\pi_{k}\left(\widetilde{{\mathcal{X}}_{t}}\right)$.
 14:     end for
@@
 16:     Define ${\mathbf{H}}_{k}=\lambda{\mathbf{I}}+\sum_{t\in\mathcal{A}}\frac{\dot{\mu}\left(\langle x,\widehat{\theta}_{w}\rangle\right)}{\beta(x_{t})}x_{t}x_{t}^{\mathsf{T}}$, and $\widehat{\theta}_{k}=\operatorname*{arg\,min}_{\theta}\sum_{s\in\mathcal{A}}\ell(\theta,x_{s},r_{s})$.
-17:     Compute Distributional optimal design policy $\pi_{k}$ using the arm sets $\{{\mathcal{X}}_{t}\}_{t\in\mathcal{B}}$.
+17:     Compute Distributional optimal design policy $\pi_{k+1}$ using the arm sets $\{\widetilde{{\mathcal{X}}_{t}}\}_{t\in\mathcal{B}}$.
 18:  end for
@@
 Given limited adaptivity budget $M\in\mathbb{N}$, our algorithm first computes the batch length for each of the $M$ batches (i.e., determine rounds where the policy remains constant).
-We build upon the batch length construction in Gao et al.
-(2019); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.
+We build upon the batch length construction in GHRZ (19); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.
 [^4] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm.
@@
 $\tau_{1}:=\left(\frac{\sqrt{\kappa}\ e^{3S}d^{2}\gamma^{2}}{S}\alpha\right)^{2/3},$ $\quad\tau_{2}:=\alpha,\quad\tau_{k}:=\alpha\sqrt{\tau_{k-1}},\;\text{for }k\in[3,M]$ (3)
-where $\gamma\coloneqq 30RS\sqrt{d\log{T}}$  [^6] Recall that $R$ provides an upper bound on the stochastic rewards and $S$ is an upper bound on the norm of $\theta^{*}$.
+where $\gamma\coloneqq 30(RS)^{3/2}\sqrt{d\log{T}}$  [^5] Recall that $R$ provides an upper bound on the stochastic rewards and $S$ is an upper bound on the norm of $\theta^{*}$.
 and $\alpha=T^{\frac{1}{2(1-2^{-M+1})}}$ if $M\leq\log{\log T}$ and $\alpha=2\sqrt{T}$ otherwise.
@@
 Theorem 3.2.
-Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$  [^8] Note that $\textsc{R}_{T}$ is expected regret. See the 2., where
-$\textsc{R}_{1}$ $=O\left(RSd\ \left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)T^{\frac{1}{2\left(1-2^{1-M}\right)}}\log{T}\right)\text{ and }$
-$\textsc{R}_{2}$ $=O\left(\kappa^{1/3}d^{2}e^{2S}(RS\log{T})^{2/3}T^{\frac{1}{3\left(1-2^{1-M}\right)}}\right).$
+Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$  [^7] Note that $\textsc{R}_{T}$ is expected regret. See the Section 2., where
+$\textsc{R}_{1}$ $=O\left((RS)^{3/2}d\ \left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)T^{\frac{1}{2\left(1-2^{1-M}\right)}}\log{T}\right)\text{ and }$
+$\textsc{R}_{2}$ $=O\left(\kappa^{1/3}d^{2}e^{2S}RS(\log{T})^{2/3}T^{\frac{1}{3\left(1-2^{1-M}\right)}}\right).$
 Corollary 3.3.
 When the number of batches $M\geq\log{\log{T}}$, Algorithm 1 achieves a regret bound of
-$\textsc{R}_{T}$ $\leq\widetilde{O}\left(\left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)dRS\sqrt{T}+d^{2}e^{2S}(S^{2}R^{2}\kappa T)^{1/3}\right).$
+$\textsc{R}_{T}$ $\leq\widetilde{O}\left(\left(\sqrt{\frac{d}{{\widehat{\kappa}}}}\land\sqrt{\frac{1}{\kappa^{*}}}\right)d(RS)^{3/2}\sqrt{T}+d^{2}e^{2S}RS(\kappa T)^{1/3}\right).$
 Remark 3.4.
@@
 Algorithm 2 RS-GLinCB: Rarely-Switching GLM Bandit Algorithm
-1:  Initialize: $\mathbf{V}={\mathbf{H}}_{1}=\lambda{\mathbf{I}}$, ${\mathcal{T}}_{o}=\emptyset$, $\tau=1$, $\lambda\coloneqq d\log({T/\delta})/R^{2}$ and $\gamma\coloneqq 25RS\sqrt{d\log\left(\frac{T}{\delta}\right)}$.
+1:  Initialize: $\mathbf{V}={\mathbf{H}}_{1}=\lambda{\mathbf{I}}$, ${\mathcal{T}}_{o}=\emptyset$, $\tau=1$, $\lambda\coloneqq dR^{2}\log({T/\delta})/S$ and $\gamma\coloneqq 25R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log\left(\frac{ST}{\delta}\right)}$.
 2:  for rounds $t=1,\dots,T$ do
@@
 6:        Update ${\mathcal{T}}_{o}\leftarrow{\mathcal{T}}_{o}\cup\{t\}$, $\mathbf{V}\leftarrow\mathbf{V}+x_{t}x_{t}^{\mathsf{T}}$ and ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}$.
-7:        Compute $\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\theta}\sum_{s\in{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$.
-8:     else
+7:        Compute $\widetilde{\theta}_{o}=\operatorname*{arg\,min}_{\theta}\sum_{s\in{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$.
+8:        $\widehat{\theta}_{o}\leftarrow\texttt{Project}(\widetilde{\theta}_{o})$ via (8).
+9:     else
 […]
-10:           Set $\tau=t$ and $\widetilde{\theta}\leftarrow\operatorname*{arg\,min}_{\theta}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$ and
-11:           $\widehat{\theta}_{\tau}\leftarrow\texttt{Project}(\widetilde{\theta})$.
-12:        end if
-13:        Update ${\mathcal{X}}_{t}\leftarrow{\mathcal{X}}_{t}\setminus\{x\in{\mathcal{X}}_{t}:UCB_{o}(x)<\max_{z\in{\mathcal{X}}_{t}}LCB_{o}(z)\}$.
-14:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})$ and observe reward $r_{t}$.
-15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{o}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
-16:     end if
-17:  end for
+11:           Set $\tau=t$ and $\widetilde{\theta}\leftarrow\operatorname*{arg\,min}_{\theta}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$ and
+12:           $\widehat{\theta}_{\tau}\leftarrow\texttt{Project}(\widetilde{\theta})$ via (9).
+13:        end if
+14:        Update ${\mathcal{X}}_{t}\leftarrow{\mathcal{X}}_{t}\setminus\{x\in{\mathcal{X}}_{t}:UCB_{o}(x)<\max_{z\in{\mathcal{X}}_{t}}LCB_{o}(z)\}$.
+15:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})$ and observe reward $r_{t}$.
+16:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{o}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
+17:     end if
+18:  end for
 In this section we present RS-GLinCB (Algorithm 2) that solves Problem 2 described in Section 2, which enforces constraints of limited adaptivity setting M2.
-This algorithm incorporates a novel switching criterion (Line 4), extending the determinant-doubling approach of Abbasi-Yadkori et al.
-(2011).
-Additionally, we introduce an arm-elimination step (Line 13) to obtain tighter regret guarantees.
-Throughout this section, we set $\lambda=d\log({T/\delta})/R^{2}$ and $\gamma=25RS\sqrt{d\log\left({T}/{\delta}\right)}$.
+This algorithm incorporates a novel switching criterion (Line 4), extending the determinant-doubling approach of AYPS (11).
+Additionally, we introduce an arm-elimination step (Line 14) to obtain tighter regret guarantees.
+Throughout this section, we set $\lambda=dR^{2}\log({T/\delta})/S$ and $\gamma=25R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log\left({ST}/{\delta}\right)}$.
 At round $t$, on receiving an arm set ${\mathcal{X}}_{t}$, RS-GLinCB first checks the Switching Criterion I (Line 4).
@@
 Subsequently in Line 6, the set ${\mathcal{T}}_{o}$ is updated to include $t$; the design matrix $\mathbf{V}$ is updated as $\mathbf{V}\leftarrow\mathbf{V}+x_{t}x_{t}^{\intercal}$; and the scaled design matrix ${\mathbf{H}}_{t+1}$ is set to ${\mathbf{H}}_{t}$.
-The MLE is computed (Line 7) based on the data in the rounds in ${\mathcal{T}}_{o}$ to obtain $\widehat{\theta}_{o}$.
-When Switching Criterion I is not triggered, the algorithm first checks (Line 9) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered).
-If Switching Criterion II is triggered at round $t$, then in Line 10, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$.
+The MLE is computed (Line 7) based on the data in the rounds in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}_{o}$.
+This is subsequently projected to the ball $\{\theta\in\mathbb{R}^{d}:\left\lVert\theta\right\rVert\leq S\}$ via the following optimization problem [^8] This optimization problem is non-convex.
+However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
+$\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq S}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widetilde{\theta}_{o}\rangle\right)\right)-\lambda\widetilde{\theta}_{o}\right\rVert_{{\mathbf{H}}_{w}(\theta)^{-1}}$ (8)
+where ${\mathbf{H}}_{w}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
+When Switching Criterion I is not triggered, the algorithm first checks (Line 10) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered).
+If Switching Criterion II is triggered at round $t$, then in Line 11, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$.
 Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^9] This optimization problem is non-convex.
 However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
-$\min_{\theta}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}\right\rVert_{{\mathbf{H}}(\theta)}\text{s.t.
+$\widehat{\theta}_{\tau}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\left\lVert\sum_{s\in{[t-1]\setminus{\mathcal{T}}_{o}}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}-\lambda\widetilde{\theta}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}\text{s.t.
 }\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (9)
-Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$.
-After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ [^10] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm..
+Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$.
+After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 14) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ [^10] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm..
 Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$.
-For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}\sqrt{d\log{({T}/{\delta})}}$, and plays the arm $x_{t}$ with the highest index (Line 14).
-After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 15) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$.
+For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}R\sqrt{dS\log(ST/\delta)}$, and plays the arm $x_{t}$ with the highest index (Line 15).
+After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 16) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$.
 With this, the round $t$ ends and the algorithm moves to the next round.
@@
 Lemma 4.1.
-RS-GLinCB (Algorithm 2), during its entire execution, updates its policy at most $O(R^{4}S^{2}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
+RS-GLinCB (Algorithm 2), during its entire execution, updates its policy at most $O(\kappa d^{2}R^{6}S^{3}\log^{2}(T/\delta))$ times.
 Theorem 4.2.
-Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}d\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{RT}/{\delta}}\right)+\ \kappa d^{2}R^{5}S^{2}\log^{2}\left({{T}/{\delta}}\right)\big{)}.$
+Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}dRS^{\nicefrac{{1}}{{2}}}\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{ST}/{\delta}}\right)+\ \kappa d^{2}R^{7}S^{3}\log^{2}\left({{ST}/{\delta}}\right)\big{)}.$
 Remark 4.3.
 Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting.
-Unlike existing literature Faury et al.
-(2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.3 and B.4).
+Unlike existing literature FAJC (22), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.4 and B.5).
 We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.12), hence incurring a small regret in these rounds.
 Remark 4.4.
-Unlike Abbasi-Yadkori et al.
-(2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$).
+Unlike AYPS (11), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$).
 The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret.
@@
 Remark 4.5.
-The feasible set for the optimization stated in 8 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability.
-Deviating from existing literature on GLM Bandits which projects the estimate into the ball set of radius $S$ ($\{\theta:\left\lVert\theta\right\rVert\leq S\}$), our projection step leads to tighter regret guarantees; notably, the leading $\sqrt{T}$ term is free of parameters $S$ (and $R$).
-This resolves the conjecture made in Lee et al.
-(2024) regarding the possibility of obtaining $S$-free regret in the $\sqrt{T}$ term in logistic bandits.
-Remark 4.6.
-The regret guarantees of the logistic bandit algorithms in Abeille et al.
-(2021); Lee et al.
-(2024) have a second-order term that is minimum of an arm-geometry dependent quantity (see Theorem 3 of Lee et al.
-(2024)) and a $\kappa$-dependent term similar to our regret guarantee.
+The regret guarantees of the logistic bandit algorithms in AFC (21); LYJ (24) have a second-order term that is minimum of an arm-geometry dependent quantity (see Theorem 3 of LYJ (24)) and a $\kappa$-dependent term similar to our regret guarantee.
 Although our analysis is not able to accommodate this arm-geometry dependent quantity, we underscore that our algorithm is computationally efficient while the above works are not.
-In fact, to the best of our knowledge, the other known efficient algorithms for logistic bandits Faury et al.
-(2022); Zhang and Sugiyama (2023) also do not achieve the arm-geometry dependent regret term.
-It can be interesting to design an efficient algorithm that is able to achieve the same guarantees in the second-order regret term as in Abeille et al.
-(2021); Lee et al.
-(2024).
+In fact, to the best of our knowledge, the other known efficient algorithms for logistic bandits FAJC (22); ZS (23) also do not achieve the arm-geometry dependent regret term.
+It can be interesting to design an efficient algorithm that is able to achieve the same guarantees in the second-order regret term as in AFC (21); LYJ (24).
 5 Experiments
@@
 Logistic.
-We compared RS-GLinCB against ECOLog Faury et al.
-(2022) and GLOC Jun et al.
-(2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting.
+We compared RS-GLinCB against ECOLog FAJC (22) and GLOC JBNW (17), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting.
 The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$.
@@
 Probit.
-For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB Filippi et al.
-(2010).
+For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB FCGS (10).
 The dimension was set to $d=5$ and number of arms per round to $K=20$.
@@
 To the best of our knowledge, our paper provides the first algorithms for the CB problem with GLM rewards under limited adaptivity (and otherwise) that achieve $\kappa$-independent regret.
-The regret guarantee of RS-GLinCB, not only aligns with the best-known guarantees for Logistic Bandits but enhances them by removing the dependence on $S$ (upper bound on $\left\lVert\theta^{*}\right\rVert$) in the leading term of the regret and therefore resolves a conjecture in Lee et al.
-(2024).
+The regret guarantee of RS-GLinCB, aligns with the best-known guarantees for Logistic Bandits.
 The batch learning algorithm B-GLinCB, for $M=\Omega(\log{(\log{T})})$, achieves a regret of $\widetilde{O}\left(dRS\left(\sqrt{{d/{\widehat{\kappa}}}}\land\sqrt{1/\kappa^{*}}\right)\sqrt{T}\right)$.
@@
 We write c to denote absolute constant(s) that appears throughout our analysis. Our analysis also utilizes the following function
-$\gamma(\lambda)=24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}.$ (9)
+$\gamma(\lambda)=24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+S\sqrt{(1+2RS)\lambda}.$ (10)
 Note that $\gamma(\lambda)$ is a ‘parameterized’ version of $\gamma$ (which was defined in section 3).
@@
 Then, with probability at least than $1-\frac{1}{T^{2}}$, the following inequality holds
-$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}\leq 24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+2S\sqrt{\lambda}$ (11)
+$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}\leq 24RS\left(\sqrt{\log{\left(T\right)}+d}+\frac{R\left(\log{\left(T\right)}+d\right)}{\sqrt{\lambda}}\right)+S\sqrt{(1+2RS)\lambda}$ (12)
 Proof.
 We first define the following quantities
-[…]
-${\mathbf{G}}$ $\coloneqq\sum_{j=1}^{s}\alpha(x,\theta^{*},\widehat{\theta})x_{j}x_{j}^{\mathsf{T}}+\frac{\lambda}{1+2RS}{\mathbf{I}}$
+$\alpha(x,\theta^{*},\widehat{\theta})$ $\coloneqq\int_{v=0}^{1}\dot{\mu}\left(\langle x,\theta^{*}\rangle+v\langle x,~{}\rangle{\left(\widehat{\theta}-\theta^{*}\right)}\right)dv$
+${\mathbf{G}}$ $\coloneqq\sum_{j=1}^{s}\alpha(x,\theta^{*},\widehat{\theta})x_{j}x_{j}^{\mathsf{T}}+\lambda{\mathbf{I}}$
 Using Lemma C.2 we have
@@
 Hence we write
-$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}$ $\leq\sqrt{(1+2RS)}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}}$
+$\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}}$ $\leq\sqrt{1+2RS}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}}$
 $=\sqrt{1+2RS}\left\lVert{\mathbf{G}}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
-$=\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\langle\theta^{*},x_{j}\rangle-\langle\widehat{\theta},x_{j}\rangle\right)\alpha(x,\theta^{*},\widehat{\theta})x_{j}+\frac{\lambda}{1+2RS}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
-$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{\mathbf{G}}^{-1}}+\frac{\lambda}{\sqrt{1+2RS}}\left\lVert\theta^{*}-\widehat{\theta}\right\rVert_{{\mathbf{G}}^{-1}}$
-$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{\mathbf{G}}^{-1}}+2S\sqrt{\lambda}$
-$\leq 3RS\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+2S\sqrt{\lambda}$ (Using (12) and assuming $RS\geq 1$)
-Now by the optimality condition on $\widehat{\theta}$ we have $\sum_{j=1}^{s}\mu\left(\langle x_{j},\widehat{\theta}\rangle\right)x_{j}=\sum_{j=1}^{s}r_{j}x_{j}$ (see equation (3) Filippi et al.
-[2010]).
+$=\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\langle\theta^{*},x_{j}\rangle-\langle\widehat{\theta},x_{j}\rangle\right)\alpha(x,\theta^{*},\widehat{\theta})x_{j}+{\lambda}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
+$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}+{\lambda}\left(\theta^{*}-\widehat{\theta}\right)\right\rVert_{{\mathbf{G}}^{-1}}$ (14)
+$\leq\sqrt{1+2RS}\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}-{\lambda}\widehat{\theta}\right\rVert_{{\mathbf{G}}^{-1}}+\sqrt{1+2RS}\lambda\left\lVert\theta^{*}\right\rVert_{{\mathbf{G}}^{-1}}$ (15)
+$\leq(1+2RS)\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}-{\lambda}\widehat{\theta}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}+S\sqrt{(1+2RS)\lambda}$ (16)
+Now by the optimality condition on $\widehat{\theta}$ we have $\sum_{j=1}^{s}\mu\left(\langle x_{j},\widehat{\theta}\rangle\right)x_{j}+\lambda~{}\widehat{\theta}=\sum_{j=1}^{s}r_{j}x_{j}$.
 Hence, we write
-$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}=\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ (13)
+$\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-\mu\left(\langle\widehat{\theta},x_{j}\rangle\right)\right)x_{j}-\lambda~{}\widehat{\theta}~{}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ $=\left\lVert\sum_{j=1}^{s}\left(\mu\left(\langle\theta^{*},x_{j}\rangle\right)-r_{j}\right)x_{j}~{}\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}$ (17)
 Let ${\mathcal{B}}$ denote the unit ball in $\mathbb{R}^{d}$. We can write
@@
 Proof.
-This result is derived directly from Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). By applying the lemma, we obtain
+This result is derived directly from Lemma A.2 and the definition of $\gamma(\lambda)$ (see 10). By applying the lemma, we obtain
 $|\langle x,\theta^{*}-\widehat{\theta}_{w}\rangle|\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{w}}\right\rVert_{{\mathbf{H}}^{*}}\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\gamma(\lambda)$
 Considering the definition of $\kappa$, we have $\dot{\mu}\left(\langle\mathbf{x},\theta^{*}\rangle\right)\geq\frac{1}{\kappa}$.
-This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$  [^11] For logistic rewards, we note that instead of using the worst case bound of $\kappa$, one can make use of an upper bound on $S\coloneqq\left\lVert\theta\right\rVert$ as done in Mason et al.
-[2022] That is, we lower bound $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\geq\dot{\mu}\left({S\left\lVert x\right\rVert}\right)$ and use ${\mathbf{H}}^{*}\succeq\sum_{t}\dot{\mu}\left({S\left\lVert x_{t}\right\rVert}\right)x_{t}x_{t}^{\intercal}$.
+This implies that ${\mathbf{H}}^{*}\succeq\frac{1}{\kappa}\mathbf{V}$  [^11] For logistic rewards, we note that instead of using the worst case bound of $\kappa$, one can make use of an upper bound on $S\coloneqq\left\lVert\theta\right\rVert$ as done in [20] That is, we lower bound $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)\geq\dot{\mu}\left({S\left\lVert x\right\rVert}\right)$ and use ${\mathbf{H}}^{*}\succeq\sum_{t}\dot{\mu}\left({S\left\lVert x_{t}\right\rVert}\right)x_{t}x_{t}^{\intercal}$.
 which in turn leads to the inequality $\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}}^{-1}}\leq\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$.
@@
 Proof.
-This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 9). According to the lemma, we have
+This result is a direct consequence of Lemma A.2 and the definition of $\gamma(\lambda)$ (see 10). According to the lemma, we have
 $|\langle x,\theta^{*}-\widehat{\theta}_{k}\rangle|$ $\leq\left\lVert x\right\rVert_{{{\mathbf{H}}^{*}_{k}}^{-1}}\left\lVert{\theta^{*}-\widehat{\theta}_{k}}\right\rVert_{{\mathbf{H}}^{*}_{k}}$
@@
 $\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\sum_{\begin{subarray}{c}y\in\{x,x^{*}\}\\ \widetilde{y}\in\{\widetilde{x},\widetilde{x}^{*}\}\end{subarray}}\left\lVert y\right\rVert_{\mathbf{V}^{-1}}\left\lVert\widetilde{y}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}$ $\leq 4\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}\right]$
-$\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Jensen’s inequality)
+$\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Cauchy–Schwarz)
 $\leq 4\sqrt{\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}\right]\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k-1}}\left[\max_{\mathbf{x}\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k-1}^{-1}}^{2}\right]}$ (via Claim A.11)
@@
 Now with the choice of $\lambda=20dR\log T$, we have
-$\gamma(\lambda)\leq\gamma\quad\text{ where }\gamma=30RS\sqrt{d\log{T}}.$
+$\gamma(\lambda)\leq\gamma\quad\text{ where }\gamma=30(RS)^{3/2}\sqrt{d\log{T}}.$
 Substituting $\alpha=T^{\frac{1}{2\left(1-2^{-M}\right)}}$ and $\phi(\lambda)=\frac{\sqrt{\kappa}\ e^{3S}\ \gamma^{2}}{S}$ we get
@@
 ${\mathbf{W}}_{G}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi_{G}({\mathcal{X}})}\left[xx^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
-${\mathbf{W}}_{k}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi_{k}({\mathcal{X}})}\left[\widetilde{x}\widetilde{x}^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
+${\mathbf{W}}_{k}$ $\coloneqq\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\left[\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}(\widetilde{{\mathcal{X}}})}\left[\widetilde{x}\widetilde{x}^{\mathsf{T}}|{\mathcal{X}}\right]\right]$
 Recall, for all batches starting from the second batch ($k\geq 2$), we employ the scaled arm set, denoted as $\widetilde{{\mathcal{X}}}$, for learning and action selection under the Distributional optimal design policy.
@@
 This is due to the fact that the set of surviving arms in batch $k$ is always a smaller set than the previous batches.
-Lemma A.12 (Lemma 4 Ruan et al. [2021]).
+Lemma A.12 (Lemma 4 [23]).
 The expected data matrix ${\mathbf{W}}_{G}$ satisfies. We have
 $\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}^{2}\right]\leq d^{2}$
-Lemma A.13 (Theorem 5 Ruan et al. [2021]).
+Lemma A.13 (Theorem 5 [23]).
 Let the Distributional optimal design $\pi$ which has been learnt from $s$ independent samples ${\mathcal{X}}_{1},\ldots{\mathcal{X}}_{s}\sim{\mathcal{D}}$ and let ${\mathbf{W}}$ denote the expected data matrix, ${\mathbf{W}}=\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\operatorname*{\mathbb{E}}_{x\sim\pi({\mathcal{X}})}\left[xx^{\mathsf{T}}|{\mathcal{X}}\right]\right]$.
@@
 $\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}_{G}}^{2}\right]\leq 2d^{2}.$
-Lemma A.15 (Matrix Chernoff Tropp et al. [2015], Ruan et al. [2021]).
+Lemma A.15 (Matrix Chernoff [27, 23]).
 Let $x_{1},x_{3},\ldots,x_{n}\sim{\mathcal{D}}$ be vectors, with $\left\lVert x_{t}\right\rVert\leq 1$, then we have
@@
 Similarly ${\mathbf{H}}_{k}$, with $\lambda\geq 6\log{(Td)}$ satisfies the following for each batch $k\geq 2$ with probability greater than $1-\frac{1}{T^{2}}$
-${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
+${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}(\widetilde{{\mathcal{X}}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
 Proof.
@@
 ∎
-We note that the analysis of Ruan et al.
-[2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$.
+We note that the analysis of [23] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$.
 We obtain such a bound here and use it in the analysis.
@@
 First, we note from Corollary A.16 that the following holds with high probability
-$\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$
-$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
-We obtain the first two inequalities, (18) and (19), by a direct use of Corollary A.16.
+$\left\lVert x\right\rVert_{\mathbf{V}^{-1}}\leq\frac{8}{\sqrt{\tau_{1}}}\left\lVert x\right\rVert_{{\mathbf{W}}_{G}^{-1}}$
+$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\leq\frac{8}{\sqrt{\tau_{k}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
+We obtain the first two inequalities, (22) and (23), by a direct use of Lemma A.12 and Lemma A.14.
 For (24) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M).
@@
 We will specify the regularizer $\lambda$ later. We also define
-$\gamma\coloneqq\textbf{c}RS\sqrt{d\log{\frac{T}{\delta}}}$
+$\gamma\coloneqq\textbf{c}R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log{\left(\frac{ST}{\delta}\right)}}$
 Below, we state the main concentration bound used in the proof.
-Lemma B.1 (Theorem 1 of Faury et al. [2020]).
+Lemma B.1 (Theorem 1 of [6]).
 Let $\{\mathcal{F}_{t}\}_{t=1}^{\infty}$ be a filtration.
@@
 B.1 Confidence Sets for Switching Criterion I rounds
+Let us define $g_{o}(\theta)=\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda\theta$ and ${\mathbf{H}}_{w}(\theta)=\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
 Lemma B.2.
-At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$. With probability at least $1-\delta$ we have
+At any round $t$, let $\widetilde{\theta}_{o}$ be the unconstrained maximum likelihood estimate calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$ (Algorithm 2 Line 7).
+With probability at least $1-\delta$ we have,
+$\left\lVert g(\widetilde{\theta}_{o})-g(\theta^{*})\right\rVert_{{\mathbf{H}}_{w}^{*-1}}\leq\textbf{c}RS\sqrt{d\log(ST/\delta)}~{}.$
+Proof.
+First note that first order optimality condition of the MLE, we have $g(\widetilde{\theta}_{o})=\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}$. Therefore,
+$\left\lVert g(\widetilde{\theta}_{o})-g(\theta^{*})\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$ $=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}-\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
+$\leq\left\lVert\sum_{s\in{\mathcal{T}}_{o}}(r_{s}-\mu\left(\langle x_{s},\theta^{*}\rangle\right))x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+\left\lVert\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
+$\leq\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta^{\prime}_{s}x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+S\sqrt{\lambda}$ (${\mathbf{H}}_{w}^{*}\succeq\lambda{\mathbf{I}}$ and $\left\lVert\theta^{*}\right\rVert\leq S$; $\eta^{\prime}_{s}\coloneqq r_{s}-\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)$)
+$=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\frac{\eta^{\prime}_{s}}{R}x_{s}\right\rVert_{(\frac{1}{R^{2}}{\mathbf{H}}^{*}_{w})^{-1}}+S\sqrt{\lambda}$
+$\leq\left(\frac{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}{2}+\frac{2d}{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}\log\left(1+\frac{\tau_{o}}{d\left(\nicefrac{{\lambda}}{{R^{2}}}\right)}\right)+\frac{2}{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{\nicefrac{{\lambda}}{{R^{2}}}}}d\log(2)\right)+S\sqrt{\lambda}$ (Lemma B.1 with $\eta_{s}=\frac{\eta^{\prime}_{s}}{R}$)
+Simplifying constants and setting $\sqrt{\frac{\lambda}{R^{2}}}=\textbf{c}\sqrt{d\log(T/\delta)/S}$, we have,
+$\left\lVert g(\widetilde{\theta}_{o})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}$ $\leq\textbf{c}R\sqrt{dS\log(ST/\delta)}~{}.$
+Hereon, we will fix $\lambda=\textbf{c}R\sqrt{dS\log(ST/\delta)}$ as shown in the proof above.
+Recall that $\widehat{\theta}_{o}$ is obtained from $\widetilde{\theta}_{o}$ via the following projection:
+$\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq S}\left\lVert g_{o}(\theta)-g_{o}(\widetilde{\theta}_{o})\right\rVert_{{\mathbf{H}}_{w}(\theta)^{-1}}~{}.$
+Lemma B.3.
+At any round $t$, let $\widehat{\theta}_{o}$ be the projection of the maximum likelihood estimate $\widetilde{\theta}_{o}$ calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$ (Algorithm 2, Lines 7-8).
+With probability at least $1-\delta$ we have
 $\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma.$
 Proof.
-Let us define the matrix ${\mathbf{G}}_{w}=\sum_{s\in{\mathcal{T}}_{o}}\alpha(x,\theta^{*},\widehat{\theta}_{o})x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
-First, we note that by self-concordance property of $\mu$ (Lemma C.2), ${\mathbf{G}}_{w}\succeq\frac{1}{1+2RS}{\mathbf{H}}^{*}_{w}$.
-Hence,
-$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq\sqrt{1+2RS}\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{G}}_{w}}$
-$=\sqrt{1+2RS}\left\lVert{\mathbf{G}}_{w}\left(\widehat{\theta}_{o}-\theta^{*}\right)\right\rVert_{{\mathbf{G}}_{w}^{-1}}$
-$=\sqrt{1+2RS}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\langle\widehat{\theta}_{o},x_{s}\rangle-\langle\theta^{*},x_{s}\rangle\right)\alpha(x_{s},\widehat{\theta}_{o},\theta^{*})x_{s}+\lambda\widehat{\theta}_{o}-\lambda_{t}\theta^{*}\right\rVert_{{\mathbf{G}}_{w}^{-1}}$
-$=\sqrt{1+2RS}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle\widehat{\theta}_{o},x_{s}\rangle\right)-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}+\lambda\widehat{\theta}_{o}-\lambda_{t}\theta^{*}\right\rVert_{{\mathbf{G}}_{w}^{-1}}$ (Taylor’s theorem)
-$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle\widehat{\theta}_{o},x_{s}\rangle\right)-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}+\lambda\widehat{\theta}_{o}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$ (${\mathbf{G}}_{w}\succeq\frac{1}{1+2RS}{\mathbf{H}}^{*}_{w}$)
-Since $\widehat{\theta}_{o}$ is the maximum likelihood estimate, by optimality condition, we have the following relation: $\sum_{s\in{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\widehat{\theta}_{o}\rangle\right)x_{s}+\lambda\widehat{\theta}_{o}=\sum_{s\in{\mathcal{T}}_{o}}r_{s}x_{s}$.
-Substituting this above, we get
-$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}-\lambda\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
-$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+\lambda\left\lVert\theta^{*}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}$
-$\leq(1+2RS)\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta_{s}x_{s}\right\rVert_{{\mathbf{H}}_{w}^{*-1}}+S\sqrt{\lambda}.$ (${\mathbf{H}}^{*}_{w}\succeq\lambda_{t}{\mathbf{I}}$, $\left\lVert\theta^{*}\right\rVert_{2}\leq S$)
-where $\eta_{s}\coloneqq\left(r_{s}-\mu\left(\langle\theta^{*},x_{s}\rangle\right)\right)$.
-We will now apply Lemma B.1 $\eta_{s}$ scaled by $R$.
-First note that $\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\eta_{s}x_{s}\right\rVert_{({\mathbf{H}}^{*}_{w})^{-1}}=\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\frac{\eta_{s}}{R}x_{s}\right\rVert_{(R^{2}{\mathbf{H}}^{*}_{w})^{-1}}$ which, in turn ensures that the noise variable is upper bounded by $1$.
-Applying B.1 we get
-$\lVert\widehat{\theta}_{o}-\theta^{*}$ $\rVert_{{\mathbf{H}}^{*}_{w}}\leq S\sqrt{R^{2}\lambda}$
-$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{o}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
-Simplifying constants and setting $\sqrt{R^{2}\lambda}=\textbf{c}\sqrt{d\log(T)+\log(1/\delta)}$, we have $\left\lVert\widehat{\theta}_{o}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{w}}\leq\textbf{c}RS\sqrt{d\log(T/\delta)+\log(1/\delta)}\leq\textbf{c}RS\sqrt{d\log{(T/\delta)}}$.
-∎
+Observe that $\lvert\langle x_{t},\theta^{*}-\widehat{\theta}_{o}\rangle\rvert\leq 2S$ for all $t\in{\mathcal{T}}_{o}$. Hence, by Lemma E.1 with $c=2RS$,
+$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}$ $\leq 2(1+2RS)\textbf{c}R\sqrt{dS\log(ST/\delta)}+2R\sqrt{S(1+2RS)}\leq\textbf{c}R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log(ST/\delta)}~{}.$
 B.2 Confidence Sets for non-Switching Criterion I rounds
@@
 $\leq\frac{1}{R}$
-Recall that ${\mathbf{H}}_{t}$ is defined in line 15 of Algorithm 2. Further, we define
+Recall that ${\mathbf{H}}_{t}$ is defined in line 16 of Algorithm 2. Further, we define
 ${\mathbf{H}}^{*}_{t}=\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$
@@
 ∎
-Recall that $\tau$ is the round when the Switching Criterion II (Line 9) is satisfied. Now we define the following quantities:
-$g_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda_{\tau}\theta$
+Recall that $\tau$ is the round when the Switching Criterion II (Line 10) is satisfied. Now we define the following quantities:
+$g_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda\theta$
 ${{\mathbf{H}}}_{\tau}(\theta)$ $=\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$
 $\Theta$ =$=\left\{\theta:\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}\right\}$
-$\tilde{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})$
-$\beta$ $\coloneqq\textbf{c}\sqrt{d\log(T/\delta)}$
+$\tilde{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\lambda\left\lVert\theta\right\rVert_{2}^{2}$
+$\beta$ $\coloneqq\textbf{c}R\sqrt{dS\log(ST/\delta)}$
 Moreover, recall the following definition $\widehat{\theta}_{\tau}$:
@@
 Also note that $\theta^{*}\in\Theta$. Hence,
-$\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}$ $\leq 2(1+2)\left\lVert g_{\tau}(\theta^{*})-\sum_{s\in[\tau-1]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}$ (by Lemma E.1)
-$\leq 6\textbf{c}\sqrt{d\log(T/\delta)}$ (by Lemma B.1)
+$\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}$ $\leq 2\left(1+2\right)\left\lVert\sum_{s\in[\tau-1]\setminus{\mathcal{T}}_{o}}\left(r_{s}-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{H}}_{\tau}^{*-1}}+2S\sqrt{\lambda\left(1+2\right)}$ (by Lemma E.1)
+$\leq\textbf{c}R\sqrt{dS\log(ST/\delta)}$ (by Lemma B.1)
 Let the event in lemma B.7 be denoted by ${\mathcal{E}}_{\tau}$, or in other words, ${\mathcal{E}}_{\tau}=\{\lVert\widehat{\theta}_{\tau}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta\}$.
@@
 In this subsection, we will only consider rounds $t\in[T]$ which does not satisfy Switching Criterion I.
-Let $x_{t}\in{\mathcal{X}}_{t}$ be the played arm defined via line 14 Algorithm 2.
+Let $x_{t}\in{\mathcal{X}}_{t}$ be the played arm defined via line 15 Algorithm 2.
 Further, let $x^{*}_{t}\in{\mathcal{X}}_{t}$ be the best available arm in that round.
@@
 $\langle x^{*}_{t},\theta^{*}\rangle-\langle x_{t},\theta^{*}\rangle$ $\leq\left(\langle x^{*}_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x^{*}_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}\right)$ (by Corollary B.8)
-$\leq\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)$ (optimistic $x_{t}$, see line 14 algo.
+$\leq\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle+\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)-\left(\langle x_{t},\widehat{\theta}_{\tau}\rangle-\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{\tau}}\right)$ (optimistic $x_{t}$, see line 15 algo.
 2)
@@
 Lemma B.10.
-The arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$, (b) $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
+The arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 14 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$, (b) $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
 Proof.
@@
 Theorem B.11 (Theorem 4.2).
-Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}d\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{RT}/{\delta}}\right)+\ \kappa d^{2}R^{5}S^{2}\log^{2}\left({{T}/{\delta}}\right)\big{)}.$
+Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}dRS^{\nicefrac{{1}}{{2}}}\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{ST}/{\delta}}\right)+\ \kappa d^{2}R^{7}S^{3}\log^{2}\left({{ST}/{\delta}}\right)\big{)}.$
 Proof.
@@
 $\leq R\tau_{o}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$ (Upper bound of $R$ for rounds in ${\mathcal{T}}_{o}$)
-$\leq\textbf{c}R^{3}\kappa\gamma^{2}\log(T/\delta)+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ (some $z\in[\langle x_{t},\theta^{*}\rangle,\langle x_{t}^{*},\theta^{*}\rangle]$; lemma B.11)
+$\leq 2dR^{3}\kappa\gamma^{2}\log{(ST/\delta)}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ (some $z\in[\langle x_{t},\theta^{*}\rangle,\langle x_{t}^{*},\theta^{*}\rangle]$; lemma B.12)
 Now, let $R_{1}(T)=\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}(z)\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$.
@@
 $\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot 2d\log\left(1+\frac{RT}{\lambda d}\right)}$ (Lemma B.14; $\left\lVert\tilde{x}_{t}\right\rVert_{2}\leq R$)
-$\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}.$
+$\leq\textbf{c}RS^{\nicefrac{{1}}{{2}}}d\log(ST/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}.$
 Putting things back,
-$\textsc{R}_{T}\leq\textbf{c}d\log(RT/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}+\textbf{c}R^{5}S^{2}\kappa\log(T/\delta)^{2}.$
+$\textsc{R}_{T}\leq\textbf{c}RS^{\nicefrac{{1}}{{2}}}d\log(ST/\delta)\sqrt{\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)}+\textbf{c}R^{7}S^{3}\kappa d^{2}\log(ST/\delta)^{2}.$
 B.5 Bounding number of policy updates: Proof of Lemma 4.1
@@
 Combining (27) and (26) we have
-$\tau_{o}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{o}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
-Lemma (4.1).
-Algorithm 2, during its entire execution, updates its policy at most $O(R^{4}S^{2}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
+$\tau_{o}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{o}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(ST/\delta)}$
+Lemma B.13.
+Algorithm 2, during its entire execution, updates its policy at most $O(R^{6}S^{3}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
 Proof.
@@
 On the other hand, the number of times Switching Criterion II is triggered is equal to the number of times determinant of ${\mathbf{H}}_{t}$ doubles, which is bounded by Lemma B.17.
-Thus in total, the number of policy changes in Algorithm 2 is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
+Thus in total, the number of policy changes in Algorithm 2 is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(ST/\delta)+\textbf{c}d\log(T)~{}.$
 B.6 Some Useful Lemmas
-Lemma B.12 (Elliptic Potential Lemma (Lemma 10 Abbasi-Yadkori et al. [2011])).
+Lemma B.14 (Elliptic Potential Lemma (Lemma 10 [4])).
 Let $x_{1},x_{2},\ldots x_{t}$ be a sequence of vectors in $\mathbb{R}^{d}$ and let $\left\lVert x_{s}\right\rVert_{2}\leq L$ for all $s\in[t]$.
@@
 $\sum_{s=1}^{t}\left\lVert x_{s}\right\rVert_{\mathbf{V}_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{L^{2}t}{\lambda d}\right)$ (28)
-Lemma B.13 (Lemma 12 of Abbasi-Yadkori et al. [2011]).
+Lemma B.15 (Lemma 12 of [4]).
 Let $A\succeq B\succ 0$. Then
 $\sup_{x\neq 0}\frac{x^{\intercal}Ax}{x^{\intercal}Bx}\leq\frac{\det(A)}{\det(B)}$
-Lemma B.14 (Lemma 10 of Abbasi-Yadkori et al. [2011]).
+Lemma B.16 (Lemma 10 of [4]).
 Let $\{x_{s}\}_{s=1}^{t}$ be a set of vectors.
@@
 $=R\ddot{b}(z)$ (Lemma C.3)
-As a consequence, we have the following simple modification of the self-concordance results of Faury et al. [2020].
+As a consequence, we have the following simple modification of the self-concordance results of [6].
 Lemma C.2.
@@
 $\dot{\mu}(z_{1})\exp(-R(z-z_{1}))\leq$ $\dot{\mu}(z)\leq\dot{\mu}(z_{1})\exp(R(z-z_{1}))$
-Putting $z=z_{2}$ establishes 26. To show 25, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
+Putting $z=z_{2}$ establishes 30. To show 29, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
 […]
@@
 Note that similar exercise can be repeated with $z_{2}\leq z_{1}$ to get the same result for $z_{2}$.
-For 27, we have, by application of 26, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
+For 31, we have, by application of 30, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
 […]
@@
 $=\dot{\mu}(z_{1})\left(\frac{1}{R|z_{1}-z_{2}|}+\frac{\exp(-R|z_{1}-z_{2}|)-1}{R^{2}|z_{1}-z_{2}|^{2}}\right)$
-$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of Abeille et al. [2021])
+$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of [2])
 Next we state some nice properties of the GLM family that is the key in deriving Lemma C.1.
@@
 Remark C.5.
-In Faury [2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set.
-Thereafter the author notes that the techniques developed in Faury [2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter).
+In [8] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set.
+Thereafter the author notes that the techniques developed in [8] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter).
 However, the claim regarding self-concordance of GLMs is not true in general.
@@
 Let the maximum number of arms available every round be $K$.
-Furher, let the computational cost of an oracle that solves the non-convex optimization 8 be $NC_{opt}$.
+Furher, let the computational cost of an oracle that solves the non-convex optimization 9 be $NC_{opt}$.
 Computational Cost of B-GLinCB: In the B-GLinCB algorithm, we employ the log-loss oracle at the end of each batch.
@@
 We allowed no operations for 10 seconds after every run to let the CPU temperature come back to normal, in case the execution heats up the CPU, thereby causing subsequent runs to slow down.
-Comparison with ECOLog Faury et al.
-[2022] shows that execution time for RS-GLinCB is significantly smaller.
+Comparison with ECOLog [7] shows that execution time for RS-GLinCB is significantly smaller.
 We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round.
@@
 Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
-While comparing with GLOC Jun et al.
-[2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes.
+While comparing with GLOC [14], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes.
 Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$.
@@
 Appendix E Projection
-We describe the projection step used in Algorithms 1 and 2. We present arguments similar to the ones made in Appendix B.3 of Faury et al. [2020]. We write
+We describe the projection step used in Algorithms 1 and 2. We present arguments similar to the ones made in Appendix B.3 of [6]. We write
 ${\mathbf{H}}(\theta)=\sum_{s=1}^{t}\dot{\mu}\left(\langle\theta,x_{s}\rangle\right)x_{s}x_{s}^{\mathsf{T}}+\lambda{\mathbf{I}}$
@@
 We project $\widehat{\theta}$ to a set $\Theta$ by solving the following optimization problem
-$\widetilde{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widehat{\theta}\rangle\right))x_{s}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}$ (28)
+$\widetilde{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}\left\lVert\sum_{s=1}^{t}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widehat{\theta}\rangle\right)\right)x_{s}-\lambda\widehat{\theta}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}$ (32)
 Lemma E.1.
 Using the notations described above, if $\theta^{*}\in\Theta$ and $\max_{i\in[t]}|\langle x_{i},\widetilde{\theta}-\theta^{*}\rangle|\leq c/R$, then we have
-$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}(\theta^{*})}\leq 2(1+c)\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta^{*}\rangle\right)-r_{s})x_{s}\right\rVert_{{\mathbf{H}}(\theta^{*})^{-1}}$
+$\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}(\theta^{*})}\leq 2(1+c)\left\lVert\sum_{s=1}^{t}(\mu\left(\langle x_{s},\theta^{*}\rangle\right)-r_{s})x_{s}\right\rVert_{{\mathbf{H}}(\theta^{*})^{-1}}+2S\sqrt{\lambda(1+c)}$
 Proof.
@@
 Similarly, we have $\alpha(x_{s},\widetilde{\theta},\theta^{*})\geq\frac{\dot{\mu}\left(\langle x_{s},\widetilde{\theta}\rangle\right)}{1+c}$.
-Let us define the matrix ${\mathbf{G}}=\sum_{s\in[t]}\alpha(x,\widetilde{\theta},\theta^{*})x_{s}x_{s}^{\intercal}$.
+Let us define the matrix ${\mathbf{G}}=\sum_{s\in[t]}\alpha(x,\widetilde{\theta},\theta^{*})x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
 Using the above fact, we obtain the relation: ${\mathbf{G}}\succeq\frac{1}{1+c}{\mathbf{H}}^{*}$ and ${\mathbf{G}}\succeq\frac{1}{1+c}{{\mathbf{H}}}(\widetilde{\theta})$.
-Also define the vector $g(\theta)=\sum_{s\in[t]}\mu\left(\langle\theta,x_{s}\rangle\right)x_{s}$.
 Now,
@@
 $=\sqrt{1+c}\left\lVert{\mathbf{G}}\left(\widetilde{\theta}-\theta^{*}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
-$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\alpha(x_{s},\widetilde{\theta},\theta^{*})\langle\widetilde{\theta}-\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}$
-$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\mu\left(\langle x_{s},\widetilde{\theta}\rangle\right)-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}$ (Taylor’s theorem)
-$=\sqrt{1+c}\left\lVert\left(\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}\right)-\left(\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
-Let $g(\theta)=\sum_{s=1}^{t}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}$ for any $\theta$. Therefore, we have,
+$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\alpha(x_{s},\widetilde{\theta},\theta^{*})\langle\widetilde{\theta}-\theta^{*},x_{s}\rangle\right)x_{s}+\lambda(\widetilde{\theta}-\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
+$=\sqrt{1+c}\left\lVert\sum_{s\in[t]}\left(\mu\left(\langle x_{s},\widetilde{\theta}\rangle\right)-\mu\left(\langle x_{s},\theta^{*}\rangle\right)\right)x_{s}+\lambda(\widetilde{\theta}-\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$ (Taylor’s theorem)
+$=\sqrt{1+c}\left\lVert\left(\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}+\lambda\widetilde{\theta}\right)-\left(\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}+\lambda\theta^{*}\right)\right\rVert_{{\mathbf{G}}^{-1}}$
+Let $g(\theta)=\sum_{s=1}^{t}\mu\left(\langle x_{s},\theta\rangle\right)x_{s}+\lambda\theta$ for any $\theta$. Therefore, we have,
 $\left\lVert\widetilde{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}$ $\leq\sqrt{1+c}\left\lVert g(\widetilde{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
-$=\sqrt{1+c}\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})+g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}$
+$\leq\sqrt{1+c}\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})+g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}\right)$
 $\leq\sqrt{1+c}\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}^{-1}}+\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{G}}^{-1}}\right)$ ($\triangle$ inequality)
-$\leq(1+c)\left(\left\lVert g(\widetilde{\theta})-g(\widehat{\theta})\right\rVert_{{{\mathbf{H}}(\widetilde{\theta})}^{-1}}+\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}\right)$ (${\mathbf{H}}^{*-1}\succeq(\sqrt{1+c}){\mathbf{G}}^{-1}$)
-$\leq 2(1+c)\left\lVert g(\widehat{\theta})-g(\theta^{*})\right\rVert_{{\mathbf{H}}^{*-1}}$ (by (28))
-$=2(1+c)\left\lVert g(\theta^{*})-\sum_{s\in[t]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}$ ($\widehat{\theta}$ is the unconstrained MLE, $g(\widehat{\theta})=\sum_{s\in[t]}r_{s}x_{s}$.)
+$\leq\sqrt{1+c}\left(\left\lVert\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}^{-1}}+\left\lVert g(\widehat{\theta})-\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{G}}^{-1}}\right)+2S\sqrt{\lambda(1+c)}$ ($\triangle$ inequality)
+$\leq(1+c)\left(\left\lVert\sum_{s\in[t]}\mu\left(\langle\widetilde{\theta},x_{s}\rangle\right)x_{s}-g(\widehat{\theta})\right\rVert_{{{\mathbf{H}}(\widetilde{\theta})}^{-1}}+\left\lVert g(\widehat{\theta})-\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}\right)+2S\sqrt{\lambda(1+c)}$ (${\mathbf{H}}^{*-1}\succeq({1+c}){\mathbf{G}}^{-1}$)
+$\leq 2(1+c)\left\lVert g(\widehat{\theta})-\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}+2S\sqrt{\lambda(1+c)}$ (by (32))
+$=2(1+c)\left\lVert\sum_{s\in[t]}\mu\left(\langle\theta^{*},x_{s}\rangle\right)x_{s}-\sum_{s\in[t]}r_{s}x_{s}\right\rVert_{{\mathbf{H}}^{*-1}}+2S\sqrt{\lambda(1+c)}$ ($\widehat{\theta}$ is the unconstrained MLE, $g(\widehat{\theta})=\sum_{s\in[t]}r_{s}x_{s}+\lambda\widehat{\theta}$.)
 E.1 Convex Relaxation
@@
 However, it is possible to substitute this optimization problem with a convex one, whose computational complexity can be better tractable.
-The process is similar to the one detailed in [Abeille et al., 2021, section 6].
+The process is similar to the one detailed in [2, section 6].
 Here we briefly outline the steps.
-Let ${\mathcal{L}}_{t}(\theta)=\sum_{s=1}^{t}\ell(\theta,x_{s},r_{s})$ and $\breve{\theta}$ be defined as follows:
+Let ${\mathcal{L}}_{t}(\theta)=\sum_{s=1}^{t}\ell(\theta,x_{s},r_{s})+\lambda\left\lVert\theta^{*}\right\rVert^{2}$ and $\breve{\theta}$ be defined as follows:
 $\breve{\theta}\coloneqq\operatorname*{arg\,min}_{\theta\in\Theta}{\mathcal{L}}_{t}(\theta)$ (33)
@@
 $\tilde{\alpha}(x_{s},\theta^{*},\breve{\theta})$ $\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R|\langle x_{s},\breve{\theta}-\theta^{*}\rangle|}$ (Lemma C.2)
-$\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R(c/R)}$ ($\max_{i\in[s]}|\langle x_{s},\widetilde{\theta}-\theta^{*}\rangle|\leq c/R$)
+$\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R(c/R)}$ ($\max_{i\in[s]}|\langle x_{s},\breve{\theta}-\theta^{*}\rangle|\leq c/R$)
 $=\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+c}$
-Let us define $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\coloneqq\sum_{s=1}^{t}\tilde{\alpha}(x_{s},\theta^{*},\theta)x_{s}x_{s}^{\intercal}$.
-Using the above fact, we obtain $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\succeq\frac{1}{2+c}{\mathbf{H}}^{*}$.
-We now follow closely the proof outlined in Appendix B.3 of Abeille et al. [2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
-${\mathcal{L}}_{t}(\theta)-{\mathcal{L}}_{t}(\theta^{*})-\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\theta-\theta^{*}\rangle$ $=\left\lVert\theta-\theta^{*}\right\rVert_{\tilde{{\mathbf{G}}}(\theta,\theta^{*})}^{2}$
-$\geq\frac{1}{2+c}\left\lVert\theta-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}$
+Let us define $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\coloneqq\sum_{s=1}^{t}\tilde{\alpha}(x_{s},\theta^{*},\theta)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
+Using the above fact, we obtain $\tilde{{\mathbf{G}}}(\theta^{*},\breve{\theta})\succeq\frac{1}{2+c}{\mathbf{H}}^{*}$.
+We now follow closely the proof outlined in Appendix B.3 of [2] with minor changes. By second-order Taylor’s expansion, we can write
+${\mathcal{L}}_{t}(\breve{\theta})-{\mathcal{L}}_{t}(\theta^{*})-\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\theta-\theta^{*}\rangle$ $=\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{\tilde{{\mathbf{G}}}(\breve{\theta},\theta^{*})}^{2}$
+$\geq\frac{1}{2+c}\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}$
 Taking absolute value on both sides, and substituting $\theta=\breve{\theta}$,
@@
 Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$.
-By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of Abeille et al.
-[2021]), we have, for $\theta^{*}$.
-${\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})$ $\leq\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}(\theta^{*},\widehat{\theta})^{-1}}^{2}$
-$\leq\frac{R}{\sqrt{\lambda}}\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}^{2}+\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$
+By a similar Taylor expansion as above and some algebraic manipulations (Lemma 2 in Appendix B.4  [^12] The proof proceeds analogously, except that we retain the upper bound in terms of $\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$ instead of substituting it with $\gamma(\delta)$ at the end.
+and Appendix B.3 of [2] ), we have, for $\theta^{*}$.
+${\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})$ $\leq\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}(\theta^{*},\widehat{\theta})^{-1}}^{2}$ (see Appendix B.3 of [2])
+$\leq\frac{R}{\sqrt{\lambda}}\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}^{2}+\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$ (Lemma 2 in Appendix B.4 [2] )
 $\leq\frac{R}{\sqrt{\lambda}}\gamma^{2}+\gamma$ (Lemma B.1)